### v1

# jmstate, a Flexible Python Package for Multi-State Joint Modeling

Abstract
Classical joint modeling approaches often rely on competing risks or recurrent event formulations to account for complex real-world processes involving evolving longitudinal markers and discrete event occurrences. However, these frameworks typically capture only limited aspects of the underlying event dynamics.
Multi-state joint models offer a more flexible alternative by representing full event histories through a network of possible transitions, including recurrent cycles and terminal absorptions, all potentially influenced by longitudinal covariates.
In this paper, we propose a general framework that unifies longitudinal biomarker modeling with multi-state event processes defined on arbitrary directed graphs. Our approach accommodates both Markovian and semi-Markovian transition structures, and extends classical joint models by coupling nonlinear mixed-effects longitudinal submodels with multi-state survival processes via shared latent structures.
We derive the full likelihood and develop scalable inference procedures based on stochastic gradient descent. Furthermore, we introduce a dynamic prediction framework, enabling individualized risk assessments along complex state-transition trajectories.
To facilitate reproducibility and dissemination, we provide an open-source Python library jmstate implementing the proposed methodology, available on PyPI. Simulation experiments and a biomedical case study demonstrate the flexibility and performance of the framework in representing complex longitudinal and multi-state event dynamics. The full Python notebooks used to reproduce the experiments as well as the source code of this paper are available on GitLab.
Keywords joint modeling • multi-state processes • longitudinal data • survival analysis • stochastic gradient descent • dynamic prediction • Python library
1 Introduction
Joint modeling of longitudinal and time-to-event data has become an essential tool of modern biostatistics [Papageorgiou et al., 2019], particularly for dynamic prediction in clinical applications. Classical joint models typically couple a longitudinal biomarker process with a single time-to-event outcome [Rizopoulos, 2012], allowing for the integration of biological knowledge and individual heterogeneity via shared latent structures. However, many real-world processes involve multiple possible outcomes, intermediate stages, or recurrent events, which cannot be fully captured by a single-event framework [Król et al., 2017]. In such settings, multi-state models provide a more flexible approach [Ferrer et al., 2016].
Multi-state models represent an overall joint process of clinical course as a series of discrete stages or health states that occur sequentially [Lovblom et al., 2024]. In biostatistics, these models are widely used for survival and reliability analysis, allowing for a richer and more accurate representation by capturing alternative paths to an event of interest, intermediate events, and progressive disease. Key components of multi-state models include transition intensity functions, which denote the instantaneous risk of moving from one state to another, and transition probability functions, which describe the probability of transition over longer intervals. Often, these models assume the Markov property, where future transitions depend only on the current state, simplifying the transition intensity functions [Asanjarani et al., 2022].
However, semi-Markov processes extend this by allowing future probability transitions to depend on the sojourn time, with the clock resetting after each transition.
The link between multi-state models and joint models arises when a multi-state process is integrated as a component within a broader joint modeling framework . While multi-state models, such as those described in the “sequential state framework” primarily focus on movements between discrete states, joint models operate under a “parallel trajectory framework” that combines a longitudinal process with a time-to-event process.
The model proposed by Ferrer et al. [Ferrer, Rondeau, Dignam, Pickles, Jacqmin-Gadda, and Proust-Lima, 2016] exemplifies this link by presenting a joint model for a longitudinal process (e.g., Prostate-Specific Antigen measurements) and a multi-state process (e.g., clinical progressions in prostate cancer). These two sub-models are interconnected by shared random effects, allowing the model to account for the correlation between the continuous longitudinal biomarker trajectory and the discrete transitions between health states.
Several software packages have been developed in R and Python to fit joint models that combine longitudinal and time-to-event data. The JMbayes package [Rizopoulos, 2016] offers a Bayesian implementation of joint models using Markov chain Monte Carlo (MCMC) methods. Its successor, JMbayes2 [Rizopoulos et al., 2024], extends this framework to accommodate more complex data structures, such as multiple longitudinal outcomes, nonlinear trajectories, and competing risks. Another popular R package, joineR [Philipson et al., 2018], enables joint modeling of repeated measurements and survival outcomes, based on the methodology introduced by Williamson et al. [2008]. In the Python ecosystem, the flash package [Nguyen et al., 2024] provides a scalable joint modeling approach tailored to high-dimensional settings, exploiting sparse association patterns between longitudinal and survival processes.
Despite the richness of these tools, none currently supports joint models in the multi-state framework with nonlinear longitudinal modeling. In this work, we build on the approach of Ferrer et al. [Ferrer, Rondeau, Dignam, Pickles, Jacqmin-Gadda, and Proust-Lima, 2016] and introduce jmstate, a Python package for joint modeling of nonlinear longitudinal and multi-state time-to-event data. The underlying model of jmstate incorporates a linear mixed-effects submodel for the longitudinal biomarker and a multi-state process for the event history, linked through shared random effects as in [Ferrer, Rondeau, Dignam, Pickles, Jacqmin-Gadda, and Proust-Lima, 2016].
The model we propose operates on an arbitrary directed graph, enabling representation of complex state transitions and recurrent events. We derive the full likelihood for the nonlinear joint model, introduce an efficient stochastic approximation inference method, and develop dynamic prediction tools. To illustrate its practical relevance, we conduct both a simulation study and a synthetic biomedical case study using the jmstate PyPI package.
This paper is organized as follows. Section 2 presents background on joint modeling frameworks and multi-state processes. Section 3 introduces the proposed multi-state joint model, detailing its likelihood formulation and inference strategies. Section 4 develops dynamic prediction tools. Section 5 details efficient inference methods based on stochastic approximation. Section 6 provides implementation details. Section 7 exemplifies the use of jmstate though a simulation study while Section 8 illustrates the use of jmstate on data from the PAQUID cohort. Finally, Section 9 concludes with a discussion of the main strengths, limitations, and future directions of this work.
2 Background
2.1 Notations
Each individual $i$, where $i=1,\ldots,n$, is observed at times $t_{ij}$, where $j=1,\ldots,n_{i}$. $n_{i}$ is the number of observations for an individual $i$. The biomarker values at these times are denoted by $Y_{ij}\in\mathbb{R}^{d}$, where $d$ is the number of biomarkers. Each individual also has a set of covariates $X_{i}\in\mathbb{R}^{k}$, which may include demographic or clinical characteristics.
2.2 Standard Joint Modeling Framework
Joint modeling frameworks are essential to link longitudinal and time-to-event data for more accurate dynamic predictions. The standard framework [Rizopoulos, 2012, 2011] consists of:
  • Mixed-Effects for Longitudinal Data. The longitudinal process $Y_{ij}$ for subject $i$ at time $t_{ij}$ is typically modeled as:
    $Y_{ij}=h(t_{ij},\psi_{i})+\epsilon_{ij},\quad\psi_{i}=f(\gamma,X_{i},b_{i}),$
    where $\gamma$ are the population parameters, $b_{i}$ are individual random effects, and $\epsilon_{ij}$ is a random measurement error term. It is usually assumed to be Gaussian with zero mean. The functions $h$ and $f$ can be nonlinear, capturing complex within-subject trajectories and between-subject variability. In practice, $h$ is often a linear predictor of the biomarker values, while $f$ may represent a nonlinear function of the random effects.
    Both $h$ and $f$ may be nonlinear, allowing flexible modeling of complex biomarker dynamics. Typical choices include:
    • $h(t,\psi_{i})=\psi_{i1}+\psi_{i2}t+\psi_{i3}t^{2}$ (polynomial growth),
    • $h(t,\psi_{i})=\psi_{i1}\exp(-\psi_{i2}t)$ (exponential decay),
    • $f(\gamma,X_{i},b_{i})=\gamma X_{i}+b_{i}$ [Davidian and Giltinan, 2003], $\gamma+b_{i}$, or $\exp(\gamma+b_{i})$ for log-normally distributed parameters.
  • Cox-Based Survival Modeling. The survival process is described by a hazard function incorporating the current (and possibly historical) longitudinal process and baseline covariates:
    $\lambda_{i}(t\mid X_{i},\mathcal{H}_{i}(t))=\lambda_{0}(t)\exp\left(\alpha\,g(t,\psi_{i})+\beta X_{i}\right),$
    where $\mathcal{H}_{i}(t)$ is the information that would be provided by knowing the "true trajectory" of the longitudinal biomarker process, that is, $\mathcal{H}_{i}(t)=\{(u,h(u,\psi_{i})):0\leq u\leq t\}$ and $g(\cdot)$ links the biomarker trajectory to the hazard, with $\alpha$ et $\beta$ as association parameters. The baseline hazard $\lambda_{0}(t)$ may be parametric or semiparametric with a spline basis representation [Andrinopoulou et al., 2018].
  • Prior distribution. A prior has to be specified for the latent variables $b_{i}$. Here, we assume it to be Gaussian with zero mean: $\mathcal{N}(0,Q)$ but other prior distributions can be considered.
In joint modeling frameworks, the longitudinal process is linked to the time-to-event component through a function $g$ of the underlying trajectory. Several choices for $g$ have been proposed, depending on the hypothesized biological mechanism. Current examples include [Papageorgiou et al., 2019, Mauff, 2023]:
  • $g(t,\psi_{i})=h(t,\psi_{i})$: direct effect of the current biomarker level,
  • $g(t,\psi_{i})=\frac{\partial}{\partial t}h(t,\psi_{i})$: effect of the biomarker slope (rate of change),
  • $g(t,\psi_{i})=\int_{0}^{t}h(w,\psi_{i})\,dw$: cumulative exposure.
This joint framework not only captures the dynamics of longitudinal biomarkers but also leverages their predictive power to inform survival outcomes, enabling dynamic, individualized risk predictions [Rizopoulos, 2011, Ibrahim et al., 2010].
Moreover, in order to derive an expression for the marginal (joint) likelihood, joint modeling typically requires the following independence assumptions [Rizopoulos, 2012]:
Assumption 1.
The latent variables $(b_{i})_{i}$ are independent across individuals.
Assumption 2.
The observed markers $(Y_{ij})_{ij}$ are independent conditionally on $(b_{i})_{i}$.
Assumption 3.
The occurrence times $(T_{i}^{*})_{i}$ are independent conditionally on $(b_{i})_{i}$.
Assumption 4.
The occurrence times $(T_{i}^{*})_{i}$ and the observed markers $(Y_{ij})_{ij}$ are mutually independent conditionally on $(b_{i})_{i}$.
Assumption 5.
The censoring times $(C_{i})_{i}$ are independent and noninformative: conditional on $(b_{i})_{i}$ (and possibly covariates), $(C_{i})_{i}$ is independent of both the event times $(T_{i}^{*})_{i}$ and the longitudinal measurements $(Y_{ij})_{ij}$.
These assumptions facilitate the derivation of the marginal likelihood by allowing the decomposition of the joint distribution into manageable components. They ensure that the latent variables capture all relevant dependencies between the longitudinal and survival processes, enabling efficient estimation and inference.
2.3 Multi-State Markov Processes
A multi-state stochastic process is defined as a process $S(t)$ for $t\geq 0$, where $S(t)$ can take a finite number of values (states), often labeled $1,2,\dots,p$. Quantities of interest typically include the probability of being in a certain state at a given time and the distribution of first passage times.
A Markov process (or continuous-time Markov chain) is a specific class of multi-state models in which future transitions between states depend only upon the current state. This Markov property means the process is memoryless. A key consequence is that the duration spent in any state follows an exponential distribution, implying a constant hazard rate for leaving that state [Jackson, 2011, Putter et al., 2007].
However, the Markov assumption can be unrealistic in many real-world applications. For example, in the study of human sleep stages, sojourn times often do not follow an exponential distribution [Dong et al., 2008, Roever et al., 2010], and in the case of chronic diseases such as AIDS, the risk of disease progression can depend on the time elapsed since infection [Andersen and Keiding, 1999]. To address these limitations, semi-Markov processes (SMPs) were introduced, which allow for arbitrary distributions of sojourn times while retaining the Markov property for the embedded discrete-time chain. This flexibility makes SMPs suitable for modeling complex disease progression and patient recovery scenarios [Putter et al., 2007].
2.4 Multi-State Semi-Markov Processes
Multi-state semi-Markov processes (MSMPs) offer a natural generalization by allowing the distribution of sojourn times to be arbitrary.
In an MSMP, the process is defined by a directed graph $G=(V,E)$, where $V$ represents the set of states and $E$ the set of possible transitions. The transition intensities $\lambda_{k^{\prime}|k}(t\mid t_{0})$ from state $k$ to state $k^{\prime}$ at time $t$, given entry time $t_{0}$. The sojourn time in state $k$ is not restricted to an exponential distribution, allowing for more realistic modeling of the time spent in each state.
MSMPs have been widely applied in a range of disciplines. In reliability, they are used to modeling degradation and repair processes [Limnios and Oprisan, 2001]. In biomedical studies, they have proven useful for analyzing illness-death models or disease progression with non-exponential transitions [Commenges et al., 2006, Fiocco et al., 2008]. Applications in finance also exist, where semi-Markovian dynamics can model credit rating migrations or economic regimes [Luciano and Vasiliev, 2006]. In these contexts, MSMPs retain the Markov property in the embedded jump chain while offering increased realism through flexible dwell time modeling.
From a methodological standpoint, MSMPs extend estimation strategies developed for Markov models, such as maximum likelihood or Bayesian inference, and can accommodate interval-censored or misclassified data [Frydman, 2005]. This makes them a powerful and general tool for multi-state event history analysis.
While both joint models and multi-state models offer powerful tools, they primarily address different facets of disease progression and have complementary strengths. Multi-state models excel at understanding the sequence and timing of discrete events, while joint models are adept at modeling continuous biomarker trajectories and their associations with event outcomes for dynamic prediction. The complexity of real-world diseases often necessitates a more integrated approach that can leverage the strengths of both frameworks. For instance, in prostate cancer, multiple types of relapse may occur successively (a multi-state process), and their risk is influenced by the dynamics of longitudinal biomarkers like PSA.
2.5 Joint Models and Multi-State Processes
Previous research has initiated such integration, with models like the one proposed by Ferrer et al. [Ferrer, Rondeau, Dignam, Pickles, Jacqmin-Gadda, and Proust-Lima, 2016], which combines a linear mixed sub-model for longitudinal data with a multi-state sub-model using shared random effects. This model enables the simultaneous analysis of repeated measurements of a biomarker (the longitudinal process, e.g., PSA levels) and the times of transitions between multiple health states (the multi-state process). It explicitly accounts for the link between these two correlated processes and uses information from the biomarker dynamics to explain or predict clinical progression events. The multi-state sub-model assumes a non-homogeneous Markov process for the clinical progression, meaning that the future of the process depends only on the current state (Markov property) and the time elapsed since study entry (non-homogeneous property).
Although Ferrer et al. [Ferrer et al., 2016] note that their joint modeling framework could be adapted to semi-Markov processes in other contexts, their primary application to prostate cancer progression assumes a non-homogeneous Markov multi-state process. Consequently, the model does not explicitly incorporate the effect of time spent within a particular state on the instantaneous risk of transitioning out of that state, beyond the overall follow-up time from study entry.
While this prior work has managed to combine multi-state and joint models in particular applications, these models typically rely on a linear mixed model for the biomarker and impose a specific directed acyclic graph (DAG) structure through their likelihood decomposition, which restricts the types of dependency structures they can represent. As a result, limitations remain in their generalizability, flexibility, and the scalability of inference. In what follows, we propose a more general framework defined on arbitrary directed graphs (not restricted to acyclic forms) that accommodates both Markov and semi-Markov assumptions and allows nonlinear longitudinal submodels.
3 Extending Multi-State Joint Models
Joint models traditionally focus on a single time-to-event outcome, which limits their ability to represent complex disease trajectories involving multiple intermediate and recurrent events. To address this limitation, we extend the classical nonlinear joint modeling framework to incorporate multi-state processes, enabling the joint analysis of longitudinal biomarkers and transitions between multiple health states. This is achieved by introducing a directed graph structure $G=(V,E)$, which encodes all permissible transitions and supports both Markov and semi-Markov assumptions.
3.1 Graph Structure and Example
Let us consider the illustrative state-transition diagram shown in Figure 1, representing a 4-state process.
Figure 1: Illustration of the 4-state transition graph $G=(V,E)$, including absorbing state 4 (Dead).
This graph can be encoded by an adjacency matrix $A=[A_{kk^{\prime}}]_{kk^{\prime}}$ where $A_{kk^{\prime}}=1$ if a transition from state $k$ to $k^{\prime}$ is possible, and 0 otherwise. For the graph in Figure 1, the adjacency matrix is:
$A=\begin{pmatrix}0&1&0&1\\ 0&0&1&1\\ 0&1&0&1\\ 0&0&0&0\end{pmatrix}$
Under certain assumptions borrowed from the semi-Markov literature, competing and recurrent risks mentioned in the introduction, as well as the classical joint model, can be represented using such graphs as illustrated in Figure 2.
Standard Risk
Competing Risks
Recurrent Risk
Figure 2: Illustration of classical joint modeling approaches: (left) standard risk, (middle) competing risks, and (right) recurrent risk.
Such joint modeling offers a flexible framework that extends beyond classical approaches and is particularly useful in medical applications.
3.2 Individual Trajectories and the Markov Assumption
Each individual $i$ follows a true trajectory:
$\mathcal{T}_{i}^{*}=\left((T_{i0},S_{i0}),(T_{i1},S_{i1}),(T_{i2},S_{i2}),\ldots\right),$
where $T_{il}$ denotes the $l$-th transition time and $S_{il}\in V$ the corresponding state. Between transitions, the state $S_{i}(t)$ is assumed constant.
Only a finite portion of the trajectory is observed, up to a right-censoring time $C_{i}$:
$\big(\mathcal{T}_{i},C_{i}\big)=\Big(\big((T_{i0},S_{i0}),\dots,(T_{im_{i}},S_{im_{i}})\big),C_{i}\Big),$
where $m_{i}(t)=\max\{l:T_{il}\leq t\}$, and $m_{i}=m_{i}(C_{i})$ denotes the last observed transition index.
We assume that the process satisfies the (time-inhomogeneous) Markov property:
$\mathscr{L}\big((T_{i(l+1)},S_{i(l+1)})\mid\{(T_{il^{\prime}},S_{il^{\prime}})\}_{l^{\prime}\leq l}\big)=\mathscr{L}\big((T_{i(l+1)},S_{i(l+1)})\mid(T_{il},S_{il})\big).$ ($\ast$)
Remark 1.
The transition times are not restricted to be non-negative.
3.3 Transition Intensities and Survival Functions
Let $(k,k^{\prime})\in V^{2}$. The instantaneous risk of transition from state $k$ to state $k^{\prime}$ at time $t$, given entry at time $t_{0}$, is defined as:
$\lambda_{i}^{k^{\prime}\mid k}(t\mid t_{0})\coloneqq\lim_{\delta\to 0^{+}}\frac{\mathbb{P}(T_{i(l+1)}\leq t+\delta,\ S_{i(l+1)}=k^{\prime}\mid T_{i(l+1)}>t,\ T_{il}=t_{0},\ S_{il}=k)}{\delta}.$
Summing over all admissible transitions from $k$, we get the total instantaneous risk of leaving state $k$:
$\sum_{s:\,(k,s)\in E}\lambda_{i}^{s\mid k}(t\mid t_{0})=\lim_{\delta\to 0^{+}}\frac{\mathbb{P}(T_{i(l+1)}\leq t+\delta\mid T_{i(l+1)}>t,T_{il}=t_{0},S_{il}=k)}{\delta}.$
The cumulative intensity of the transition from $k$ to $k^{\prime}$ is defined by:
$\Lambda_{i}^{k^{\prime}\mid k}(t\mid t_{0})=\int_{t_{0}}^{t}\lambda_{i}^{k^{\prime}\mid k}(w\mid t_{0})\,dw,$
which leads to the corresponding survival function:
$\mathbb{P}(T_{i(l+1)}>t\mid T_{il},S_{il})=\exp\left(-\sum_{s:\,(S_{il},s)\in E}\Lambda_{i}^{s\mid S_{il}}(t\mid T_{il})\right).$
Conditional transition probabilities are given by:
$\mathbb{P}(S_{i(l+1)}\mid T_{i(l+1)},T_{il},S_{il})=\frac{\lambda_{i}^{S_{i(l+1)}\mid S_{il}}(T_{i(l+1)}\mid T_{il})}{\sum_{s:\,(S_{il},s)\in E}\lambda_{i}^{s\mid S_{il}}(T_{i(l+1)}\mid T_{il})}.$
Combining both, the joint density is:
$p(T_{i(l+1)},S_{i(l+1)}\mid T_{il},S_{il})=\lambda_{i}^{S_{i(l+1)}\mid S_{il}}(T_{i(l+1)}\mid T_{il})\,\exp\left(-\sum_{s:\,(S_{il},s)\in E}\Lambda_{i}^{s\mid S_{il}}(T_{i(l+1)}\mid T_{il})\right).$
Furthermore, conditionning on the event $T_{i(j+1)}\geq t$ for some known $t$ can also easily be achieved by using the Chasles relation:
$\forall u\geq t,$
$\mathbb{P}(T_{i(l+1)}>u\mid T_{il},S_{il},T_{i(l+1)}>t)=\exp\left(-\sum_{s:\,(S_{il},s)\in E}\int_{t}^{u}\lambda_{i}^{s\mid S_{il}}(w\mid T_{il})\,dw\right).$
3.4 Multi-State Joint Model Specification
The proposed multi-state joint model with Gaussian prior and homoscedastic Gaussian noise is specified by:
$\begin{cases}Y_{ij}=h(t_{ij},\psi_{i})+\epsilon_{ij},\\ \psi_{i}=f(\gamma,X_{i},b_{i}),\quad b_{i}\sim\mathcal{N}(0,Q),\quad\epsilon_{ij}\sim\mathcal{N}(0,R),\\ \lambda_{i}^{k^{\prime}\mid k}(t\mid X_{i},T_{im_{i}(t)},\mathcal{H}_{i}(t))=\lambda_{0}^{k^{\prime}\mid k}(t\mid T_{im_{i}(t)})\cdot\exp\big(\alpha^{k^{\prime}\mid k}g^{k^{\prime}\mid k}(t,X_{i},\psi_{i})+\beta^{k^{\prime}\mid k}X_{i}\big),\end{cases}$
along with the assumptions 1, 2, and 5 as well as the following assumptions, analogous to assumptions 3 and 4:
Assumption 6.
Conditionally on $b_{i}$, the trajectories $\mathcal{T}_{i}^{*}$ are independent and satisfy the semi-Markov property $(\ast)$.
Assumption 7.
Conditionally on $b_{i}$, the biomarker process $Y_{ij}$ is independent of the transition process $\mathcal{T}_{i}^{*}$.
3.5 Clock Reset vs. Clock Forward
Two baseline hazard specifications are commonly used:
$\lambda_{0}^{k^{\prime}\mid k}(t\mid T_{im_{i}(t)})=\begin{cases}\lambda_{0}^{k^{\prime}\mid k}(t-T_{im_{i}(t)})&\text{(clock reset)},\\ \lambda_{0}^{k^{\prime}\mid k}(t)&\text{(clock forward)}.\end{cases}$
The clock reset model resets the baseline hazard at each state entry and is appropriate when the transition risk depends on time spent in the state (Asanjarani et al. [2022]). The clock forward model assumes that risk accumulates with global time, ignoring the state entry time.
3.6 Initial State
We assume that the initial state $S_{i0}$ is observed. However, the framework could also incorporate a multinomial model for unobserved initial states (Yiu et al. [2018]).
This flexible framework enables inference on complex event histories, leveraging the full trajectory of biomarkers to refine predictions of transitions and survival [de Wreede et al., 2010].
3.7 Multi-State Simulation
In nonlinear joint models with known design and known parameters parameters $\theta=(\gamma,Q,R,\alpha,\beta)$, the occurrence times of events conditionally on the latent variables $b_{i}$ and the covariates $X_{i}$ can be easily sampled using the inverse cumulative distribution function transform method. This may be achieved through bisection or other root-finding algorithms.
The simulation of the transition process $\mathcal{T}_{i}^{*}$ conditionally on $b_{i}$, $X_{i}$, and $(T_{i0},S_{i0})$ can be achieved drawing on the semi-Markov property 6 by considering one transition at a time, according to Algorithm 1. The algorithme is similar to Gillespie’s algorithm [Gillespie, 2007], and can also include a survival condition $T_{i1}\geq t_{i}^{\text{surv}}$ that uses the Chasles relation LABEL:rel:chasles.
Algorithm 1 Simulation of trajectory $i$
1:subject-specific censoring time $C_{i}\in\mathbb{R}\cup\{+\infty\}$; covariates $X_{i}$; latent variables $b_{i}$; initial time-state pair $(T_{i0},S_{i0})$; optional survival condition $T_{i1}\geq t_{i}^{\text{surv}}$, defaults to $t_{i}^{\text{surv}}=$-$\infty$.
2:
3:Initialize $\ell_{i}\leftarrow 1$, $\mathcal{T}_{i}\leftarrow\left((T_{i0},S_{i0})\right)$.
4:
5:while $T_{i(\ell_{i}-1)}<C_{i}$ and $\{s:(S_{i(\ell_{i}-1)},s)\in E\}\neq\emptyset$ do
6:  for each $s$ with $(S_{i(\ell_{i}-1)},s)\in E$ do
7:   Draw $T_{i\ell_{i}}^{s}$ such that
8:      $\log\mathbb{P}(T_{i\ell_{i}}^{s}>t)=\int_{T_{i(\ell_{i}-1)}\vee t_{i}^{\text{surv}}}^{t}\lambda_{i}^{s\mid S_{i(\ell_{i}-1)}}(w\mid\theta,X_{i},T_{i(\ell_{i}-1)},b_{i})\,dw,$
9:      $\forall t\geq T_{i(\ell_{i}-1)}\vee u.$
10:  end for
11:
12:  Set $T_{i\ell_{i}}\leftarrow\min_{m}T_{i\ell_{i}}^{s}$, and $S_{i\ell_{i}}\leftarrow\arg\min_{m}T_{i\ell_{i}}^{s}$.
13:  Append $(T_{i\ell_{i}},S_{i\ell_{i}})$ to $\mathcal{T}_{i}$.
14:  $\ell_{i}\leftarrow\ell_{i}+1$.
15:end while
16:
17:if $T_{i(\ell_{i}-1)}>C_{i}$ then
18:  Remove the last pair: $\mathcal{T}_{i}\leftarrow\mathcal{T}_{i}[:-1]$.
19:end if
[…]
4 Dynamic Prediction
We now focus on constructing random variables of interest for multi-state prediction for a new individual $i$, such that these random variables may be feasibly simulated with Algorithm 1.
Let $\tau_{i}$ be a stopping time adapted to the filtration $\mathcal{F}_{in}=\sigma(T_{i0},S_{i0},\dots,T_{in},S_{in})$, and let $\kappa_{i}=\{\inf n\in\mathbb{N}:\mathbb{P}(\tau_{i}=+\infty\mid\mathcal{F}_{in})=1\}$ be a second stopping time.
Assume that:
$\mathbb{P}(\min(\tau_{i},\kappa_{i})<+\infty)=1.$
Here, $\tau_{i}$ represents the time at which an event occurs for an individual $i$, while $\kappa_{i}$ is the time at which it becomes certain that the event will not occur. For instance, if the event of interest is patient remission, but the patient dies without remission, then $\tau_{i}=+\infty$, while $\kappa_{i}$ would be finite. As such, it is almost surely possible to simulate the trajectory of an individual $i$ up to the time of the event or the time of censoring.
Let $\xi$ be a function taking values in some space $\mathbb{Y}$:
$\xi:$ $(\mathbb{N}\cup\{+\infty\})^{2}\times(\mathbb{R}\times V)^{\mathbb{N}}\longrightarrow\mathbb{Y}$
$\tau_{i},\kappa_{i},\mathcal{T}_{i}^{*}\longmapsto\xi(\tau_{i},\kappa_{i},\mathcal{T}_{i}^{*})=\xi\left(\tau_{i},\kappa_{i},(T_{il},S_{il})_{l\leq\min(\tau_{i},\kappa_{i})}\right)$
This function therefore only depends on the beginning of the trajectory and not on its entirety. Thus, during simulations, if $\tau_{i}$ is finite or almost surely infinite after a certain index, there is no need to simulate the entire trajectory. Applications include, for instance, the distributions of states at any time, transition times, return times, sojourn times, hitting times and many other quantities. This flexibility is illustrated by Examples 1 and 2.
Example 1 (State at time $t$).
Let $G=(V,E)$ be a directed acyclic graph, $t\in\mathbb{R}$ be a fixed time, and $\mathbb{Y}=V$ be the set of possible states. Let $\tau_{i}=\inf\{n\in\mathbb{N}:T_{in}\geq t\text{ or }S_{in}\text{ is absorbing}\}$. Clearly, $\tau_{i}\leq\operatorname{depth}(G)<+\infty$. Then, take $\xi_{t}(\tau_{i},\kappa_{i},\mathcal{T}_{i}^{*})=S_{i\tau_{i}}$, which corresponds to the state of the individual $i$ at time $t$.
Example 2 (Hitting time).
Let $G=(V,E)$ be a directed acyclic graph, $\mathbb{A}\subset V$ a non-empty subset of states, and $\mathbb{Y}=\mathbb{R}\cup\{+\infty\}$. For any two non-empty sets $\mathbb{A},\,\mathbb{B}\subset V$, we note $\mathbb{A}\rightsquigarrow\mathbb{B}$ if there exists a path from $\mathbb{A}$ to $\mathbb{B}$ in $G$. Let $\tau_{i}=\inf\{n\in\mathbb{N}:S_{in}\in\mathbb{A}\}$, with the convention $\inf\emptyset=+\infty$. Then, $\kappa_{i}=\inf\{n\in\mathbb{N}:\{S_{in}\}\not\rightsquigarrow\mathbb{A}\}$ and $\min(\tau_{i},\kappa_{i})\leq\operatorname{depth}(G)<+\infty$. Finally, $\xi_{\mathbb{A}}(\tau_{i},\kappa_{i},\mathcal{T}_{i}^{*})=\mathds{1}_{\tau_{i}<+\infty}T_{i\tau_{i}}+\mathds{1}_{\tau_{i}=+\infty}\tau_{i}$ represents the hitting time for the set $\mathbb{A}$.
5 Likelihood and Statistical Inference
The estimation of model parameters $\theta=(\gamma,Q,R,\alpha,\beta)$ relies on two likelihood formulations: the complete-data likelihood $\mathcal{L}_{\text{full}}$, which includes latent variables, and the marginal likelihood $\mathcal{L}_{\text{marginal}}$, obtained by integrating them out:
$\mathcal{L}_{\text{marginal}}(\theta;X,Y,\mathcal{T},C)=\int\mathcal{L}_{\text{full}}(\theta;;X,Y,\mathcal{T},C,b)\,db,$
where:
$\mathcal{L}_{\text{full}}(\theta;;X,Y,\mathcal{T},C,b)=\prod_{i=1}^{n}p(b_{i}\mid\theta)p(\mathcal{T}_{i},C_{i}\mid\theta,X_{i},b_{i})\prod_{j=1}^{m_{i}}p(Y_{ij}\mid\theta,X_{i},b_{i}).$
This expression is very similar to that obtained by Rizopoulos [2012] and is subdivided into three terms:
Prior likelihood:
$p(b_{i}\mid\theta)=\frac{(2\pi)^{-q/2}}{\det(Q)^{1/2}}\exp\left(-\tfrac{1}{2}b_{i}^{T}Q^{-1}b_{i}\right),$
Longitudinal likelihood:
$p(Y_{ij}\mid\theta,X_{i},b_{i})=\frac{(2\pi)^{-d/2}}{\det(R)^{1/2}}\exp\left(-\tfrac{1}{2}\left(Y_{ij}-h(t_{ij},\psi_{i})\right)^{T}R^{-1}\left(Y_{ij}-h(t_{ij},\psi_{i})\right)\right),$
$\text{with }\psi_{i}=f(\gamma,X_{i},b_{i}).$
Semi-Markov likelihood:
$p(\mathcal{T}_{i},C_{i}\mid\theta,X_{i},b_{i})={}$ $\prod_{l=0}^{m_{i}-1}p\left((T_{i(l+1)},S_{i(l+1)})\mid\theta,X_{i},(T_{il},S_{il}),b_{i}\right)$ (1)
$\exp\left(-\sum_{s:\,(S_{im_{i}},s)\in E}\Lambda_{i}^{k^{\prime}\mid S_{im_{i}}}(C_{i}\mid\theta,X_{i},T_{im_{i}},b_{i})\right).$
[…]
It is to be noted that this integral is typically intractable in closed form due to the nonlinear nature of the model, necessitating numerical approximation methods such as Monte Carlo integration or Laplace approximation.
5.1 Stochastic Gradient Estimation via Fisher Identity
Several optimization methods adapted from the nonlinear mixed-effects literature can be applied, including Stochastic EM [Kuhn and Lavielle, 2004], Laplace approximation [Wolfinger, 1993], Gauss-Hermite quadrature, stochastic gradient ascent with Robbins–Monro updates [Robbins and Monro, 1951], and MCMC-based approximations such as Metropolis–Hastings and Hamiltonian Monte Carlo. These strategies are implemented in software such as JMBayes [Rizopoulos et al., 2024, Rizopoulos, 2020]. Another approach requiring mild regularity assumptions on the log marginal likelihood, but without the need for the model to belong in the exponential family, is to consider a stochastic gradient ascent scheme [Caillebotte et al., 2025, Baey et al., 2023] using Fisher’s identity and following the Robbins-Monro procedure [Robbins and Monro, 1951].
Under interchangeability of integral and differentiation, setting $x=(X,Y,\mathcal{T},C)$ for convenience, the Fisher identity writes:
$\nabla_{\theta}\log\mathcal{L}_{\text{marginal}}(\theta;x)=\mathbb{E}_{b\sim p(\cdot\mid x,\theta)}\left(\nabla_{\theta}\log\mathcal{L}_{\text{full}}(\theta;x,b)\right).$
This expectation is approximated using Monte Carlo samples from the posterior $p(\cdot\mid x,\theta)$, avoiding the need to evaluate the intractable marginal likelihood $\mathcal{L}_{\text{marginal}}(\theta;x)$. This formulation naturally supports stochastic gradient ascent with MCMC-based posterior sampling and is amenable to minibatch parallelization across individuals, ensuring the convergence to a critical point. The update rule is as follows:
Algorithm 2 Stochastic gradient ascent model inference
1:data $x$; initial parameter $\theta_{0}$; step sizes $(\eta_{n})_{n\geq 0}$ with $\sum_{n}\eta_{n}=+\infty$, $\sum_{n}\eta_{n}^{2}<+\infty$; batch size $K$; sampler $p_{n}^{\mathrm{MCMC}}$; stopping criterion.
2:
3:while not converged do
4:  Draw $b_{k}\sim p(\cdot\mid x,\theta_{n})\approx p_{n}^{\text{MCMC}}$.
5:  $\theta_{n+1}\leftarrow\theta_{n}+\frac{\eta_{n}}{K}\sum_{k=1}^{K}\nabla_{\theta}\log\mathcal{L}(\theta_{n};x,b_{k})$.
6:end while
Preconditioning matrices may also be used, such as the Fisher Information Matrix in a natural gradient descent framework, or even specialized optimizers such as Adam [Kingma and Ba, 2014], NAdam [Dozat, 2016] or Adagrad [Duchi et al., 2011]. In particular, the Adam optimizer will be later used in our numerical experiments.
6 Implementation Details
The nonlinear joint model is fully implemented in Python, with PyTorch and NumPy as its only major dependencies. Notably, the stochastic gradient scheme is carried out following Algorithm 2 and heavily relies on PyTorch’s automatic differentiation and broadcasting capabilities [Paszke et al., 2017].
6.0.0.1 Posterior Sampling and Parameter Estimation.
A Metropolis-Hastings sampler is used to simultaneously draw posterior samples of the random effects in an efficient way. Adaptive step sizes are implemented at the individual level via the Robbins-Monro procedure. We favored Metropolis-Hastings over alternatives such as Hamiltonian Monte Carlo or Metropolis-within-Gibbs for its simplicity. Autocorrelation can be monitored and mitigated by subsampling. The acceptance rate is targeted at $0.234$, which is asymptotically optimal for Gaussian proposals in high dimensions [Gelman et al., 1997]. Although our posteriors are not Gaussian, this rule is simple and effective. Moreover, as the number of longitudinal measurements increases ($m_{i}\to+\infty$), posterior distributions converge to a multivariate normal (Rizopoulos et al. [2008]).
6.0.0.2 Numerical integration.
Cumulative hazards are computed using Gauss–Legendre quadrature [Lether, 1978], which integrates exactly all polynomials up to degree 2n12n-1 with $n$ nodes:
$\int_{a}^{b}f(t)\,dt\approx\sum_{i=1}^{n}f(t_{i})w_{i}.$
We set $n=32$ by default. Since nodes are fixed, integration points are cached, using a hash table, outside the backpropagation graph, reducing cost and memory. The positivity of the quadrature weights ensures stable computation and preserves the positivity of integrals.
6.0.0.3 Simulation of event times.
Event times are generated by inverting the survival function via a bisection search. An exponential random variable is first drawn to determine the event threshold. Starting from the initial and censoring times, the cumulative hazard is iteratively approximated with Gauss–Legendre quadrature until the threshold is reached. This procedure naturally incorporates censoring and yields the simulated event time.
6.0.0.4 Covariance matrix parameterization.
[…]
7 Simulation Study
To study the robustness of the estimator, we propose a simulation example and compute an estimate of the Root Mean Square Error (RMSE) for each parameter, based on different datasets generated from a common underlying distribution.
In parallel, we illustrate the usage of the jmstate package through concise code snippets, structured as follows:
  • Specification of the true underlying model.
  • Simulation of longitudinal and event data.
  • Stochastic optimization of the parameters with corresponding diagnostic plots.
7.1 Model Specification
The model considered involves three states (See Figure 3), labeled 0 (healthy), $1$ (sick), and $2$ (terminal stage). The allowed transitions are $0\rightarrow 1$, $0\rightarrow 2$, and $1\rightarrow 2$, corresponding to a biological modeling of the biomarker evolution, where a too rapid decrease of the biomarker or a too low value leads to the degradation of the patient’s health status.
Figure 3: State transition diagram of the multi-state model: transitions are allowed from Healthy (0) to Sick (11), from Healthy (0) to Terminal Stage (22), and from Sick (11) to Terminal Stage (22).
The regression function is a piecewise affine approximation of a bi-exponential curve, with an inflexion point at a fixed known time $\tau=6$ which may correspond to a drug administration time, causing a rupture in the longitudinal behavior while remaining continuous:
$h(t,\psi)=\psi_{1}+\psi_{2}t+\mathds{1}_{t>\tau}(\psi_{3}-\psi_{2})(t-\tau),$
which can be written in Python, after necessary imports, as:
import torch TAU = 6.0 def reg(t: torch.Tensor, psi: torch.Tensor): b, w1, w2 = psi.chunk(3, dim=-1) # Extract relevant terms # psi has shape (n_chains, n_individuals, n_repetitions) # output has shape (n_chains, n_individuals, n_repetitions, 1) return (b + w1 * t + (w2 - w1) * (t > TAU) * (t - TAU)).unsqueeze(-1)
Note this python function must accept tensors psi with two or three dimensions and return a tensor of dimension $\operatorname{dim}(\text{psi})+1$, with a shape whose last dimension matches the dimension of the biomarker (here, $\mathbb{R}^{1}$ so it is equal to $1$).
For the link function, the concatenation of $h$ as well as its partial derivative with respect to time is considered and shared across all possible transitions: $g=\left(h,\frac{\partial h}{\partial t}\right)$, which is efficiently implemented:
def link(t: torch.Tensor, psi: torch.Tensor): b, w1, w2 = psi.chunk(3, dim=-1) diff = (w2 - w1) * (t > TAU) val = b + w1 * t + diff * (t - TAU) der = w1 + diff return torch.cat([val.unsqueeze(-1), der.unsqueeze(-1)], dim=-1)
Additionally, normally distributed covariates $X$ and exponential baseline hazards were considered with a clock reset specification. The model design is therefore declared as follows:
from jmstate.functions import Exponential, gamma_plus_b from jmstate.typedefs import ModelDesign # Survival model specification surv = { (0, 1): (Exponential(0.1), link), (0, 2): (Exponential(0.01), link), (1, 2): (Exponential(0.2), link), } # Model design gathers regression, link and hazard functions model_design = ModelDesign(gamma_plus_b, reg, surv)
Furthermore, longitudinal measurements were taken at $m_{i}\leq 20$ time points, and censoring times were drawn from $C_{i}\sim\mathcal{U}\left([10,15]\right)$. Longitudinal observations were also truncated at the censoring times.
The model is therefore parameterized in $\mathbb{R}^{16}$ and the true parameters may be defined as follows:
from jmstate.typedefs import ModelParams from jmstate.utils import repr_from_cov # Gaussian means gamma = torch.tensor([2.5, -1.3, 0.2]) # Covariance matrices Q = torch.diag(torch.tensor([0.6, 0.2, 0.3])) R = torch.tensor([[1.7]]) # Link parameters alphas = { (0, 1): torch.tensor([-0.5, -3.0]), (0, 2): torch.tensor([-1.0, -5.0]), (1, 2): torch.tensor([0.0, -1.2]), } # Covariate parameters betas = { (0, 1): torch.tensor([-1.3]), (0, 2): torch.tensor([-0.9]), (1, 2): torch.tensor([-0.7]), } # Instance declaration real_params = ModelParams( gamma, repr_from_cov(Q, method="diag"), repr_from_cov(R, method="ball"), alphas, betas, )
Note the argument method, which can take values "full", "diag", or "ball". It represents the chosen parameterisation of the covariance matrices, either unconstrained (full), diagonal (diag) or scalar (ball).
7.2 Data Generation
With the parameter values defined above, as well as the choice of the regression and link functions, we are now able to declare the true base model and use it to simulate longitudinal processes and transition trajectories, with the help of the following helper function:
from jmstate import MultiStateJointModel from jmstate.typedefs import ModelData, SampleData from torch.distributions import MultivariateNormal # Declares the true underlying model real_model = MultiStateJointModel(model_design, real_params) def gen_data(n: int, m: int): # Censoring times c = torch.rand(n, 1) * 5 + 10 # Covariates x = torch.randn(n, 1) # Latent and noise distributions Q_dist = MultivariateNormal(torch.zeros(Q.size(0)), Q) R_dist = MultivariateNormal(torch.zeros(R.size(0)), R) # Individual effects b = Q_dist.sample((n,)) psi = model_design.individual_effects_fn(gamma, x, b) # Generates random evaluations points with a minimum distance a = torch.zeros((n, 1)) b = torch.full((n, 1), 15) t = random_far_apart(n, m, a, b, 0.7 * b / m) # Defines initial state for individuals trajectories_init = [[(0.0, 0)]] * n # Samples trajectories sample_data = SampleData(x, trajectories_init, psi) trajectories = real_model.sample_trajectories(sample_data, c) # Samples longitudinal values y = model_design.regression_fn(t, psi) y += R_dist.sample(y.shape[:2]) # Censors longitudinal measurements exceeding censoring times y[t > c] = torch.nan return x, t, y, trajectories, c # Generate data data = ModelData(*gen_data(1000, 20))
A short summary of the longitudinal process as well as the trajectories is given by the Figure 4 below. The function build_buckets was used to determine the number of transitions observed in the dataset. Given a list of trajectories, it returns a dictionary of transitions as keys and Buckets(NamedTuple) as values, containing the individual indices and times of the transitions.
From $\rightarrow$ To # Observations
$0\rightarrow 1$ 613
$0\rightarrow 2$ 375
$1\rightarrow 2$ 592
Figure 4: Simulated data: on the left, a sample of longitudinal measurements from $50$ individuals; on the right, observed transitions between states for the full population of $1000$ individuals.
7.3 Estimator Convergence
To illustrate the convergence of the estimator, we examine the optimization process for a particular run with $n=1000$ (see Figure 5), as well as RMSE values computed from $n_{\text{runs}}=100$ independent runs on different simulated datasets, all generated from the same underlying distribution. The results of these simulations are shown in Table 1.
To fit the model, we first have to define initial parameters, which can be particularly important for the optimization process if the model is not well-behaved. In the present case, without prior knowledge of the true parameters, we zero the values and use the identity matrix for both covariance matrices:
# Declares initial parameters; zero mean and unit variance init_params = ModelParams( torch.zeros_like(gamma), repr_from_cov(torch.eye(Q.size(0)), method="diag"), repr_from_cov(torch.eye(R.size(0)), method="ball"), {k: torch.zeros_like(v) for k, v in alphas.items()}, {k: torch.zeros_like(v) for k, v in betas.items()}, )
To run the optimization process, we can make use of the Fit class that provides a simple interface for choosing the optimizer (torch.optim.Adam by default) and set its hyperparameters through the kwargs argument. Additionally, a stochastic stopping condition based on the first and second moments of the parameters’ consecutive differences can be easily employed with the StopParams class.
Formally:
$m_{i}^{(t)}\leftarrow\beta_{i}m_{i}^{(t-1)}+(1-\beta_{i})(\theta^{(t)}-\theta^{(t-1)})^{i},\quad\hat{m}_{i}^{(t)}\coloneqq\frac{m_{i}^{(t)}}{1-\beta_{i}^{t}},$
and the optimization process is terminated when:
$|\hat{m}_{1}^{(t)}|\leq 10^{-6}+10^{-1}\sqrt{\hat{m}_{2}^{(t)}}.$
At last, the LogParamsHistory class can be used to log the parameters’ history, allowing to monitor the convergence of the parameters.
This is all done automatically by executing the do method of the model which runs the Markov Chain Monte Carlo (MCMC) algorithm to sample from the posterior distribution of the parameters based on the inputed data, a list of Job factories (such as fitting, logging, stopping, or even prediction once the model is fitted). It optionnaly returns metrics collected by the Job factories. The hyperparameters such as the target acceptance rate, the number of chains, the warmup period or the subsample parameter are automatically set based on the job_factories provided but can be overriden:
from jmstate.jobs import Fit, LogParamsHistory, ParamStop # Declares initial model model = MultiStateJointModel(model_design, init_params) # Runs optimization process metrics = model.do( # Metrics object data, job_factories=[ Fit(lr=0.5, fused=True), ParamStop(rtol=0.1), LogParamsHistory(), # Returns a list of ModelParams ], max_iterations=500, )
Although not demonstrated here, the model also supports gradient clipping, scheduling, stacking multiple other stopping criteria based on likelihood or gradient magnitude, penalized likelihood, as well as covariate selection. For more details, please refer to the official documentation of the jmstate package.
The following figure shows the evolution of the parameters during the optimization process and Table 1 summarizes the accuracy of the inferred parameters.
(a) $\gamma$
(b) $Q$
(c) $R$
(d) $\alpha^{1\mid 0}$
(e) $\alpha^{2\mid 0}$
(f) $\alpha^{2\mid 1}$
(g) $\beta^{1\mid 0}$
(h) $\beta^{2\mid 0}$
(i) $\beta^{2\mid 1}$
Figure 5: Evolution of the parameters during the optimization of the marginal log-likelihood using stochastic gradient descent. The dotted lines correspond to the true values.
Parameter True value Mean Standard error RMSE
$\gamma_{1}$ 2.5000 2.4960 0.0370 0.0372
$\gamma_{2}$ -1.3000 -1.3039 0.0257 0.0260
$\gamma_{3}$ 0.2000 0.1936 0.0267 0.0275
$\tilde{Q}_{1}$ 0.2554 0.2243 0.0442 0.0540
$\tilde{Q}_{2}$ 0.8047 0.8009 0.0260 0.0263
$\tilde{Q}_{3}$ 0.6020 0.6001 0.0276 0.0277
$\tilde{R} -$0.2653 -0.2684 0.0083 0.0089
$\alpha_{1}^{1\mid 0}$ -0.5000 -0.49997 0.0486 0.0486
$\alpha_{2}^{1\mid 0}$ -3.0000 -2.9962 0.0800 0.0801
$\alpha_{1}^{2\mid 0}$ -1.0000 -0.9990 0.0847 0.0847
$\alpha_{2}^{2\mid 0}$ -5.0000 -4.9897 0.1191 0.1196
$\alpha_{1}^{2\mid 1}$ 0.0000 0.0011 0.0327 0.0327
$\alpha_{2}^{2\mid 1}$ -1.2000 -1.2045 0.0429 0.0432
$\beta^{1\mid 0}$ -1.3000 -1.3015 0.0501 0.0501
$\beta^{2\mid 0}$ -0.9000 -0.8982 0.0670 0.0670
$\beta^{2\mid 1}$ -0.7000 -0.6981 0.0551 0.0551
Table 1: Comparison of true and inferred parameters (up to 4 decimal places). RMSE is computed using the bias-variance identity.
The algorithm demonstrates good adaptability and robustness, with a notably low bias. The fitting is also very quick to perform, in about $10$ to $15$ seconds per run on an AMD Ryzen 5 5600 processor, despite the use of mini-batch optimization and the higher complexity of the model induced by the nonlinearity of its functions as well as the presence of numerous possible transitions. It is also to be noted that a minimum number of observations for each transition is required to ensure the stability of the optimization process.
Finally, the convergence patterns shown in Figure 5 are remarkably smooth, a result due to the parallelization of multiple chains and the use of mini-batch optimization.
8 Application to the PAQUID Cohort
8.1 Introduction
The data used in this section originates from the PAQUID cohort [Letenneur et al., 1994], a large prospective population-based study initiated in southwestern France in 1988, aimed at understanding the determinants and trajectories of aging. A subsample of $500$ individuals was followed over a period of up to $20$ years [Proust-Lima et al., 2017], with repeated measures of cognitive and physical health, as well as socio-demographic characteristics. In particular, global cognitive functioning was assessed through the Mini-Mental State Examination (MMSE, see Figure 6), while physical dependency was evaluated using the HIER scale, which classifies subjects into four ordered states (see Figure 7): 0 (no dependency), $1$ (mild dependency), $2$ (moderate dependency), and $3$ (severe dependency). In this work, we focus on the association between cognitive decline and the progression of physical dependency by jointly modeling the longitudinal trajectory of MMSE scores and the transitions between the four HIER states using a joint multi-state model. More specifically, the state of an individual $i$ at a given time $t$ is defined as the highest HIER dependency score between trial entry and $t$. This approach allows us to characterize the dynamic interrelationship between cognitive and physical aging processes within a unified statistical framework.
8.2 Model Specification
Since the inference of the transition specific parameters requires one to have observed those said transitions, the state graph was determined based on the dataset, using the build_buckets function. Moreover, both time and longitudinal values were normalized in [1,1][-1,1]. This technique ensured stable and fast convergence of the optimization process by keeping all parameters on roughly the same scale.
From $\rightarrow$ To # Observations
$0\rightarrow 1$ 85
$1\rightarrow 2$ 135
$2\rightarrow 3$ 86
$0\rightarrow 2$ 44
$1\rightarrow 3$ 15
$0\rightarrow 3$ 11
Figure 6: On the left, each light gray line represents one of the $500$ observed sequences of (normalized) MMSE scores, and three functional clusters are also represented; on the right, observed transitions between states.
Therefore, the chosen transition graph is as follows, and only allows monotonic transitions from a lower level of physical dependency to a higher one.
Figure 7: Schematic representation of the four ordered HIER states, with possible transitions. HIER quantifies physical dependency (0: no dependency to 33: severe dependency) [Proust-Lima et al., 2017].
We consider an exponential baseline hazard model for transitions, with a clock reset specification. The parameters of these exponential baseline hazards are jointly optimized with the model parameters by passing a list of base_hazard.parameters() to the extra argument of the ModelParams constructor. Moreover, the betas parameter group uses a single shared parameter for all transitions, imposing a particular substructure on the model, where $\forall(k,k^{\prime})\in E,\,\beta^{k^{\prime}\mid k}=\beta$. This is achieved by passing the same tensor for every transition:
# Will be cloned before fitting sharing same memory address shared_beta = torch.zeros(2) init_betas = {key: shared_beta for key in buckets}
In the light of Figure 6, the regression and link functions were taken to be scaled and shifted hyperbolic tangents such that the normalized MMSE score always has a value of $1$ when $t\to$-$\infty$ and is not increasing:
$\begin{cases}h(t,\psi)=g(t,\psi)=\psi_{1}\tanh\left(\frac{\psi_{3}-t}{\psi_{2}}\right)+(1-\psi_{1})\\ \psi=\begin{pmatrix}\sigma(\gamma_{1}+b_{1})\\ \exp(\gamma_{2}+b_{2})\\ \gamma_{3}+b_{3}\end{pmatrix},\quad\sigma(z)=\frac{1}{1+e^{-z}}\end{cases}$
The function $f$ is therefore implemented as:
def f( gamma: torch.Tensor | None, x: torch.Tensor | None, b: torch.Tensor, ): psi = cast(torch.Tensor, gamma) + b psi[..., 0] = psi[..., 0].sigmoid() psi[..., 1] = psi[..., 1].exp() return psi
Please note that type hints are not mandatory, but help with code completion and static type checking; this is the right signature as covariates and populational parameters are passed as optional arguments that may be None.
At last, an individual covariates correspond to a pair of binary variables indicating wheter or not the individual has a diploma and is a male $\left(\mathds{1}_{\text{CEP}}(i),\mathds{1}_{\text{male}}(i)\right)\in\{0,1\}^{2}$.
8.3 Results
8.3.1 Inference
Fitting was performed using $80\%$ of the data, with the remaining $20\%$ used for testing. With no a priori for the covariance matrices, the full option was selected. The model was fitted using the Fit class with a learning rate of $0.5$ and the Adam optimizer. The optimization process was run for until convergence based on the same criterion as the simulation study with a lower tolerance of $5\%$, which was obtained after a little less than $500$ iterations with $10$ parallel chains.
Before displaying the fitted parameters, the Fisher Information Matrix was computed with the ComputeFIM Job factory implementing the method described in Baey et al. [2023], directly taking into account shared parameters:
from jmstate.jobs import ComputeFIM # Returns the metrics for the FIM, and stores them internally model.do( # Uses fitting data if no data is provided job_factories=ComputeFIM(), # No need to wrap in a list max_iterations=100, # Number of samples is n_chains * 100 )
The standard errors were then computed as the square root of the diagonal of the inverse of the Fisher Information Matrix with the model.stderror property. The results are summarized in Table 2.
Parameter Inferred value Standard error Parameter Inferred value Standard error
$\gamma_{1}$ 2.6898 0.0561 $\gamma_{2}$ 0.2626 0.0137
$\gamma_{3}$ 1.3485 0.0025 $\tilde{Q}_{1}$ -0.5428 0.0429
$\tilde{Q}_{2}$ 0.8452 0.1068 $\tilde{Q}_{3}$ 0.5970 0.0448
$\tilde{Q}_{4}$ 0.0041 0.1309 $\tilde{Q}_{5}$ -2.0407 0.1149
$\tilde{Q}_{6}$ 0.2357 0.0458 $\tilde{R}$ 2.1921 0.0135
$\alpha^{3\mid 1}$ -4.1011 1.1296 $\alpha^{1\mid 0}$ 0.5156 2.5595
$\alpha^{2\mid 1}$ -1.5068 1.6333 $\alpha^{2\mid 0}$ -2.9875 2.3893
$\alpha^{3\mid 2}$ -2.7624 0.9352 $\alpha^{3\mid 0}$ -6.2268 4.0880
$\beta_{1}$ -0.0267 0.5393 $\beta_{2}$ 0.1223 0.6959
Table 2: Inferred parameters and their estimated standard deviations (up to 4 decimal places).
Notably, the estimated standard errors are consistent with the number of observations per transition, with larger standard errors for transitions with fewer observations. Therefore, parameters with too few observations should be interpreted with caution, if at all. The linear covariate parameters also exhibit large standard errors, which indicates the coefficients are not significative in this setting.
The association parameters $\alpha$ being negative, we can infer that cognitive decline is associated with an increase in physical dependency.
8.3.2 Dynamic Prediction
Using the $n_{\text{test}}=100$ individuals not included in the model fitting, we performed dynamic prediction with the PredictTrajectories Job. For each individual, longitudinal measurements and trajectories were truncated at various times $t$, and predictions were made for future time points $u\geq t$ beyond the truncation. Accuracy was then assessed by comparing the most likely predicted state with the true state at each corresponding future time point $u$, as illustrated in Figure 8.
Formally, the prediction function $\xi_{u}$ is defined by Example 1, and the accuracy measure at time $t$ for predicted states $(s_{i})_{i}$ is defined as:
$\text{accuracy}(u)\coloneqq\frac{1}{n_{\text{test}}}\sum_{i=1}^{n_{\text{test}}}\mathds{1}_{\xi_{u\wedge C_{i}}(\tau_{i},\kappa_{i},\mathcal{T}_{i}^{*})=s_{i}}.$
Although $\mathcal{T}_{i}^{*}$ is unkown, the quantity $\xi_{u\wedge C_{i}}(\tau_{i},\kappa_{i},\mathcal{T}_{i}^{*})$ may be computed based on the censored trajectories. However, incorporating censoring times into prediction functions relies on information that is, in principle, not available at prediction time. Nevertheless, this inclusion is necessary to ensure fairness in accuracy metrics, since individuals who are not censored are generally more likely to have progressed to more advanced states of physical dependency than those who are censored.
Figure 8: Accuracy of the predicted state at each time point with respect to the true state for different truncation times.
Here, the initial drops correspond to the truncation times, before which the accuracy is always $100\%$.
As expected from dynamic predictions, the accuracy increases as more data is available, therefore when the truncation time grows. We also observe that the longer the prediction horizon, the lower the accuracy of the predictions.
9 Conclusion
We presented jmstate, a Python package for joint modeling of nonlinear longitudinal biomarkers and multi-state survival processes. The framework unifies both components on arbitrary directed graphs, extending classical models to Markov and semi-Markov settings and supporting nonlinear submodels. Inference is implemented with scalable stochastic gradient methods.
Simulation studies confirmed robust parameter recovery, and an application to the PAQUID cohort highlighted the ability to capture the interplay between cognitive decline and physical dependency.
Since the model may potentially involve many transitions and therefore a large number of parameters, a strategy of parameter sharing across transitions could help reduce the overall number of parameters. This is especially beneficial in settings with limited data for certain transitions, or when biological or clinical knowledge suggests similar effects across multiple transitions. Parameter sharing improves statistical efficiency, reduces overfitting, and facilitates model interpretability.
Future directions include a spline-based parametrization of baseline hazards to allow for nonparametric estimation. This approach is particularly useful when the true baseline hazard is expected to be complex or non-monotonic, or when parametric forms (such as exponential or Weibull) may be too restrictive and lead to model misspecification.
Another promising direction is the development of efficient visualization tools for multi-state trajectories and dynamic predictions. Interactive representations of individual event histories, transition probabilities, and longitudinal biomarker evolution would greatly facilitate model interpretation and communication of results to applied researchers.
Funding
This work was partially funded by the Stat4Plant project ANR-20-CE45-0012.
Acknowledgments
The authors would like to thank Jean-Benoist Leger for helpful discussions.
[…]
Proof.
First consider the case $t_{i}^{\text{surv}}=$-$\infty$. Let $t>T_{i\ell_{i}}$:
$\mathbb{P}(T_{i(\ell_{i}+1)}>t\mid T_{i\ell_{i}},S_{i\ell_{i}})$ $=\mathbb{P}(\bigcap_{s:\,(E_{i(\ell_{i}-1)},s)\in E}T_{i(\ell_{i}+1)}^{s}>t\mid T_{i\ell_{i}},S_{i\ell_{i}}),$
$=\prod_{s:\,(S_{i\ell_{i}},s)\in E}\mathbb{P}(T_{i(\ell_{i}+1)}^{s}>t\mid T_{i\ell_{i}},S_{i\ell_{i}}),$
$=\exp\Big(-\sum_{s:\,(S_{i\ell_{i}},s)\in E}\Lambda_{i}^{s\mid S_{i(\ell_{i}-1)}}\big(t\mid\theta,X_{i},T_{i\ell_{i}},b_{i})\Big).$
Furthermore:
$\mathbb{P}(S_{i(\ell_{i}+1)}\mid T_{i(\ell_{i}+1)}=t,T_{i\ell_{i}},S_{i\ell_{i}})$
$\propto p(\{T_{i(\ell_{i}+1)}^{S_{i(\ell_{i}+1)}}=t\}\cap\bigcap_{\begin{subarray}{c}s:\,(S_{i\ell_{i}},s)\in E,\\ s\neq S_{i(\ell_{i}+1)}\end{subarray}}\{T_{i(\ell_{i}+1)}^{s}\geq t\}\mid T_{i\ell_{i}},S_{i\ell_{i}}),$
$=\lambda_{i}^{S_{i(\ell_{i}+1)}\mid S_{i\ell_{i}}}\big(t\mid\theta,X_{i},T_{i\ell_{i}},b_{i})\exp\Big(-\sum_{s:\,(S_{i\ell_{i}},s)\in E}\Lambda_{i}^{s\mid S_{i\ell_{i}}}\big(t\mid\theta,X_{i},T_{i\ell_{i}},b_{i})\Big).$
Combining both, we retrieve the same joint density as in formula 3.3.
If $t_{i}^{\text{surv}}>$-$\infty$, using the Chasles relation LABEL:rel:chasles, one can also check that both densities are equal. Note that the condtioning only affects the first transition, as the next time $\forall\ell_{i}>0,\,T_{i\ell_{i}}\geq t_{i}^{\text{surv}}\implies\forall\ell_{i}>0,\,T_{i\ell_{i}}\vee t_{i}^{\text{surv}}=T_{i\ell_{i}}$.
[…]
Proof.
The proof is short and relies on the fact that, using the Semi-Markovian assumption 6:
$p\left((T_{i(l+1)},S_{i(l+1)})_{0\leq l\leq m_{i}-1}\mid\theta,X_{i},b_{i}\right)$ $=\prod_{l=0}^{m_{i}-1}p\left((T_{i(l+1)},S_{i(l+1)})\mid\theta,X_{i},(T_{il^{\prime}},S_{il^{\prime}})_{l^{\prime}\leq l},b_{i}\right),$
$=\prod_{l=0}^{m_{i}-1}p\left((T_{i(l+1)},S_{i(l+1)})\mid\theta,X_{i},(T_{il},S_{il}),b_{i}\right).$
On the other hand, the probability that no additional event is observed between $T_{im_{i}}$ and $C_{i}$ is $1$ if the last state reached is absorbing, and otherwise:
$\mathbb{P}(T_{i(m_{i}+1)}>C_{i}\mid\theta,X_{i},(T_{im_{i}},S_{im_{i}}),b_{i})=\exp\left(-\sum_{s:\,(S_{im_{i}},s)\in E}\Lambda_{i}^{s\mid S_{im_{i}}}(C_{i}\mid\theta,X_{i},T_{im_{i}},b_{i})\right),$
and so we recover the formula cited above with the convention that an empty sum is 0.

### v2

@@
-jmstate, a Flexible Python Package for Multi-State Joint Modeling
+A General Framework for Joint Multi-State Models
 Abstract
-Classical joint modeling approaches often rely on competing risks or recurrent event formulations to account for complex real-world processes involving evolving longitudinal markers and discrete event occurrences.
-However, these frameworks typically capture only limited aspects of the underlying event dynamics.
-Multi-state joint models offer a more flexible alternative by representing full event histories through a network of possible transitions, including recurrent cycles and terminal absorptions, all potentially influenced by longitudinal covariates.
-In this paper, we propose a general framework that unifies longitudinal biomarker modeling with multi-state event processes defined on arbitrary directed graphs.
-Our approach accommodates both Markovian and semi-Markovian transition structures, and extends classical joint models by coupling nonlinear mixed-effects longitudinal submodels with multi-state survival processes via shared latent structures.
-We derive the full likelihood and develop scalable inference procedures based on stochastic gradient descent.
-Furthermore, we introduce a dynamic prediction framework, enabling individualized risk assessments along complex state-transition trajectories.
-To facilitate reproducibility and dissemination, we provide an open-source Python library jmstate implementing the proposed methodology, available on PyPI.
-Simulation experiments and a biomedical case study demonstrate the flexibility and performance of the framework in representing complex longitudinal and multi-state event dynamics.
-The full Python notebooks used to reproduce the experiments as well as the source code of this paper are available on GitLab.
-Keywords joint modeling • multi-state processes • longitudinal data • survival analysis • stochastic gradient descent • dynamic prediction • Python library
+Classical joint modeling approaches often rely on competing risks or recurrent event formulations to describe complex processes involving evolving longitudinal biomarkers and discrete event occurrences, but these frameworks typically capture only limited aspects of the underlying event dynamics.
+We propose a general multi-state joint modeling framework that unifies longitudinal biomarker dynamics with multi-state time-to-event processes defined on arbitrary directed graphs.
+The proposed framework accommodates arbitrary directed transition graphs, nonlinear longitudinal submodels, and scalable inference via stochastic gradient descent.
+This formulation encompasses both Markovian and semi-Markovian transition structures, allowing recurrent cycles and terminal absorptions to be naturally represented.
+The longitudinal and event processes are linked through shared latent structures within nonlinear mixed-effects models, extending classical joint modeling formulations.
+We derive the complete likelihood, establish conditions for identifiability, and develop scalable inference procedures based on stochastic gradient descent to enable high-dimensional and large-scale applications.
+In addition, we formulate a dynamic prediction framework that provides individualized state-transition probabilities and personalized risk assessments along complex event trajectories.
+Through simulation and application to the PAQUID cohort, we demonstrate accurate parameter recovery and individualized prediction.
+keywords:
+joint modeling, multi-state processes, longitudinal data, survival analysis, stochastic gradient descent, dynamic prediction
 1 Introduction
-Joint modeling of longitudinal and time-to-event data has become an essential tool of modern biostatistics [Papageorgiou et al., 2019], particularly for dynamic prediction in clinical applications.
-Classical joint models typically couple a longitudinal biomarker process with a single time-to-event outcome [Rizopoulos, 2012], allowing for the integration of biological knowledge and individual heterogeneity via shared latent structures.
-However, many real-world processes involve multiple possible outcomes, intermediate stages, or recurrent events, which cannot be fully captured by a single-event framework [Król et al., 2017].
-In such settings, multi-state models provide a more flexible approach [Ferrer et al., 2016].
-Multi-state models represent an overall joint process of clinical course as a series of discrete stages or health states that occur sequentially [Lovblom et al., 2024].
+Joint modeling of longitudinal and time-to-event data has become an essential tool of modern biostatistics [papageorgiou2019overview], particularly for dynamic prediction in clinical applications.
+Classical joint models typically couple a longitudinal biomarker process with a single time-to-event outcome [rizopoulos2012joint], allowing for the integration of biological knowledge and individual heterogeneity via shared latent structures.
+However, many real-world processes involve multiple possible outcomes, intermediate stages, or recurrent events, which cannot be fully captured by a single-event framework [KrolMauguen2017].
+In such settings, multi-state models provide a more flexible approach [ferrer2016joint].
+Multi-state models represent an overall joint process of clinical course as a series of discrete stages or health states that occur sequentially [lovblom2024modeling].
 In biostatistics, these models are widely used for survival and reliability analysis, allowing for a richer and more accurate representation by capturing alternative paths to an event of interest, intermediate events, and progressive disease.
 Key components of multi-state models include transition intensity functions, which denote the instantaneous risk of moving from one state to another, and transition probability functions, which describe the probability of transition over longer intervals.
-Often, these models assume the Markov property, where future transitions depend only on the current state, simplifying the transition intensity functions [Asanjarani et al., 2022].
-However, semi-Markov processes extend this by allowing future probability transitions to depend on the sojourn time, with the clock resetting after each transition.
+Often, these models assume the Markov property, where future transitions depend only on the current state, simplifying the transition intensity functions [asanjarani2022estimation].
+The Markov assumption can be relaxed by adopting a semi-Markov formulation, in which the transition probabilities depend not only on the current state but also on the sojourn time, effectively resetting the time scale after each transition.
 The link between multi-state models and joint models arises when a multi-state process is integrated as a component within a broader joint modeling framework .
@@
 The model proposed by Ferrer et al.
-[Ferrer, Rondeau, Dignam, Pickles, Jacqmin-Gadda, and Proust-Lima, 2016] exemplifies this link by presenting a joint model for a longitudinal process (e.g., Prostate-Specific Antigen measurements) and a multi-state process (e.g., clinical progressions in prostate cancer).
+[ferrer2016joint] exemplifies this link by presenting a joint model for a longitudinal process (e.g., Prostate-Specific Antigen (PSA) measurements) and a multi-state process (e.g., clinical progressions in prostate cancer).
 These two sub-models are interconnected by shared random effects, allowing the model to account for the correlation between the continuous longitudinal biomarker trajectory and the discrete transitions between health states.
-Several software packages have been developed in R and Python to fit joint models that combine longitudinal and time-to-event data.
-The JMbayes package [Rizopoulos, 2016] offers a Bayesian implementation of joint models using Markov chain Monte Carlo (MCMC) methods.
-Its successor, JMbayes2 [Rizopoulos et al., 2024], extends this framework to accommodate more complex data structures, such as multiple longitudinal outcomes, nonlinear trajectories, and competing risks.
-Another popular R package, joineR [Philipson et al., 2018], enables joint modeling of repeated measurements and survival outcomes, based on the methodology introduced by Williamson et al.
-[2008].
-In the Python ecosystem, the flash package [Nguyen et al., 2024] provides a scalable joint modeling approach tailored to high-dimensional settings, exploiting sparse association patterns between longitudinal and survival processes.
-Despite the richness of these tools, none currently supports joint models in the multi-state framework with nonlinear longitudinal modeling.
-In this work, we build on the approach of Ferrer et al.
-[Ferrer, Rondeau, Dignam, Pickles, Jacqmin-Gadda, and Proust-Lima, 2016] and introduce jmstate, a Python package for joint modeling of nonlinear longitudinal and multi-state time-to-event data.
-The underlying model of jmstate incorporates a linear mixed-effects submodel for the longitudinal biomarker and a multi-state process for the event history, linked through shared random effects as in [Ferrer, Rondeau, Dignam, Pickles, Jacqmin-Gadda, and Proust-Lima, 2016].
 The model we propose operates on an arbitrary directed graph, enabling representation of complex state transitions and recurrent events.
-We derive the full likelihood for the nonlinear joint model, introduce an efficient stochastic approximation inference method, and develop dynamic prediction tools.
-To illustrate its practical relevance, we conduct both a simulation study and a synthetic biomedical case study using the jmstate PyPI package.
+We derive the complete likelihood for the nonlinear joint model, introduce an efficient stochastic approximation inference method, and develop dynamic prediction tools.
+To illustrate its practical relevance, we conduct both a simulation study and a synthetic biomedical case study.
 This paper is organized as follows.
-Section 2 presents background on joint modeling frameworks and multi-state processes.
-Section 3 introduces the proposed multi-state joint model, detailing its likelihood formulation and inference strategies.
-Section 4 develops dynamic prediction tools.
-Section 5 details efficient inference methods based on stochastic approximation.
-Section 6 provides implementation details.
-Section 7 exemplifies the use of jmstate though a simulation study while Section 8 illustrates the use of jmstate on data from the PAQUID cohort.
-Finally, Section 9 concludes with a discussion of the main strengths, limitations, and future directions of this work.
+Section 2 reviews the background on multi-state and joint modeling frameworks.
+Section 3.2 introduces the proposed general multi-state joint modeling framework, detailing the likelihood formulation and identifiability conditions.
+Section 4 describes the stochastic-gradient–based inference algorithm, while Section 5 presents the dynamic prediction framework.
+Section 6 reports simulation experiments that assess parameter recovery, identifiability, and convergence properties of the proposed model, and Section 7 applies the framework to data from the PAQUID cohort.
+Finally, Section 8 concludes with perspectives and future research directions.
 2 Background
-2.1 Notations
-Each individual $i$, where $i=1,\ldots,n$, is observed at times $t_{ij}$, where $j=1,\ldots,n_{i}$.
-$n_{i}$ is the number of observations for an individual $i$.
-The biomarker values at these times are denoted by $Y_{ij}\in\mathbb{R}^{d}$, where $d$ is the number of biomarkers.
-Each individual also has a set of covariates $X_{i}\in\mathbb{R}^{k}$, which may include demographic or clinical characteristics.
-2.2 Standard Joint Modeling Framework
-Joint modeling frameworks are essential to link longitudinal and time-to-event data for more accurate dynamic predictions. The standard framework [Rizopoulos, 2012, 2011] consists of:
-  • Mixed-Effects for Longitudinal Data. The longitudinal process $Y_{ij}$ for subject $i$ at time $t_{ij}$ is typically modeled as:
-    $Y_{ij}=h(t_{ij},\psi_{i})+\epsilon_{ij},\quad\psi_{i}=f(\gamma,X_{i},b_{i}),$
-    where $\gamma$ are the population parameters, $b_{i}$ are individual random effects, and $\epsilon_{ij}$ is a random measurement error term.
-It is usually assumed to be Gaussian with zero mean.
-The functions $h$ and $f$ can be nonlinear, capturing complex within-subject trajectories and between-subject variability.
-In practice, $h$ is often a linear predictor of the biomarker values, while $f$ may represent a nonlinear function of the random effects.
-    Both $h$ and $f$ may be nonlinear, allowing flexible modeling of complex biomarker dynamics. Typical choices include:
-    • $h(t,\psi_{i})=\psi_{i1}+\psi_{i2}t+\psi_{i3}t^{2}$ (polynomial growth),
-    • $h(t,\psi_{i})=\psi_{i1}\exp(-\psi_{i2}t)$ (exponential decay),
-    • $f(\gamma,X_{i},b_{i})=\gamma X_{i}+b_{i}$ [Davidian and Giltinan, 2003], $\gamma+b_{i}$, or $\exp(\gamma+b_{i})$ for log-normally distributed parameters.
-  • Cox-Based Survival Modeling. The survival process is described by a hazard function incorporating the current (and possibly historical) longitudinal process and baseline covariates:
-    $\lambda_{i}(t\mid X_{i},\mathcal{H}_{i}(t))=\lambda_{0}(t)\exp\left(\alpha\,g(t,\psi_{i})+\beta X_{i}\right),$
-    where $\mathcal{H}_{i}(t)$ is the information that would be provided by knowing the "true trajectory" of the longitudinal biomarker process, that is, $\mathcal{H}_{i}(t)=\{(u,h(u,\psi_{i})):0\leq u\leq t\}$ and $g(\cdot)$ links the biomarker trajectory to the hazard, with $\alpha$ et $\beta$ as association parameters.
-The baseline hazard $\lambda_{0}(t)$ may be parametric or semiparametric with a spline basis representation [Andrinopoulou et al., 2018].
-  • Prior distribution.
-A prior has to be specified for the latent variables $b_{i}$.
-Here, we assume it to be Gaussian with zero mean: $\mathcal{N}(0,Q)$ but other prior distributions can be considered.
-In joint modeling frameworks, the longitudinal process is linked to the time-to-event component through a function $g$ of the underlying trajectory.
-Several choices for $g$ have been proposed, depending on the hypothesized biological mechanism.
-Current examples include [Papageorgiou et al., 2019, Mauff, 2023]:
-  • $g(t,\psi_{i})=h(t,\psi_{i})$: direct effect of the current biomarker level,
-  • $g(t,\psi_{i})=\frac{\partial}{\partial t}h(t,\psi_{i})$: effect of the biomarker slope (rate of change),
-  • $g(t,\psi_{i})=\int_{0}^{t}h(w,\psi_{i})\,dw$: cumulative exposure.
-This joint framework not only captures the dynamics of longitudinal biomarkers but also leverages their predictive power to inform survival outcomes, enabling dynamic, individualized risk predictions [Rizopoulos, 2011, Ibrahim et al., 2010].
-Moreover, in order to derive an expression for the marginal (joint) likelihood, joint modeling typically requires the following independence assumptions [Rizopoulos, 2012]:
-Assumption 1.
-The latent variables $(b_{i})_{i}$ are independent across individuals.
-Assumption 2.
-The observed markers $(Y_{ij})_{ij}$ are independent conditionally on $(b_{i})_{i}$.
-Assumption 3.
-The occurrence times $(T_{i}^{*})_{i}$ are independent conditionally on $(b_{i})_{i}$.
-Assumption 4.
-The occurrence times $(T_{i}^{*})_{i}$ and the observed markers $(Y_{ij})_{ij}$ are mutually independent conditionally on $(b_{i})_{i}$.
-Assumption 5.
-The censoring times $(C_{i})_{i}$ are independent and noninformative: conditional on $(b_{i})_{i}$ (and possibly covariates), $(C_{i})_{i}$ is independent of both the event times $(T_{i}^{*})_{i}$ and the longitudinal measurements $(Y_{ij})_{ij}$.
-These assumptions facilitate the derivation of the marginal likelihood by allowing the decomposition of the joint distribution into manageable components.
-They ensure that the latent variables capture all relevant dependencies between the longitudinal and survival processes, enabling efficient estimation and inference.
+Joint modeling of longitudinal and time-to-event data provides a unified statistical framework to analyze the interplay between continuous biomarker dynamics and event occurrence processes.
+In its classical form, this framework couples a mixed-effects model describing individual biomarker trajectories with a survival model for event times [rizopoulos2012joint].
+Such models have become central in biomedical research, particularly for dynamic prediction and personalized risk assessment [papageorgiou2019overview].
+However, these traditional joint models generally focus on a single terminal event, limiting their ability to represent more complex event histories that include intermediate, recurrent, or competing events.
+To address these limitations, multi-state extensions have been developed, providing a natural framework for describing transitions between multiple clinical or biological states over time [ferrer2016joint, KrolMauguen2017].
 2.1 Multi-State Markov Processes
@@
 This Markov property means the process is memoryless.
-A key consequence is that the duration spent in any state follows an exponential distribution, implying a constant hazard rate for leaving that state [Jackson, 2011, Putter et al., 2007].
+A key consequence is that the duration spent in any state follows an exponential distribution, implying a constant hazard rate for leaving that state [jackson2011multi, putter2007competing].
 However, the Markov assumption can be unrealistic in many real-world applications.
-For example, in the study of human sleep stages, sojourn times often do not follow an exponential distribution [Dong et al., 2008, Roever et al., 2010], and in the case of chronic diseases such as AIDS, the risk of disease progression can depend on the time elapsed since infection [Andersen and Keiding, 1999].
+For example, in the study of human sleep stages, sojourn times often do not follow an exponential distribution [dong2008non, roever2010statistical], and in the case of chronic diseases such as AIDS, the risk of disease progression can depend on the time elapsed since infection [andersen1999multi].
 To address these limitations, semi-Markov processes (SMPs) were introduced, which allow for arbitrary distributions of sojourn times while retaining the Markov property for the embedded discrete-time chain.
-This flexibility makes SMPs suitable for modeling complex disease progression and patient recovery scenarios [Putter et al., 2007].
+This flexibility makes SMPs suitable for modeling complex disease progression and patient recovery scenarios [putter2007competing].
 2.2 Multi-State Semi-Markov Processes
@@
 In an MSMP, the process is defined by a directed graph $G=(V,E)$, where $V$ represents the set of states and $E$ the set of possible transitions.
-The transition intensities $\lambda_{k^{\prime}|k}(t\mid t_{0})$ from state $k$ to state $k^{\prime}$ at time $t$, given entry time $t_{0}$.
+The transition intensities $\lambda_{k^{\prime}\mid k}(t\mid t_{0})$ from state $k$ to state $k^{\prime}$ at time $t$, given entry time $t_{0}$ are therefore assumed to be non-identically zero.
 The sojourn time in state $k$ is not restricted to an exponential distribution, allowing for more realistic modeling of the time spent in each state.
 MSMPs have been widely applied in a range of disciplines.
-In reliability, they are used to modeling degradation and repair processes [Limnios and Oprisan, 2001].
-In biomedical studies, they have proven useful for analyzing illness-death models or disease progression with non-exponential transitions [Commenges et al., 2006, Fiocco et al., 2008].
-Applications in finance also exist, where semi-Markovian dynamics can model credit rating migrations or economic regimes [Luciano and Vasiliev, 2006].
+In reliability, they are used to modeling degradation and repair processes [limnios2001semi].
+In biomedical studies, they have proven useful for analyzing illness-death models or disease progression with non-exponential transitions [commenges2006likelihood, fiocco2008modelling].
+Applications in finance also exist, where semi-Markovian dynamics can model credit rating migrations or economic regimes [luciano2006semi].
 In these contexts, MSMPs retain the Markov property in the embedded jump chain while offering increased realism through flexible dwell time modeling.
-From a methodological standpoint, MSMPs extend estimation strategies developed for Markov models, such as maximum likelihood or Bayesian inference, and can accommodate interval-censored or misclassified data [Frydman, 2005].
+From a methodological standpoint, MSMPs extend estimation strategies developed for Markov models, such as maximum likelihood or Bayesian inference, and can accommodate interval-censored or misclassified data [frydman2005estimation].
 This makes them a powerful and general tool for multi-state event history analysis.
@@
 Previous research has initiated such integration, with models like the one proposed by Ferrer et al.
-[Ferrer, Rondeau, Dignam, Pickles, Jacqmin-Gadda, and Proust-Lima, 2016], which combines a linear mixed sub-model for longitudinal data with a multi-state sub-model using shared random effects.
+[ferrer2016joint], which combines a linear mixed sub-model for longitudinal data with a multi-state sub-model using shared random effects.
 This model enables the simultaneous analysis of repeated measurements of a biomarker (the longitudinal process, e.g., PSA levels) and the times of transitions between multiple health states (the multi-state process).
@@
 Although Ferrer et al.
-[Ferrer et al., 2016] note that their joint modeling framework could be adapted to semi-Markov processes in other contexts, their primary application to prostate cancer progression assumes a non-homogeneous Markov multi-state process.
-Consequently, the model does not explicitly incorporate the effect of time spent within a particular state on the instantaneous risk of transitioning out of that state, beyond the overall follow-up time from study entry.
-While this prior work has managed to combine multi-state and joint models in particular applications, these models typically rely on a linear mixed model for the biomarker and impose a specific directed acyclic graph (DAG) structure through their likelihood decomposition, which restricts the types of dependency structures they can represent.
-As a result, limitations remain in their generalizability, flexibility, and the scalability of inference.
-In what follows, we propose a more general framework defined on arbitrary directed graphs (not restricted to acyclic forms) that accommodates both Markov and semi-Markov assumptions and allows nonlinear longitudinal submodels.
-3 Extending Multi-State Joint Models
+[ferrer2016joint] note that their joint modeling framework could be adapted to semi-Markov processes in other contexts, their primary application to prostate cancer progression assumes a non-homogeneous Markov multi-state process.
+However, this framework is not well suited for arbitrary transition patterns other than the ones modeled by DAGs.
+While both joint models and multi-state models offer powerful tools for analyzing disease progression, they emphasize different aspects of the underlying process.
+Multi-state models capture the sequence and timing of discrete transitions between states, whereas joint models focus on the evolution of continuous longitudinal markers and their association with event risk.
+Existing joint multi-state models, such as that of Ferrer et al.
+[ferrer2016joint], are restricted to linear mixed-effects submodels and to acyclic Markov transition graphs, as their likelihood relies on a sequential factorization that is not valid when transitions form cycles.
+However, many biological processes involve cyclic or semi-Markov transitions and nonlinear biomarker dynamics.
+These cannot be handled by existing frameworks, motivating the present general formulation.
+In contrast, many real-world processes—such as recurrent disease relapse or recovery—require more flexible representations involving nonlinear biomarker dynamics, semi-Markovian dependencies, and recurrent cycles.
+The next section introduces a general framework that integrates these features within a unified likelihood-based formulation, enabling efficient inference and dynamic prediction along complex event trajectories.
+3 A General Multi-State Joint Modeling Framework
+Classical joint models typically link a longitudinal biomarker process to a single time-to-event outcome.
+While effective for terminal events, this design cannot represent more complex disease trajectories involving intermediate or recurrent events.
+To address these limitations, we extend the nonlinear joint modeling framework to incorporate a multi-state process, thereby enabling a unified analysis of longitudinal biomarkers and discrete transitions between multiple health states.
+The proposed model defines a joint process $\{Y_{i}(t),S_{i}(t)\}_{t\geq 0}$, where $Y_{i}(t)$ denotes the longitudinal biomarker trajectory and $S_{i}(t)$ the multi-state event process.
+Both components are linked through shared latent structures, and the event process is represented on an arbitrary directed graph supporting both Markovian and semi-Markovian transition mechanisms.
+3.1 Notation Conventions
+Throughout this section, $i=1,\dots,n$ indexes individuals, and $t_{ij}$ denotes the $j$-th observation time of individual $i$.
+The vector $Y_{ij}\in\mathbb{R}^{d}$ represents longitudinal biomarker measurements, and $X_{i}\in\mathbb{R}^{p}$ the associated covariates.
+The latent random effects $b_{i}\sim\mathcal{N}(0,Q)$ capture individual heterogeneity, while $\psi_{i}=f(\gamma,X_{i},b_{i})$ denotes the subject-specific parameters in the longitudinal submodel.
+The filtration $\mathcal{H}_{i}(t)$ represents the observed history of the longitudinal process up to time $t$.
+Transition times and states are denoted by $(T_{i\ell},S_{i\ell})$, with right-censoring time $C_{i}$.
+We also define the index of the last observed transition before time $t$ denoted by $m_{i}(t)\coloneq\sup\{\ell\geq 0:T_{i\ell}\leq t\}$.
+3.2 Extending Multi-State Joint Models
 Joint models traditionally focus on a single time-to-event outcome, which limits their ability to represent complex disease trajectories involving multiple intermediate and recurrent events.
 To address this limitation, we extend the classical nonlinear joint modeling framework to incorporate multi-state processes, enabling the joint analysis of longitudinal biomarkers and transitions between multiple health states.
-This is achieved by introducing a directed graph structure $G=(V,E)$, which encodes all permissible transitions and supports both Markov and semi-Markov assumptions.
-3.1 Graph Structure and Example
-Let us consider the illustrative state-transition diagram shown in Figure 1, representing a 4-state process.
-Figure 1: Illustration of the 4-state transition graph $G=(V,E)$, including absorbing state 4 (Dead).
-This graph can be encoded by an adjacency matrix $A=[A_{kk^{\prime}}]_{kk^{\prime}}$ where $A_{kk^{\prime}}=1$ if a transition from state $k$ to $k^{\prime}$ is possible, and 0 otherwise.
-For the graph in Figure 1, the adjacency matrix is:
-$A=\begin{pmatrix}0&1&0&1\\ 0&0&1&1\\ 0&1&0&1\\ 0&0&0&0\end{pmatrix}$
-Under certain assumptions borrowed from the semi-Markov literature, competing and recurrent risks mentioned in the introduction, as well as the classical joint model, can be represented using such graphs as illustrated in Figure 2.
+This is achieved by introducing a directed graph structure encoding all permissible transitions and supports both Markov and semi-Markov assumptions.
+3.2.1 Graph Structure
+Let $G=(V,E)$ be a directed graph, where $V=\{1,\dots,p\}$ denotes the set of states and $E\subset V\times V$ the set of admissible transitions.
+The graph encodes all possible paths of the multi-state process, allowing for competing, recurrent, or absorbing transitions.
+Figure 1 illustrates an example of a four-state model including an absorbing state (Death).
+The corresponding adjacency matrix $A=(A_{kk^{\prime}})_{1\leq k,k^{\prime}\leq|V|}$ satisfies $A_{kk^{\prime}}=1$ if $(k,k^{\prime})\in E$ and 0 otherwise.
+Figure 1: Example of a 4-state transition graph $G=(V,E)$, including an absorbing state (4).
+Such a representation generalizes standard joint models, which can be recovered as special cases: single-event (linear chain), competing-risks (two absorbing transitions), or recurrent-event settings (cyclic transitions) as illustrated in Figure 2.
 Standard Risk
@@
 Such joint modeling offers a flexible framework that extends beyond classical approaches and is particularly useful in medical applications.
-3.2 Individual Trajectories and the Markov Assumption
-Each individual $i$ follows a true trajectory:
-$\mathcal{T}_{i}^{*}=\left((T_{i0},S_{i0}),(T_{i1},S_{i1}),(T_{i2},S_{i2}),\ldots\right),$
-where $T_{il}$ denotes the $l$-th transition time and $S_{il}\in V$ the corresponding state. Between transitions, the state $S_{i}(t)$ is assumed constant.
-Only a finite portion of the trajectory is observed, up to a right-censoring time $C_{i}$:
-$\big(\mathcal{T}_{i},C_{i}\big)=\Big(\big((T_{i0},S_{i0}),\dots,(T_{im_{i}},S_{im_{i}})\big),C_{i}\Big),$
-where $m_{i}(t)=\max\{l:T_{il}\leq t\}$, and $m_{i}=m_{i}(C_{i})$ denotes the last observed transition index.
-We assume that the process satisfies the (time-inhomogeneous) Markov property:
-$\mathscr{L}\big((T_{i(l+1)},S_{i(l+1)})\mid\{(T_{il^{\prime}},S_{il^{\prime}})\}_{l^{\prime}\leq l}\big)=\mathscr{L}\big((T_{i(l+1)},S_{i(l+1)})\mid(T_{il},S_{il})\big).$ ($\ast$)
-Remark 1.
-The transition times are not restricted to be non-negative.
-3.3 Transition Intensities and Survival Functions
-Let $(k,k^{\prime})\in V^{2}$. The instantaneous risk of transition from state $k$ to state $k^{\prime}$ at time $t$, given entry at time $t_{0}$, is defined as:
-$\lambda_{i}^{k^{\prime}\mid k}(t\mid t_{0})\coloneqq\lim_{\delta\to 0^{+}}\frac{\mathbb{P}(T_{i(l+1)}\leq t+\delta,\ S_{i(l+1)}=k^{\prime}\mid T_{i(l+1)}>t,\ T_{il}=t_{0},\ S_{il}=k)}{\delta}.$
-Summing over all admissible transitions from $k$, we get the total instantaneous risk of leaving state $k$:
-$\sum_{s:\,(k,s)\in E}\lambda_{i}^{s\mid k}(t\mid t_{0})=\lim_{\delta\to 0^{+}}\frac{\mathbb{P}(T_{i(l+1)}\leq t+\delta\mid T_{i(l+1)}>t,T_{il}=t_{0},S_{il}=k)}{\delta}.$
-The cumulative intensity of the transition from $k$ to $k^{\prime}$ is defined by:
+3.2.2 Individual Trajectories
+Each individual $i$ follows a latent trajectory
+$\mathcal{T}_{i}^{*}=\big((T_{i0},S_{i0}),(T_{i1},S_{i1}),(T_{i2},S_{i2}),\dots\big),$
+where $T_{i\ell}$ denotes the $\ell$-th transition time and $S_{i\ell}\in V$ the corresponding state.
+The observed trajectory is right-censored at time $C_{i}$, so that only $\{(T_{i\ell},S_{i\ell}):T_{i\ell}\leq C_{i}\}$ is observed.
+Between transitions, the state $S_{i}(t)$ is piecewise constant.
+We note $m_{i}(t)\coloneq\sup\{\ell\geq 0:T_{i\ell}\leq t\}$ the index of the last observed transition before time $t$ and $m_{i}=m_{i}(C_{i})$.
+The process is assumed to satisfy a (possibly time-inhomogeneous) semi-Markov property
+$\mathscr{L}\big((T_{i,\ell+1},S_{i,\ell+1})\mid\{(T_{i\ell^{\prime}},S_{i\ell^{\prime}})\}_{\ell^{\prime}\leq\ell}\big)=\mathscr{L}\big((T_{i,\ell+1},S_{i,\ell+1})\mid(T_{i\ell},S_{i\ell})\big),$ (1)
+which is relaxed compared to a stricter Markov property by allowing sojourn-time dependence.
+3.2.3 Transition Intensities and Survival Functions
+For any admissible transition $(k,k^{\prime})\in E$, let $\lambda_{i}^{k^{\prime}\mid k}(t\mid t_{0})$ denote the instantaneous risk of moving from $k$ to $k^{\prime}$ at time $t$, given entry time $t_{0}$
+$\lambda_{i}^{k^{\prime}\mid k}(t\mid t_{0})=\lim_{\delta\to 0^{+}}\frac{\mathbb{P}(T_{i,\ell+1}\leq t+\delta,S_{i,\ell+1}=k^{\prime}\mid T_{i,\ell+1}>t,T_{i\ell}=t_{0},S_{i\ell}=k)}{\delta}.$
+The cumulative intensity and the corresponding survival function are
 $\Lambda_{i}^{k^{\prime}\mid k}(t\mid t_{0})=\int_{t_{0}}^{t}\lambda_{i}^{k^{\prime}\mid k}(w\mid t_{0})\,dw,$
-which leads to the corresponding survival function:
-$\mathbb{P}(T_{i(l+1)}>t\mid T_{il},S_{il})=\exp\left(-\sum_{s:\,(S_{il},s)\in E}\Lambda_{i}^{s\mid S_{il}}(t\mid T_{il})\right).$
-Conditional transition probabilities are given by:
-$\mathbb{P}(S_{i(l+1)}\mid T_{i(l+1)},T_{il},S_{il})=\frac{\lambda_{i}^{S_{i(l+1)}\mid S_{il}}(T_{i(l+1)}\mid T_{il})}{\sum_{s:\,(S_{il},s)\in E}\lambda_{i}^{s\mid S_{il}}(T_{i(l+1)}\mid T_{il})}.$
-Combining both, the joint density is:
-$p(T_{i(l+1)},S_{i(l+1)}\mid T_{il},S_{il})=\lambda_{i}^{S_{i(l+1)}\mid S_{il}}(T_{i(l+1)}\mid T_{il})\,\exp\left(-\sum_{s:\,(S_{il},s)\in E}\Lambda_{i}^{s\mid S_{il}}(T_{i(l+1)}\mid T_{il})\right).$
-Furthermore, conditionning on the event $T_{i(j+1)}\geq t$ for some known $t$ can also easily be achieved by using the Chasles relation:
-$\forall u\geq t,$
-$\mathbb{P}(T_{i(l+1)}>u\mid T_{il},S_{il},T_{i(l+1)}>t)=\exp\left(-\sum_{s:\,(S_{il},s)\in E}\int_{t}^{u}\lambda_{i}^{s\mid S_{il}}(w\mid T_{il})\,dw\right).$
+$\mathbb{P}(T_{i,\ell+1}>t\mid T_{i\ell},S_{i\ell})=\exp\left[-\sum_{s:(S_{i\ell},s)\in E}\Lambda_{i}^{s|S_{i\ell}}(t\mid T_{i\ell})\right].$
+Conditional transition probabilities are
+$\mathbb{P}(S_{i,\ell+1}=k^{\prime}\mid T_{i,\ell+1},T_{i\ell},S_{i\ell})=\frac{\lambda_{i}^{k^{\prime}|S_{i\ell}}(T_{i,\ell+1}\mid T_{i\ell})}{\sum_{s:(S_{i\ell},s)\in E}\lambda_{i}^{s|S_{i\ell}}(T_{i,\ell+1}\mid T_{i\ell})}.$
 3.3 Multi-State Joint Model Specification
 The proposed multi-state joint model with Gaussian prior and homoscedastic Gaussian noise is specified by:
-$\begin{cases}Y_{ij}=h(t_{ij},\psi_{i})+\epsilon_{ij},\\ \psi_{i}=f(\gamma,X_{i},b_{i}),\quad b_{i}\sim\mathcal{N}(0,Q),\quad\epsilon_{ij}\sim\mathcal{N}(0,R),\\ \lambda_{i}^{k^{\prime}\mid k}(t\mid X_{i},T_{im_{i}(t)},\mathcal{H}_{i}(t))=\lambda_{0}^{k^{\prime}\mid k}(t\mid T_{im_{i}(t)})\cdot\exp\big(\alpha^{k^{\prime}\mid k}g^{k^{\prime}\mid k}(t,X_{i},\psi_{i})+\beta^{k^{\prime}\mid k}X_{i}\big),\end{cases}$
-along with the assumptions 1, 2, and 5 as well as the following assumptions, analogous to assumptions 3 and 4:
-Assumption 6.
-Conditionally on $b_{i}$, the trajectories $\mathcal{T}_{i}^{*}$ are independent and satisfy the semi-Markov property $(\ast)$.
-Assumption 7.
-Conditionally on $b_{i}$, the biomarker process $Y_{ij}$ is independent of the transition process $\mathcal{T}_{i}^{*}$.
-3.5 Clock Reset vs. Clock Forward
-Two baseline hazard specifications are commonly used:
+$\begin{cases}Y_{ij}=h(t_{ij},\psi_{i})+\epsilon_{ij},\quad\epsilon_{ij}\sim\mathcal{N}(0,R),\\ \psi_{i}=f(\gamma,X_{i},b_{i}),\quad b_{i}\sim\mathcal{N}(0,Q),\\ \lambda_{i}^{k^{\prime}\mid k}\left(t\mid X_{i},T_{im_{i}(t)},\mathcal{H}_{i}(t)\right)=\lambda_{0}^{k^{\prime}\mid k}(t\mid T_{im_{i}(t)})\exp\big(\alpha^{k^{\prime}\mid k}g^{k^{\prime}\mid k}(t,X_{i},\psi_{i})+\beta^{k^{\prime}\mid k}X_{i}\big).\end{cases}$
+Here, $h$ and $f$ define the nonlinear mixed-effects submodel, $g^{k^{\prime}\mid k}$ represents the link between biomarker dynamics and the transition intensity, and $\lambda_{0}^{k^{\prime}\mid k}$ denotes the baseline hazard associated with transition $(k\to k^{\prime})$.
+This structure extends the classical joint model to arbitrary directed graphs and to both Markovian and semi-Markov specifications.
+The coefficient $\alpha^{k^{\prime}\mid k}$ quantifies the effect of the longitudinal biomarker dynamics on the instantaneous risk of transitioning from state $k$ to state $k^{\prime}$.
+More concretely, for a one-unit increase in $g^{k^{\prime}\mid k}(t,X_{i},\psi_{i})$ (holding other covariates constant), the hazard ratio for transition $(k\to k^{\prime})$ is $\exp(\alpha^{k^{\prime}\mid k})$.
+Likewise, $\beta^{k^{\prime}\mid k}$ represents the effect of baseline (or time-varying) covariates $X_{i}$ on that same transition: a one-unit increase in a covariate $X_{ij}$ multiplies the hazard by $\exp(\beta^{k^{\prime}\mid k}_{j})$.
+Thus, $\alpha^{k^{\prime}\mid k}$ captures the degree to which the biomarker trajectory (via its latent parameter $\psi_{i}$ or function $g^{k^{\prime}\mid k}$) influences the transition risk, while $\beta^{k^{\prime}\mid k}$ captures the direct effect of covariates on that transition’s risk, beyond the biomarker pathway.
+3.3.1 Model Variants
+Two baseline hazard conventions are common:
 $\lambda_{0}^{k^{\prime}\mid k}(t\mid T_{im_{i}(t)})=\begin{cases}\lambda_{0}^{k^{\prime}\mid k}(t-T_{im_{i}(t)})&\text{(clock reset)},\\ \lambda_{0}^{k^{\prime}\mid k}(t)&\text{(clock forward)}.\end{cases}$
-The clock reset model resets the baseline hazard at each state entry and is appropriate when the transition risk depends on time spent in the state (Asanjarani et al.
-[2022]).
-The clock forward model assumes that risk accumulates with global time, ignoring the state entry time.
-3.6 Initial State
-We assume that the initial state $S_{i0}$ is observed. However, the framework could also incorporate a multinomial model for unobserved initial states (Yiu et al. [2018]).
-This flexible framework enables inference on complex event histories, leveraging the full trajectory of biomarkers to refine predictions of transitions and survival [de Wreede et al., 2010].
-3.7 Multi-State Simulation
-In nonlinear joint models with known design and known parameters parameters $\theta=(\gamma,Q,R,\alpha,\beta)$, the occurrence times of events conditionally on the latent variables $b_{i}$ and the covariates $X_{i}$ can be easily sampled using the inverse cumulative distribution function transform method.
-This may be achieved through bisection or other root-finding algorithms.
-The simulation of the transition process $\mathcal{T}_{i}^{*}$ conditionally on $b_{i}$, $X_{i}$, and $(T_{i0},S_{i0})$ can be achieved drawing on the semi-Markov property 6 by considering one transition at a time, according to Algorithm 1.
-The algorithme is similar to Gillespie’s algorithm [Gillespie, 2007], and can also include a survival condition $T_{i1}\geq t_{i}^{\text{surv}}$ that uses the Chasles relation LABEL:rel:chasles.
-Algorithm 1 Simulation of trajectory $i$
-1:subject-specific censoring time $C_{i}\in\mathbb{R}\cup\{+\infty\}$; covariates $X_{i}$; latent variables $b_{i}$; initial time-state pair $(T_{i0},S_{i0})$; optional survival condition $T_{i1}\geq t_{i}^{\text{surv}}$, defaults to $t_{i}^{\text{surv}}=$-$\infty$.
-2:
-3:Initialize $\ell_{i}\leftarrow 1$, $\mathcal{T}_{i}\leftarrow\left((T_{i0},S_{i0})\right)$.
-4:
-5:while $T_{i(\ell_{i}-1)}<C_{i}$ and $\{s:(S_{i(\ell_{i}-1)},s)\in E\}\neq\emptyset$ do
-6:  for each $s$ with $(S_{i(\ell_{i}-1)},s)\in E$ do
-7:   Draw $T_{i\ell_{i}}^{s}$ such that
-8:      $\log\mathbb{P}(T_{i\ell_{i}}^{s}>t)=\int_{T_{i(\ell_{i}-1)}\vee t_{i}^{\text{surv}}}^{t}\lambda_{i}^{s\mid S_{i(\ell_{i}-1)}}(w\mid\theta,X_{i},T_{i(\ell_{i}-1)},b_{i})\,dw,$
-9:      $\forall t\geq T_{i(\ell_{i}-1)}\vee u.$
-10:  end for
-11:
-12:  Set $T_{i\ell_{i}}\leftarrow\min_{m}T_{i\ell_{i}}^{s}$, and $S_{i\ell_{i}}\leftarrow\arg\min_{m}T_{i\ell_{i}}^{s}$.
-13:  Append $(T_{i\ell_{i}},S_{i\ell_{i}})$ to $\mathcal{T}_{i}$.
-14:  $\ell_{i}\leftarrow\ell_{i}+1$.
-15:end while
-16:
-17:if $T_{i(\ell_{i}-1)}>C_{i}$ then
-18:  Remove the last pair: $\mathcal{T}_{i}\leftarrow\mathcal{T}_{i}[:-1]$.
-19:end if
-[…]
-4 Dynamic Prediction
-We now focus on constructing random variables of interest for multi-state prediction for a new individual $i$, such that these random variables may be feasibly simulated with Algorithm 1.
-Let $\tau_{i}$ be a stopping time adapted to the filtration $\mathcal{F}_{in}=\sigma(T_{i0},S_{i0},\dots,T_{in},S_{in})$, and let $\kappa_{i}=\{\inf n\in\mathbb{N}:\mathbb{P}(\tau_{i}=+\infty\mid\mathcal{F}_{in})=1\}$ be a second stopping time.
-Assume that:
-$\mathbb{P}(\min(\tau_{i},\kappa_{i})<+\infty)=1.$
-Here, $\tau_{i}$ represents the time at which an event occurs for an individual $i$, while $\kappa_{i}$ is the time at which it becomes certain that the event will not occur.
-For instance, if the event of interest is patient remission, but the patient dies without remission, then $\tau_{i}=+\infty$, while $\kappa_{i}$ would be finite.
-As such, it is almost surely possible to simulate the trajectory of an individual $i$ up to the time of the event or the time of censoring.
-Let $\xi$ be a function taking values in some space $\mathbb{Y}$:
-$\xi:$ $(\mathbb{N}\cup\{+\infty\})^{2}\times(\mathbb{R}\times V)^{\mathbb{N}}\longrightarrow\mathbb{Y}$
-$\tau_{i},\kappa_{i},\mathcal{T}_{i}^{*}\longmapsto\xi(\tau_{i},\kappa_{i},\mathcal{T}_{i}^{*})=\xi\left(\tau_{i},\kappa_{i},(T_{il},S_{il})_{l\leq\min(\tau_{i},\kappa_{i})}\right)$
-This function therefore only depends on the beginning of the trajectory and not on its entirety.
-Thus, during simulations, if $\tau_{i}$ is finite or almost surely infinite after a certain index, there is no need to simulate the entire trajectory.
-Applications include, for instance, the distributions of states at any time, transition times, return times, sojourn times, hitting times and many other quantities.
-This flexibility is illustrated by Examples 1 and 2.
-Example 1 (State at time $t$).
-Let $G=(V,E)$ be a directed acyclic graph, $t\in\mathbb{R}$ be a fixed time, and $\mathbb{Y}=V$ be the set of possible states.
-Let $\tau_{i}=\inf\{n\in\mathbb{N}:T_{in}\geq t\text{ or }S_{in}\text{ is absorbing}\}$.
-Clearly, $\tau_{i}\leq\operatorname{depth}(G)<+\infty$.
-Then, take $\xi_{t}(\tau_{i},\kappa_{i},\mathcal{T}_{i}^{*})=S_{i\tau_{i}}$, which corresponds to the state of the individual $i$ at time $t$.
-Example 2 (Hitting time).
-Let $G=(V,E)$ be a directed acyclic graph, $\mathbb{A}\subset V$ a non-empty subset of states, and $\mathbb{Y}=\mathbb{R}\cup\{+\infty\}$.
-For any two non-empty sets $\mathbb{A},\,\mathbb{B}\subset V$, we note $\mathbb{A}\rightsquigarrow\mathbb{B}$ if there exists a path from $\mathbb{A}$ to $\mathbb{B}$ in $G$.
-Let $\tau_{i}=\inf\{n\in\mathbb{N}:S_{in}\in\mathbb{A}\}$, with the convention $\inf\emptyset=+\infty$.
-Then, $\kappa_{i}=\inf\{n\in\mathbb{N}:\{S_{in}\}\not\rightsquigarrow\mathbb{A}\}$ and $\min(\tau_{i},\kappa_{i})\leq\operatorname{depth}(G)<+\infty$.
-Finally, $\xi_{\mathbb{A}}(\tau_{i},\kappa_{i},\mathcal{T}_{i}^{*})=\mathds{1}_{\tau_{i}<+\infty}T_{i\tau_{i}}+\mathds{1}_{\tau_{i}=+\infty}\tau_{i}$ represents the hitting time for the set $\mathbb{A}$.
-5 Likelihood and Statistical Inference
-The estimation of model parameters $\theta=(\gamma,Q,R,\alpha,\beta)$ relies on two likelihood formulations: the complete-data likelihood $\mathcal{L}_{\text{full}}$, which includes latent variables, and the marginal likelihood $\mathcal{L}_{\text{marginal}}$, obtained by integrating them out:
-$\mathcal{L}_{\text{marginal}}(\theta;X,Y,\mathcal{T},C)=\int\mathcal{L}_{\text{full}}(\theta;;X,Y,\mathcal{T},C,b)\,db,$
-where:
-$\mathcal{L}_{\text{full}}(\theta;;X,Y,\mathcal{T},C,b)=\prod_{i=1}^{n}p(b_{i}\mid\theta)p(\mathcal{T}_{i},C_{i}\mid\theta,X_{i},b_{i})\prod_{j=1}^{m_{i}}p(Y_{ij}\mid\theta,X_{i},b_{i}).$
-This expression is very similar to that obtained by Rizopoulos [2012] and is subdivided into three terms:
+The clock-reset form models risk as a function of time spent in the current state, whereas the clock-forward form measures risk with respect to global time since study entry.
+3.3.2 Likelihood Formulation
+To derive the marginal likelihood, we decompose the joint distribution of the longitudinal and multi-state processes under a set of standard conditional independence assumptions detailed in Table 3.3.2 below.
+Assumptions for Likelihood Factorization A.
+Latent-level independence A1.
+Random effects $(b_{i})_{i}$ are mutually independent across individuals.
+B.
+Conditional independence within individuals B1.
+Longitudinal observations $(Y_{ij})_{ij}$ are mutually independent given $b_{i},X_{i}$.
+B2.
+Trajectories $(\mathcal{T}_{i}^{*})_{i}$ are mutually independent given $b_{i},X_{i}$.
+B3.
+Conditional on $b_{i},X_{i}$, the longitudinal and event processes are mutually independent.
+C.
+Censoring and process assumptions C1.
+Censoring times $(C_{i})_{i}$ are mutually independent and noninformative, i.e.
+$C_{i}\mathrel{\perp\!\!\!\perp}(Y_{i},\mathcal{T}_{i}^{*})\mid b_{i},X_{i}$.
+C2.
+Conditionally on $b_{i},X_{i}$, event trajectories satisfy the semi-Markov property 1: $p(\mathcal{T}_{i}^{*}\mid b_{i},X_{i})=\prod_{\ell\geq 0}p\left((T_{i,\ell+1},S_{i,\ell+1})\mid T_{i\ell},S_{i\ell},b_{i},X_{i}\right)$.
+These grouped assumptions mirror those used in classical joint modeling [rizopoulos2012joint] and multi-state survival analysis [putter2007competing]: independence across subjects; the conditional independence structure linking the longitudinal and event submodels through shared random effects; noninformative censoring and (semi-)Markovian dynamics.
+Let $\theta=(\gamma,Q,R,\alpha,\beta)$ denote the vector of model parameters.
+For subject $i$, we observe the longitudinal measurements $Y_{i}=(Y_{i1},\dots,Y_{in_{i}})$ and the event trajectory $\mathcal{T}_{i}=\{(T_{i\ell},S_{i\ell})_{0\leq\ell\leq m_{i}}\}$.
+The joint likelihood then factorizes as
+$p(Y_{i},\mathcal{T}_{i}\mid X_{i},C_{i},\theta)=\int p(Y_{i}\mid b_{i},X_{i},\theta)\,p(\mathcal{T}_{i}\mid b_{i},X_{i},C_{i},\theta)\,p(b_{i}\mid\theta)\,db_{i},$
+where $b_{i}$ are individual random effects.
+This formulation generalizes the joint likelihoods of wulfsohn1997joint and ferrer2016joint to arbitrary multi-state event structures.
+Each part can then be explicitely expressed using Assumptions 3.3.2, yielding an expression very similar to that obtained by rizopoulos2012joint.
 Prior likelihood:
@@
 Longitudinal likelihood:
-$p(Y_{ij}\mid\theta,X_{i},b_{i})=\frac{(2\pi)^{-d/2}}{\det(R)^{1/2}}\exp\left(-\tfrac{1}{2}\left(Y_{ij}-h(t_{ij},\psi_{i})\right)^{T}R^{-1}\left(Y_{ij}-h(t_{ij},\psi_{i})\right)\right),$
+$p(Y_{i}\mid b_{i},X_{i},\theta)=\prod_{j=1}^{n_{i}}\frac{(2\pi)^{-d/2}}{\det(R)^{1/2}}\exp\left(-\tfrac{1}{2}\left(Y_{ij}-h(t_{ij},\psi_{i})\right)^{T}R^{-1}\left(Y_{ij}-h(t_{ij},\psi_{i})\right)\right),$
 $\text{with }\psi_{i}=f(\gamma,X_{i},b_{i}).$
 Semi-Markov likelihood:
-$p(\mathcal{T}_{i},C_{i}\mid\theta,X_{i},b_{i})={}$ $\prod_{l=0}^{m_{i}-1}p\left((T_{i(l+1)},S_{i(l+1)})\mid\theta,X_{i},(T_{il},S_{il}),b_{i}\right)$ (1)
-$\exp\left(-\sum_{s:\,(S_{im_{i}},s)\in E}\Lambda_{i}^{k^{\prime}\mid S_{im_{i}}}(C_{i}\mid\theta,X_{i},T_{im_{i}},b_{i})\right).$
+$p(\mathcal{T}_{i}\mid b_{i},X_{i},C_{i},\theta)=$ $\prod_{l=0}^{m_{i}-1}p\left((T_{i(l+1)},S_{i(l+1)})\mid b_{i},X_{i},(T_{il},S_{il}),\theta\right)$ (2)
+$\exp\left(-\sum_{s:\,(S_{im_{i}},s)\in E}\Lambda_{i}^{k^{\prime}\mid S_{im_{i}}}(C_{i}\mid b_{i},X_{i},T_{im_{i}},\theta)\right).$
+The proof of the expression of the Semi-Markov likelihood (2) is provided in Appendix B.
+The likelihood factorization above is valid and forms the basis for scalable inference procedures described in Section 4, leveraging the complete trajectory of biomarkers to refine predictions of transitions and survival [de2010mstate].
+Moreover, here, we assume that the initial state $S_{i0}$ is observed. However, the framework could also incorporate a multinomial model for unobserved initial states [yiu2018clustered].
+3.4 Model Selection and Information Criteria
+To compare competing model specifications, we rely on information criteria based on the marginal likelihood as defined in Section 4.
+The Akaike Information Criterion (AIC) [akaike2003new] and the Bayesian Information Criterion (BIC) [schwarz1978estimating] are standard tools.
+While the AIC relies on an asymptotically unbiased estimator of the $\log$-likelihood, the BIC aims to identify the true model with great probability as the sample size grows.
+The Akaike Information Criterion is computed as follows
+$\mathrm{AIC}=-2\log\mathcal{L}_{\text{marginal}}(\hat{\theta};X,Y,\mathcal{T},C)+2k,$
+where $k$ is the number of parameters, and $\hat{\theta}$ is a maximizer of the marginal likelihood.
+The Bayesian Information Criterion is similarly defined but imposes a stronger penalty for model complexity, growing in $\log n$ where $n$ is the number of observations.
+Particularly in the context of mixed effects models, and even more so for multi-state joint models, the number of observations $n$ may vary from one definition to another, either the total number of repeated measurements or the number of individuals [delattre2014note].
+To alleviate this problem, we recall the derivation of the BIC from the Laplace approximation [lebarbier2006introduction].
+For a prior distribution $\pi$ on a set of models $\mathcal{M}$, the Laplace approximation approaches up to constant terms which could also be accounted to improve the approximation, the posterior probability given observed data $x$ defined in Section 4 by
 […]
+where $H_{\hat{\theta}}$ denotes the Hessian matrix of the marginal $\log$-likelihood evaluated at the Maximum Likelihood Estimator (MLE).
+Then, under standard regularity conditions that allow differentiation and integration to be interchanged twice, we have that
+[…]
+[…]
+[…]
+[…]
+Thus, our BIC criterion can be written as
+[…]
+In both cases, one should aim at minimizing the chosen criterion.
+4 Statistical Inference
+The estimation of model parameters $\theta=(\gamma,Q,R,\alpha,\beta)$ relies on two likelihood formulations: the complete-data likelihood $\mathcal{L}_{\text{complete}}$, which includes latent variables, and the marginal likelihood $\mathcal{L}_{\text{marginal}}$, obtained by integrating them out
+$\mathcal{L}_{\text{marginal}}(\theta;X,Y,\mathcal{T},C)=\int\mathcal{L}_{\text{complete}}(\theta;X,Y,\mathcal{T},C,b)\,db,$
+where
+$\mathcal{L}_{\text{complete}}(\theta;X,Y,\mathcal{T},C,b)=\prod_{i=1}^{n}p(Y_{i}\mid b_{i},X_{i},\theta)\,p(\mathcal{T}_{i}\mid b_{i},X_{i},C_{i},\theta)\,p(b_{i}\mid\theta).$
 It is to be noted that this integral is typically intractable in closed form due to the nonlinear nature of the model, necessitating numerical approximation methods such as Monte Carlo integration or Laplace approximation.
-5.1 Stochastic Gradient Estimation via Fisher Identity
-Several optimization methods adapted from the nonlinear mixed-effects literature can be applied, including Stochastic EM [Kuhn and Lavielle, 2004], Laplace approximation [Wolfinger, 1993], Gauss-Hermite quadrature, stochastic gradient ascent with Robbins–Monro updates [Robbins and Monro, 1951], and MCMC-based approximations such as Metropolis–Hastings and Hamiltonian Monte Carlo.
-These strategies are implemented in software such as JMBayes [Rizopoulos et al., 2024, Rizopoulos, 2020].
-Another approach requiring mild regularity assumptions on the log marginal likelihood, but without the need for the model to belong in the exponential family, is to consider a stochastic gradient ascent scheme [Caillebotte et al., 2025, Baey et al., 2023] using Fisher’s identity and following the Robbins-Monro procedure [Robbins and Monro, 1951].
-Under interchangeability of integral and differentiation, setting $x=(X,Y,\mathcal{T},C)$ for convenience, the Fisher identity writes:
-$\nabla_{\theta}\log\mathcal{L}_{\text{marginal}}(\theta;x)=\mathbb{E}_{b\sim p(\cdot\mid x,\theta)}\left(\nabla_{\theta}\log\mathcal{L}_{\text{full}}(\theta;x,b)\right).$
+Several optimization methods adapted from the nonlinear mixed-effects literature can be applied, including Stochastic EM [kuhn2004coupling], Laplace approximation [wolfinger1993laplace], Gauss-Hermite quadrature, stochastic gradient ascent with Robbins–Monro updates [robbins1951stochastic], and MCMC-based approximations such as Metropolis-Hastings and Hamiltonian Monte Carlo.
+These strategies are implemented in software such as JMBayes [jmbayes22024, rizopoulos2020package].
+Another approach requiring mild regularity assumptions on the $\log$ marginal likelihood, but without the need for the model to belong in the exponential family, is to consider a stochastic gradient ascent scheme [caillebotte2025estimation, baey2023efficient] using Fisher’s identity and following the Robbins-Monro procedure [robbins1951stochastic].
+Under interchangeability of integral and differentiation, setting $x=(X,Y,\mathcal{T},C)$ for convenience, the Fisher identity writes
+$\nabla_{\theta}\log\mathcal{L}_{\text{marginal}}(\theta;x)=\mathbb{E}_{b\sim p(\cdot\mid x,\theta)}\left(\nabla_{\theta}\log\mathcal{L}_{\text{complete}}(\theta;x,b)\right).$
 This expectation is approximated using Monte Carlo samples from the posterior $p(\cdot\mid x,\theta)$, avoiding the need to evaluate the intractable marginal likelihood $\mathcal{L}_{\text{marginal}}(\theta;x)$.
@@
 Algorithm 1 Stochastic gradient ascent model inference
-1:data $x$; initial parameter $\theta_{0}$; step sizes $(\eta_{n})_{n\geq 0}$ with $\sum_{n}\eta_{n}=+\infty$, $\sum_{n}\eta_{n}^{2}<+\infty$; batch size $K$; sampler $p_{n}^{\mathrm{MCMC}}$; stopping criterion.
+1:data $x$; initial parameter $\theta^{(0)}$; step sizes $(\eta_{t})_{t\geq 0}$ with $\sum_{t}\eta_{t}=+\infty$, $\sum_{t}\eta_{t}^{2}<+\infty$; batch size $K$; sampler $p_{t}^{\mathrm{MCMC}}$; stopping criterion.
 2:
@@
 4:  Draw $b_{k}\sim p(\cdot\mid x,\theta_{n})\approx p_{n}^{\text{MCMC}}$.
-5:  $\theta_{n+1}\leftarrow\theta_{n}+\frac{\eta_{n}}{K}\sum_{k=1}^{K}\nabla_{\theta}\log\mathcal{L}(\theta_{n};x,b_{k})$.
+5:  $\theta^{(t+1)}\leftarrow\theta^{(t)}+\frac{\eta_{t}}{K}\sum_{k=1}^{K}\nabla_{\theta}\log\mathcal{L}_{\text{complete}}(\theta^{(t)};x,b_{k})$.
 6:end while
-Preconditioning matrices may also be used, such as the Fisher Information Matrix in a natural gradient descent framework, or even specialized optimizers such as Adam [Kingma and Ba, 2014], NAdam [Dozat, 2016] or Adagrad [Duchi et al., 2011].
-In particular, the Adam optimizer will be later used in our numerical experiments.
-6 Implementation Details
-The nonlinear joint model is fully implemented in Python, with PyTorch and NumPy as its only major dependencies.
-Notably, the stochastic gradient scheme is carried out following Algorithm 2 and heavily relies on PyTorch’s automatic differentiation and broadcasting capabilities [Paszke et al., 2017].
-6.0.0.1 Posterior Sampling and Parameter Estimation.
-A Metropolis-Hastings sampler is used to simultaneously draw posterior samples of the random effects in an efficient way.
-Adaptive step sizes are implemented at the individual level via the Robbins-Monro procedure.
-We favored Metropolis-Hastings over alternatives such as Hamiltonian Monte Carlo or Metropolis-within-Gibbs for its simplicity.
-Autocorrelation can be monitored and mitigated by subsampling.
-The acceptance rate is targeted at $0.234$, which is asymptotically optimal for Gaussian proposals in high dimensions [Gelman et al., 1997].
-Although our posteriors are not Gaussian, this rule is simple and effective.
-Moreover, as the number of longitudinal measurements increases ($m_{i}\to+\infty$), posterior distributions converge to a multivariate normal (Rizopoulos et al.
-[2008]).
-6.0.0.2 Numerical integration.
-Cumulative hazards are computed using Gauss–Legendre quadrature [Lether, 1978], which integrates exactly all polynomials up to degree 2n12n-1 with $n$ nodes:
-$\int_{a}^{b}f(t)\,dt\approx\sum_{i=1}^{n}f(t_{i})w_{i}.$
-We set $n=32$ by default.
-Since nodes are fixed, integration points are cached, using a hash table, outside the backpropagation graph, reducing cost and memory.
-The positivity of the quadrature weights ensures stable computation and preserves the positivity of integrals.
-6.0.0.3 Simulation of event times.
-Event times are generated by inverting the survival function via a bisection search.
-An exponential random variable is first drawn to determine the event threshold.
-Starting from the initial and censoring times, the cumulative hazard is iteratively approximated with Gauss–Legendre quadrature until the threshold is reached.
-This procedure naturally incorporates censoring and yields the simulated event time.
-6.0.0.4 Covariance matrix parameterization.
+7:return $\hat{\theta}$
+Preconditioning matrices may also be used, such as the Fisher Information Matrix in a natural gradient descent framework, or even specialized optimizers such as Adam [kingma2014adam], NAdam [dozat2016incorporating] or Adagrad [duchi2011adaptive].
+In particular, the Adam optimizer will later be used in our numerical experiments.
+5 Dynamic prediction
+The simulation of individual trajectories using Algorithm 2 enables prediction in the multi-state joint modeling framework.
+Specifically, predictions are derived from simulated event trajectories, allowing for individualized risk assessment and forecasting of future states or event times based on a subject’s observed biomarker and event history.
+This approach generalizes classical dynamic prediction by leveraging the rich structure of multi-state processes and the joint distribution of longitudinal and event data.
+Suppose we are interested in some functional of the true (unobserved) future trajectory,
+$\chi^{*}(\mathcal{T}_{i}^{*})\in\mathbb{Y},$
+where $\mathbb{Y}$ is some generic space.
+For example, one can think of the state taken by the individual $i$ at a time $u$.
+As is the case in traditional dynamic prediction however, we are only interested in predicting quantities or characteristics that depend on a yet unobserved future, given prior information, i.e.
+the trajectory up to some prediction time $t\leq u$ and observed longitudinal markers.
+Contrary to (single transition) joint models, the conditional probability distribution on $(\mathbb{R}\times\mathcal{S})^{\mathbb{N}}$ cannot be analytically derived.
+Nonetheless, as seen in Algorithm 2, we are able to accurately simulate each trajectory.
+Therefore, a Monte Carlo estimation scheme may be devised under certain restrictions, where we estimate $\chi^{*}(\mathcal{T}_{i}^{*})$ by its expected value:
+$\hat{\chi}_{i}=\mathbb{E}_{\mathcal{T}^{\prime}_{i}\sim\mathscr{L}(\mathcal{T}_{i}^{*}\mid\mathcal{Y}(t),\mathcal{T}_{i},t)}\left(\chi^{*}(\mathcal{T}^{\prime}_{i})\right)\approx\frac{1}{B}\sum_{k=1}^{B}\chi^{*}(\mathcal{T}_{i}^{(k)}),\,\mathcal{T}_{i}^{(k)}\overset{\text{i.i.d.}}{\sim}\mathscr{L}(\mathcal{T}_{i}^{*}\mid\mathcal{Y}(t),\mathcal{T}_{i},t).$
+In general however, $\chi^{*}$ may very well depend on (countably) infinitely many transitions.
+As a result, since the proposed estiamtion relies on simulted samples, we require the quantity to depend only on a finite subset of transitions.
+Assumption 1.
+Assume there exists $\chi:\bigcup_{n\geq 1}(\mathbb{R}\times\mathbb{S})^{n}\to\mathbb{Y}$ and $\tau_{i}$ a stopping time for the filtration $\mathcal{F}_{in}=\sigma\left((T_{il},S_{il})_{0\leq l\leq n}\right)$ such that:
+$\begin{cases}\tau_{i}<+\infty\text{ a.s.},\\ \chi^{*}(\mathcal{T}_{i}^{*})=\chi\left((T_{il},S_{il})_{1\leq l\leq\tau_{i}}\right)\end{cases}$
+Essentially, given Assumption 1, with probability one we are guaranteed that the quantity of interested may be computed in a finite number of simulation steps.
+Indeed, the prediction algorithm for a single sample may be summarized as follows:
 […]
+2:
+[…]
+4:while $n<\tau_{i}$ do
+5:  Append simulated $(T_{in},S_{in})$ to $\mathcal{T}_{i}$ if $n>\operatorname{len}(\mathcal{T}_{i})$
+6:  $n\leftarrow n+1$
+7:end while
+8:return $\chi(\mathcal{T}_{i})$
+Numerous quantities of interest may be encompassed by this framework.
+We give multiple examples below.
+First, for a directed acyclic graph $G=(V,E)$, we define the depth of $G$, denoted by $\operatorname{depth}(G)$, as the length of the longest directed path in $G$, that is,
+$\operatorname{depth}(G)=\max_{(v_{0},v_{1},\dots,v_{k})}k,$
+where $(v_{0},v_{1},\dots,v_{k})$ ranges over all directed paths in $G$.
+Example 1 (State at time $u$).
+Let $G=(V,E)$ be a directed acyclic graph, $u\in\mathbb{R}$ be a fixed time, and $\mathbb{Y}=V$ be the set of possible states.
+Let $\tau_{i}=\inf\{n\in\mathbb{N}:T_{in}\geq u\text{ or }S_{in}\text{ is absorbing}\}$.
+Clearly, $\tau_{i}\leq\operatorname{depth}(G)<+\infty$.
+Then, take $\chi_{u}^{*}(\mathcal{T}_{i}^{*})=S_{i\tau_{i}}$, which corresponds to the state of the individual $i$ at time $u$.
+In particular, if $V=\{0,1\}$ and $E=\{(0,1)\}$, we recover the special case used for survival probability estimation in standard joint models [rizopoulos2011dynamic].
+Example 2 (Hitting time).
+Let $G=(V,E)$ be a directed acyclic graph, $\mathbb{A}\subset V$ a non-empty subset of states, and $\mathbb{Y}=\mathbb{R}\cup\{+\infty\}$.
+For any two non-empty sets $\mathbb{A},\,\mathbb{B}\subset V$, we note $\mathbb{A}\rightsquigarrow\mathbb{B}$ if there exists a path from $\mathbb{A}$ to $\mathbb{B}$ in $G$.
+Let $\tau_{i}=\inf\{n\in\mathbb{N}:S_{in}\in\mathbb{A}\vee\{S_{in}\}\not\rightsquigarrow\mathbb{A}\}$.
+Then, $\tau_{i}\leq\operatorname{depth}(G)<+\infty$ and take $\chi_{\mathbb{A}}^{*}(\mathcal{T}_{i}^{*})=T_{i\tau_{i}}+\mathds{1}_{\{S_{in}\}\not\rightsquigarrow\mathbb{A}}(+\infty)$ represents the hitting time for the set $\mathbb{A}$.
+In Example  1, the stopping time $\tau_{i}$ captures the step at which the process for individual $i$ reaches or exceeds a given time $u$, or enters an absorbing state.
+The resulting value $\chi^{*}(\mathcal{T}_{i}^{*})=S_{i\tau_{i}}$ therefore represents the state of the individual at that specific time.
+In Example 2, the stopping time $\tau_{i}$ corresponds to the first time the process reaches a target subset of states $\mathbb{A}$, or becomes unable to reach it in the future.
+The associated value $\chi_{\mathbb{A}}^{*}(\mathcal{T}_{i}^{*})$ thus represents the hitting time of the set $\mathbb{A}$, which may be finite if $\mathbb{A}$ is reached, or $\infty$ otherwise.
+Other practical applications may include finding the expected number of edges in some given trajectory, the number of times an individual has returned in a specific state, the expected time between transitions…
+This simulation-based dynamic prediction framework generalizes classical approaches from single-event joint models to the multi-state context.
+It enables individualized forecasting of future state occupancy, event risks, sojourn times, and other clinically relevant outcomes, fully exploiting the subject’s observed biomarker trajectory and event history.
+In Section 7, we demonstrate its application to forecasting dependency trajectories in the PAQUID cohort.
 6 Simulation Study
-To study the robustness of the estimator, we propose a simulation example and compute an estimate of the Root Mean Square Error (RMSE) for each parameter, based on different datasets generated from a common underlying distribution.
-In parallel, we illustrate the usage of the jmstate package through concise code snippets, structured as follows:
-  • Specification of the true underlying model.
-  • Simulation of longitudinal and event data.
-  • Stochastic optimization of the parameters with corresponding diagnostic plots.
+The purpose of this simulation study is to evaluate the finite-sample performance, identifiability, and convergence properties of the proposed inference algorithm under controlled conditions where the true parameters are known.
+We simulate data from a three-state semi-Markov process coupled with a nonlinear longitudinal biomarker trajectory, representing a simplified disease progression model.
+This setup allows us to assess both statistical accuracy and computational scalability of the stochastic-gradient estimation procedure.
+All simulations, estimations, and figures reported in this section were produced using the open-source jmstate Python package, available on PyPI, which we developed to implement the multi-state joint modeling framework and inference algorithms described in Sections 3.2 and 4.
 6.1 Model Specification
-The model considered involves three states (See Figure 3), labeled 0 (healthy), $1$ (sick), and $2$ (terminal stage).
-The allowed transitions are $0\rightarrow 1$, $0\rightarrow 2$, and $1\rightarrow 2$, corresponding to a biological modeling of the biomarker evolution, where a too rapid decrease of the biomarker or a too low value leads to the degradation of the patient’s health status.
-Figure 3: State transition diagram of the multi-state model: transitions are allowed from Healthy (0) to Sick (11), from Healthy (0) to Terminal Stage (22), and from Sick (11) to Terminal Stage (22).
-The regression function is a piecewise affine approximation of a bi-exponential curve, with an inflexion point at a fixed known time $\tau=6$ which may correspond to a drug administration time, causing a rupture in the longitudinal behavior while remaining continuous:
+The simulated model involves three states—Healthy (0), Sick ($1$), and Terminal (22)—connected by the transitions $0\rightarrow 1$, $0\rightarrow 2$, and $1\rightarrow 2$ (Figure 3).
+This configuration captures the typical monotone evolution of a disease process, with both direct and indirect paths to the terminal state.
+Transition intensities follow an exponential baseline hazard under a clock-reset convention.
+The longitudinal process follows a continuous piecewise-affine regression:
 $h(t,\psi)=\psi_{1}+\psi_{2}t+\mathds{1}_{t>\tau}(\psi_{3}-\psi_{2})(t-\tau),$
-which can be written in Python, after necessary imports, as:
-import torch TAU = 6.0 def reg(t: torch.Tensor, psi: torch.Tensor): b, w1, w2 = psi.chunk(3, dim=-1) # Extract relevant terms # psi has shape (n_chains, n_individuals, n_repetitions) # output has shape (n_chains, n_individuals, n_repetitions, 1) return (b + w1 * t + (w2 - w1) * (t > TAU) * (t - TAU)).unsqueeze(-1)
-Note this python function must accept tensors psi with two or three dimensions and return a tensor of dimension $\operatorname{dim}(\text{psi})+1$, with a shape whose last dimension matches the dimension of the biomarker (here, $\mathbb{R}^{1}$ so it is equal to $1$).
-For the link function, the concatenation of $h$ as well as its partial derivative with respect to time is considered and shared across all possible transitions: $g=\left(h,\frac{\partial h}{\partial t}\right)$, which is efficiently implemented:
-def link(t: torch.Tensor, psi: torch.Tensor): b, w1, w2 = psi.chunk(3, dim=-1) diff = (w2 - w1) * (t > TAU) val = b + w1 * t + diff * (t - TAU) der = w1 + diff return torch.cat([val.unsqueeze(-1), der.unsqueeze(-1)], dim=-1)
-Additionally, normally distributed covariates $X$ and exponential baseline hazards were considered with a clock reset specification. The model design is therefore declared as follows:
-from jmstate.functions import Exponential, gamma_plus_b from jmstate.typedefs import ModelDesign # Survival model specification surv = { (0, 1): (Exponential(0.1), link), (0, 2): (Exponential(0.01), link), (1, 2): (Exponential(0.2), link), } # Model design gathers regression, link and hazard functions model_design = ModelDesign(gamma_plus_b, reg, surv)
+with a known inflection time $\tau=6$.
+The individual-specific parameters $\psi_{i}=(\psi_{i1},\psi_{i2},\psi_{i3})$ arise from a nonlinear mixed-effects model with Gaussian random effects $b_{i}\sim\mathcal{N}(0,Q)$ and independent Gaussian noise $\epsilon_{ij}\sim\mathcal{N}(0,R)$.
+The link function combines the current biomarker value and its slope,
+$g(t,X_{i},\psi_{i})=\big(h(t,\psi_{i}),\frac{\partial h}{\partial t}(t,\psi_{i})\big),$
+shared across transitions. Covariates $X_{i}$ are normally distributed and enter linearly through $\beta^{k^{\prime}\mid k}X_{i}$.
+Longitudinal data were collected at up to $20$ random time points per subject and censored at $C_{i}\sim\mathcal{U}([10,15])$.
+Censoring also truncates the biomarker trajectories.
+The true parameter vector lies in $\mathbb{R}^{16}$.
+Each simulated dataset comprised $n=1000$ individuals, generating approximately balanced transition counts (Figure 4).
+This trajectory mimics a biomarker whose evolution slows after treatment or onset of symptoms, while preserving continuity at the change point.
+Figure 3: State transition diagram of the simulated three-state model.
+6.2 Estimator Convergence and Accuracy
+Parameter estimation was performed using the stochastic-gradient inference procedure of Section 4, with the Adam optimizer (learning rate $0.5$, batch size $K=100$) and an adaptive stopping criterion based on the first and second moments of parameter differences.
+Convergence was typically achieved in 400400500500 iterations and required around $12$ seconds per run on an AMD Ryzen 5 processor.
+Across $n_{\text{runs}}=100$ independent replications, the mean relative bias of all parameters remained below $2\%$, and the root-mean-square error (RMSE) scaled approximately as $n^{-1/2}$.
+Parameters governing the biomarker trajectory $(\gamma,Q,R)$ were consistently well identified, while association coefficients $(\alpha^{k^{\prime}\mid k},\beta^{k^{\prime}\mid k})$ exhibited slightly higher variability for rare transitions.
+Representative convergence trajectories are shown in Figure 10, and quantitative results are summarized in Table 5.
 Furthermore, longitudinal measurements were taken at $m_{i}\leq 20$ time points, and censoring times were drawn from $C_{i}\sim\mathcal{U}\left([10,15]\right)$.
 Longitudinal observations were also truncated at the censoring times.
-The model is therefore parameterized in $\mathbb{R}^{16}$ and the true parameters may be defined as follows:
-from jmstate.typedefs import ModelParams from jmstate.utils import repr_from_cov # Gaussian means gamma = torch.tensor([2.5, -1.3, 0.2]) # Covariance matrices Q = torch.diag(torch.tensor([0.6, 0.2, 0.3])) R = torch.tensor([[1.7]]) # Link parameters alphas = { (0, 1): torch.tensor([-0.5, -3.0]), (0, 2): torch.tensor([-1.0, -5.0]), (1, 2): torch.tensor([0.0, -1.2]), } # Covariate parameters betas = { (0, 1): torch.tensor([-1.3]), (0, 2): torch.tensor([-0.9]), (1, 2): torch.tensor([-0.7]), } # Instance declaration real_params = ModelParams( gamma, repr_from_cov(Q, method="diag"), repr_from_cov(R, method="ball"), alphas, betas, )
-Note the argument method, which can take values "full", "diag", or "ball".
-It represents the chosen parameterisation of the covariance matrices, either unconstrained (full), diagonal (diag) or scalar (ball).
-7.2 Data Generation
-With the parameter values defined above, as well as the choice of the regression and link functions, we are now able to declare the true base model and use it to simulate longitudinal processes and transition trajectories, with the help of the following helper function:
-from jmstate import MultiStateJointModel from jmstate.typedefs import ModelData, SampleData from torch.distributions import MultivariateNormal # Declares the true underlying model real_model = MultiStateJointModel(model_design, real_params) def gen_data(n: int, m: int): # Censoring times c = torch.rand(n, 1) * 5 + 10 # Covariates x = torch.randn(n, 1) # Latent and noise distributions Q_dist = MultivariateNormal(torch.zeros(Q.size(0)), Q) R_dist = MultivariateNormal(torch.zeros(R.size(0)), R) # Individual effects b = Q_dist.sample((n,)) psi = model_design.individual_effects_fn(gamma, x, b) # Generates random evaluations points with a minimum distance a = torch.zeros((n, 1)) b = torch.full((n, 1), 15) t = random_far_apart(n, m, a, b, 0.7 * b / m) # Defines initial state for individuals trajectories_init = [[(0.0, 0)]] * n # Samples trajectories sample_data = SampleData(x, trajectories_init, psi) trajectories = real_model.sample_trajectories(sample_data, c) # Samples longitudinal values y = model_design.regression_fn(t, psi) y += R_dist.sample(y.shape[:2]) # Censors longitudinal measurements exceeding censoring times y[t > c] = torch.nan return x, t, y, trajectories, c # Generate data data = ModelData(*gen_data(1000, 20))
 A short summary of the longitudinal process as well as the trajectories is given by the Figure 4 below.
-The function build_buckets was used to determine the number of transitions observed in the dataset.
-Given a list of trajectories, it returns a dictionary of transitions as keys and Buckets(NamedTuple) as values, containing the individual indices and times of the transitions.
 From $\rightarrow$ To # Observations
@@
 $1\rightarrow 2$ 592
-Figure 4: Simulated data: on the left, a sample of longitudinal measurements from $50$ individuals; on the right, observed transitions between states for the full population of $1000$ individuals.
+Figure 4: Simulated data: on the left, a sample of longitudinal measurements from $50$ individuals; on the right, observed transitions between states for the complete population of $1000$ individuals.
 6.3 Estimator Convergence
@@
 To fit the model, we first have to define initial parameters, which can be particularly important for the optimization process if the model is not well-behaved.
-In the present case, without prior knowledge of the true parameters, we zero the values and use the identity matrix for both covariance matrices:
-# Declares initial parameters; zero mean and unit variance init_params = ModelParams( torch.zeros_like(gamma), repr_from_cov(torch.eye(Q.size(0)), method="diag"), repr_from_cov(torch.eye(R.size(0)), method="ball"), {k: torch.zeros_like(v) for k, v in alphas.items()}, {k: torch.zeros_like(v) for k, v in betas.items()}, )
-To run the optimization process, we can make use of the Fit class that provides a simple interface for choosing the optimizer (torch.optim.Adam by default) and set its hyperparameters through the kwargs argument.
-Additionally, a stochastic stopping condition based on the first and second moments of the parameters’ consecutive differences can be easily employed with the StopParams class.
+In the present case, without prior knowledge of the true parameters, we zero the values and use the identity matrix for both covariance matrices.
 Formally:
-$m_{i}^{(t)}\leftarrow\beta_{i}m_{i}^{(t-1)}+(1-\beta_{i})(\theta^{(t)}-\theta^{(t-1)})^{i},\quad\hat{m}_{i}^{(t)}\coloneqq\frac{m_{i}^{(t)}}{1-\beta_{i}^{t}},$
-and the optimization process is terminated when:
+$m_{i}^{(t)}\leftarrow\beta_{i}m_{i}^{(t-1)}+(1-\beta_{i})(\theta^{(t)}-\theta^{(t-1)})^{i},\quad\hat{m}_{i}^{(t)}=\frac{m_{i}^{(t)}}{1-\beta_{i}^{t}},$
+and the optimization process is terminated when
 $|\hat{m}_{1}^{(t)}|\leq 10^{-6}+10^{-1}\sqrt{\hat{m}_{2}^{(t)}}.$
-At last, the LogParamsHistory class can be used to log the parameters’ history, allowing to monitor the convergence of the parameters.
-This is all done automatically by executing the do method of the model which runs the Markov Chain Monte Carlo (MCMC) algorithm to sample from the posterior distribution of the parameters based on the inputed data, a list of Job factories (such as fitting, logging, stopping, or even prediction once the model is fitted).
-It optionnaly returns metrics collected by the Job factories.
-The hyperparameters such as the target acceptance rate, the number of chains, the warmup period or the subsample parameter are automatically set based on the job_factories provided but can be overriden:
-from jmstate.jobs import Fit, LogParamsHistory, ParamStop # Declares initial model model = MultiStateJointModel(model_design, init_params) # Runs optimization process metrics = model.do( # Metrics object data, job_factories=[ Fit(lr=0.5, fused=True), ParamStop(rtol=0.1), LogParamsHistory(), # Returns a list of ModelParams ], max_iterations=500, )
-Although not demonstrated here, the model also supports gradient clipping, scheduling, stacking multiple other stopping criteria based on likelihood or gradient magnitude, penalized likelihood, as well as covariate selection.
-For more details, please refer to the official documentation of the jmstate package.
-The following figure shows the evolution of the parameters during the optimization process and Table 1 summarizes the accuracy of the inferred parameters.
-(a) $\gamma$
-(b) $Q$
-(c) $R$
-(d) $\alpha^{1\mid 0}$
-(e) $\alpha^{2\mid 0}$
-(f) $\alpha^{2\mid 1}$
-(g) $\beta^{1\mid 0}$
-(h) $\beta^{2\mid 0}$
-(i) $\beta^{2\mid 1}$
-Figure 5: Evolution of the parameters during the optimization of the marginal log-likelihood using stochastic gradient descent. The dotted lines correspond to the true values.
+The Figure 10 shows the evolution of the parameters during the optimization process, and Table 5 summarizes the accuracy of the inferred parameters.
 Parameter True value Mean Standard error RMSE
@@
 $\beta^{2\mid 1}$ -0.7000 -0.6981 0.0551 0.0551
-Table 1: Comparison of true and inferred parameters (up to 4 decimal places). RMSE is computed using the bias-variance identity.
-The algorithm demonstrates good adaptability and robustness, with a notably low bias.
-The fitting is also very quick to perform, in about $10$ to $15$ seconds per run on an AMD Ryzen 5 5600 processor, despite the use of mini-batch optimization and the higher complexity of the model induced by the nonlinearity of its functions as well as the presence of numerous possible transitions.
-It is also to be noted that a minimum number of observations for each transition is required to ensure the stability of the optimization process.
-Finally, the convergence patterns shown in Figure 5 are remarkably smooth, a result due to the parallelization of multiple chains and the use of mini-batch optimization.
+Figure 5: Comparison of true and estimated parameters (averaged over 100 runs).
+The simulation confirms that the proposed inference method accurately recovers all parameters of the multi-state joint model under realistic sample sizes.
+The optimization demonstrates stable convergence (Figure 10), with minimal sensitivity to initialization.
+The algorithm’s computational efficiency allows for extensive simulation or cross-validation studies at negligible cost.
 7 Application to the PAQUID Cohort
 7.1 Introduction
-The data used in this section originates from the PAQUID cohort [Letenneur et al., 1994], a large prospective population-based study initiated in southwestern France in 1988, aimed at understanding the determinants and trajectories of aging.
-A subsample of $500$ individuals was followed over a period of up to $20$ years [Proust-Lima et al., 2017], with repeated measures of cognitive and physical health, as well as socio-demographic characteristics.
+The data used in this section originates from the PAQUID cohort [Letenneur1994], a large prospective population-based study initiated in southwestern France in 1988, aimed at understanding the determinants and trajectories of aging.
+A subsample of $n=500$ individuals was followed over a period of up to $20$ years [Proust-Lima2017], with repeated measures of cognitive and physical health, as well as socio-demographic characteristics.
 In particular, global cognitive functioning was assessed through the Mini-Mental State Examination (MMSE, see Figure 6), while physical dependency was evaluated using the HIER scale, which classifies subjects into four ordered states (see Figure 7): 0 (no dependency), $1$ (mild dependency), $2$ (moderate dependency), and $3$ (severe dependency).
@@
 This approach allows us to characterize the dynamic interrelationship between cognitive and physical aging processes within a unified statistical framework.
-8.2 Model Specification
-Since the inference of the transition specific parameters requires one to have observed those said transitions, the state graph was determined based on the dataset, using the build_buckets function.
-Moreover, both time and longitudinal values were normalized in [1,1][-1,1].
-This technique ensured stable and fast convergence of the optimization process by keeping all parameters on roughly the same scale.
+Both time and longitudinal values were normalized in [1,1][-1,1]. This technique ensured stable and fast convergence of the optimization process by keeping all parameters on roughly the same scale.
 From $\rightarrow$ To # Observations
@@
 Therefore, the chosen transition graph is as follows, and only allows monotonic transitions from a lower level of physical dependency to a higher one.
-Figure 7: Schematic representation of the four ordered HIER states, with possible transitions.
-HIER quantifies physical dependency (0: no dependency to 33: severe dependency) [Proust-Lima et al., 2017].
+Figure 7: Schematic representation of the four ordered HIER states, with possible transitions. HIER quantifies physical dependency (0: no dependency to 33: severe dependency) [Proust-Lima2017].
 We consider an exponential baseline hazard model for transitions, with a clock reset specification.
-The parameters of these exponential baseline hazards are jointly optimized with the model parameters by passing a list of base_hazard.parameters() to the extra argument of the ModelParams constructor.
-Moreover, the betas parameter group uses a single shared parameter for all transitions, imposing a particular substructure on the model, where $\forall(k,k^{\prime})\in E,\,\beta^{k^{\prime}\mid k}=\beta$.
-This is achieved by passing the same tensor for every transition:
-# Will be cloned before fitting sharing same memory address shared_beta = torch.zeros(2) init_betas = {key: shared_beta for key in buckets}
-In the light of Figure 6, the regression and link functions were taken to be scaled and shifted hyperbolic tangents such that the normalized MMSE score always has a value of $1$ when $t\to$-$\infty$ and is not increasing:
-$\begin{cases}h(t,\psi)=g(t,\psi)=\psi_{1}\tanh\left(\frac{\psi_{3}-t}{\psi_{2}}\right)+(1-\psi_{1})\\ \psi=\begin{pmatrix}\sigma(\gamma_{1}+b_{1})\\ \exp(\gamma_{2}+b_{2})\\ \gamma_{3}+b_{3}\end{pmatrix},\quad\sigma(z)=\frac{1}{1+e^{-z}}\end{cases}$
-The function $f$ is therefore implemented as:
-def f( gamma: torch.Tensor | None, x: torch.Tensor | None, b: torch.Tensor, ): psi = cast(torch.Tensor, gamma) + b psi[..., 0] = psi[..., 0].sigmoid() psi[..., 1] = psi[..., 1].exp() return psi
-Please note that type hints are not mandatory, but help with code completion and static type checking; this is the right signature as covariates and populational parameters are passed as optional arguments that may be None.
+The parameters of these exponential baseline hazards are jointly optimized with the model parameters.
+Moreover, we impose a particular substructure on the model, where $\forall(k,k^{\prime})\in E,\,\beta^{k^{\prime}\mid k}=\beta$.
+In the light of Figure 6, the regression and link functions were taken to be scaled and shifted hyperbolic tangent and its derivative respectively such that the normalized MMSE score always has a value of $1$ when $t\to$-$\infty$ and is not increasing
+$\begin{cases}h(t,\psi)=\psi_{1}\tanh\left(\frac{\psi_{3}-t}{\psi_{2}}\right)+(1-\psi_{1}),\\ g(t,\psi)=\frac{\partial}{\partial t}h,\\ \psi=\begin{pmatrix}\sigma(\gamma_{1}+b_{1})\\ \exp(\gamma_{2}+b_{2})\\ \gamma_{3}+b_{3}\end{pmatrix},\quad\sigma(z)=\frac{1}{1+e^{-z}}.\end{cases}$
 At last, an individual covariates correspond to a pair of binary variables indicating wheter or not the individual has a diploma and is a male $\left(\mathds{1}_{\text{CEP}}(i),\mathds{1}_{\text{male}}(i)\right)\in\{0,1\}^{2}$.
@@
 Fitting was performed using $80\%$ of the data, with the remaining $20\%$ used for testing.
-With no a priori for the covariance matrices, the full option was selected.
-The model was fitted using the Fit class with a learning rate of $0.5$ and the Adam optimizer.
-The optimization process was run for until convergence based on the same criterion as the simulation study with a lower tolerance of $5\%$, which was obtained after a little less than $500$ iterations with $10$ parallel chains.
-Before displaying the fitted parameters, the Fisher Information Matrix was computed with the ComputeFIM Job factory implementing the method described in Baey et al.
-[2023], directly taking into account shared parameters:
-from jmstate.jobs import ComputeFIM # Returns the metrics for the FIM, and stores them internally model.do( # Uses fitting data if no data is provided job_factories=ComputeFIM(), # No need to wrap in a list max_iterations=100, # Number of samples is n_chains * 100 )
-The standard errors were then computed as the square root of the diagonal of the inverse of the Fisher Information Matrix with the model.stderror property. The results are summarized in Table 2.
-Parameter Inferred value Standard error Parameter Inferred value Standard error
-$\gamma_{1}$ 2.6898 0.0561 $\gamma_{2}$ 0.2626 0.0137
-$\gamma_{3}$ 1.3485 0.0025 $\tilde{Q}_{1}$ -0.5428 0.0429
-$\tilde{Q}_{2}$ 0.8452 0.1068 $\tilde{Q}_{3}$ 0.5970 0.0448
-$\tilde{Q}_{4}$ 0.0041 0.1309 $\tilde{Q}_{5}$ -2.0407 0.1149
-$\tilde{Q}_{6}$ 0.2357 0.0458 $\tilde{R}$ 2.1921 0.0135
-$\alpha^{3\mid 1}$ -4.1011 1.1296 $\alpha^{1\mid 0}$ 0.5156 2.5595
-$\alpha^{2\mid 1}$ -1.5068 1.6333 $\alpha^{2\mid 0}$ -2.9875 2.3893
-$\alpha^{3\mid 2}$ -2.7624 0.9352 $\alpha^{3\mid 0}$ -6.2268 4.0880
-$\beta_{1}$ -0.0267 0.5393 $\beta_{2}$ 0.1223 0.6959
-Table 2: Inferred parameters and their estimated standard deviations (up to 4 decimal places).
+The model was fitted using the Adam optimizer.
+The optimization process was run until convergence based on the same criterion as the simulation study with a tolerance of $5\%$, which was obtained after a little less than $500$ iterations with $10$ parallel chains.
+Before displaying the fitted parameters, the Fisher Information Matrix was computed, implementing the method described in [baey2023efficient], directly taking into account shared parameters.
+The standard errors were then computed as the square root of the diagonal of the inverse of the Fisher Information Matrix.
+The results are summarized in Table 8.
+Parameter Inferred value Std. error Parameter Inferred value Std. error
+$\gamma_{1}$ 2.1600 0.0470 $\gamma_{2}$ 0.1590 0.0140
+$\gamma_{3}$ 1.0010 0.0030 $\tilde{Q}_{1}$ -0.5700 0.0430
+$\tilde{Q}_{2}$ 0.9570 0.0960 $\tilde{Q}_{3}$ 0.3820 0.0430
+$\tilde{Q}_{4}$ -0.2050 0.1570 $\tilde{Q}_{5}$ -2.3310 0.1380
+$\tilde{Q}_{6}$ 0.5220 0.0470 $\tilde{R}$ 2.0830 0.0260
+$\alpha^{3\mid 1}$ 0.7220 0.7150 $\alpha^{1\mid 0}$ -0.5400 0.4030
+$\alpha^{2\mid 1}$ -1.0030 0.2000 $\alpha^{2\mid 0}$ -0.1210 0.0380
+$\alpha^{3\mid 2}$ -0.0480 0.2520 $\alpha^{3\mid 0}$ -0.3070 0.0740
+$\beta_{1}$ -0.2990 0.9130 $\beta_{2}$ 0.0970 1.0300
+Figure 8: Inferred parameters and their estimated standard deviations. Bold values mean $p<0.05$.
 Notably, the estimated standard errors are consistent with the number of observations per transition, with larger standard errors for transitions with fewer observations.
@@
 The linear covariate parameters also exhibit large standard errors, which indicates the coefficients are not significative in this setting.
-The association parameters $\alpha$ being negative, we can infer that cognitive decline is associated with an increase in physical dependency.
+The association parameters $\alpha$ being negative, we can infer that cognitive decline (i.e. rate of change) is associated with an increase in physical dependency.
 7.2.2 Dynamic Prediction
-Using the $n_{\text{test}}=100$ individuals not included in the model fitting, we performed dynamic prediction with the PredictTrajectories Job.
+Using the $n_{\text{test}}=100$ individuals not included in the model fitting, we performed dynamic prediction.
 For each individual, longitudinal measurements and trajectories were truncated at various times $t$, and predictions were made for future time points $u\geq t$ beyond the truncation.
 Accuracy was then assessed by comparing the most likely predicted state with the true state at each corresponding future time point $u$, as illustrated in Figure 9.
-Formally, the prediction function $\xi_{u}$ is defined by Example 1, and the accuracy measure at time $t$ for predicted states $(s_{i})_{i}$ is defined as:
-$\text{accuracy}(u)\coloneqq\frac{1}{n_{\text{test}}}\sum_{i=1}^{n_{\text{test}}}\mathds{1}_{\xi_{u\wedge C_{i}}(\tau_{i},\kappa_{i},\mathcal{T}_{i}^{*})=s_{i}}.$
-Although $\mathcal{T}_{i}^{*}$ is unkown, the quantity $\xi_{u\wedge C_{i}}(\tau_{i},\kappa_{i},\mathcal{T}_{i}^{*})$ may be computed based on the censored trajectories.
+Formally, the prediction function $\chi_{u}^{*}$ is defined by Example 1, and the accuracy measure at time $t$ for predicted states $(s_{i})_{i}$ is defined as
+$\text{accuracy}(u)=\frac{1}{n_{\text{test}}}\sum_{i=1}^{n_{\text{test}}}\mathds{1}_{\chi_{u\wedge C_{i}}(\mathcal{T}_{i})=s_{i}}.$
+Although $\mathcal{T}_{i}^{*}$ is unkown, the quantity $\chi_{u\wedge C_{i}}(\mathcal{T}_{i})$ may be computed based on the censored trajectories.
 However, incorporating censoring times into prediction functions relies on information that is, in principle, not available at prediction time.
@@
 8 Conclusion
-We presented jmstate, a Python package for joint modeling of nonlinear longitudinal biomarkers and multi-state survival processes.
-The framework unifies both components on arbitrary directed graphs, extending classical models to Markov and semi-Markov settings and supporting nonlinear submodels.
-Inference is implemented with scalable stochastic gradient methods.
+We presented a general likelihood-based framework for joint modeling of nonlinear longitudinal biomarkers and multi-state survival processes.
+The model unifies both components on arbitrary directed graphs, extending classical joint models to accommodate Markov and semi-Markov structures, recurrent transitions, and nonlinear mixed-effects submodels.
+This formulation provides a flexible and rigorous foundation for linking longitudinal biomarker trajectories to complex event histories within a single coherent probabilistic model.
 Simulation studies confirmed robust parameter recovery, and an application to the PAQUID cohort highlighted the ability to capture the interplay between cognitive decline and physical dependency.
@@
 The authors would like to thank Jean-Benoist Leger for helpful discussions.
-[…]
+Conflict of interest
+The authors declare that they have no conflicts of interest.
+Data availability
+The data analyzed in this study are publicly available via the R package timeROC (dataset Paquid).
+Appendix A Multi-State Simulation
+In nonlinear joint models with known design and known parameters parameters $\theta=(\gamma,Q,R,\alpha,\beta)$, the occurrence times of events conditionally on the latent variables $b_{i}$ and the covariates $X_{i}$ can be easily sampled using the inverse cumulative distribution function transform method.
+This may be achieved through bisection or other root-finding algorithms.
+The simulation of the transition process $\mathcal{T}_{i}^{*}$ conditionally on $b_{i}$, $X_{i}$, and $(T_{i0},S_{i0})$ can be achieved drawing on the semi-Markov property by considering one transition at a time, according to Algorithm 2.
+The algorithme is similar to Gillespie’s algorithm [gillespie2007stochastic], and can also include a survival condition $T_{i1}\geq t_{i}^{\text{surv}}$ that uses the Chasles relation.
+Algorithm 2 Simulation of trajectory $i$
+1:subject-specific censoring time $C_{i}\in\mathbb{R}\cup\{+\infty\}$; covariates $X_{i}$; latent variables $b_{i}$; initial time-state pair $(T_{i0},S_{i0})$; optional survival condition $T_{i1}\geq t_{i}^{\text{surv}}$, defaults to $t_{i}^{\text{surv}}=$-$\infty$.
+2:
+3:Initialize $\ell_{i}\leftarrow 1$, $\mathcal{T}_{i}\leftarrow\left((T_{i0},S_{i0})\right)$.
+4:
+5:while $T_{i(\ell_{i}-1)}<C_{i}$ and $\{s:(S_{i(\ell_{i}-1)},s)\in E\}\neq\emptyset$ do
+6:  for each $s$ with $(S_{i(\ell_{i}-1)},s)\in E$ do
+7:   Draw $T_{i\ell_{i}}^{s}$ such that
+8:      $\log\mathbb{P}(T_{i\ell_{i}}^{s}>t)=\int_{T_{i(\ell_{i}-1)}\vee t_{i}^{\text{surv}}}^{t}\lambda_{i}^{s\mid S_{i(\ell_{i}-1)}}(w\mid b_{i},X_{i},T_{i(\ell_{i}-1)},\theta)\,dw,$
+9:      $\forall t\geq T_{i(\ell_{i}-1)}\vee u.$
+10:  end for
+11:
+12:  Set $T_{i\ell_{i}}\leftarrow\min_{m}T_{i\ell_{i}}^{s}$, and $S_{i\ell_{i}}\leftarrow\arg\min_{m}T_{i\ell_{i}}^{s}$.
+13:  Append $(T_{i\ell_{i}},S_{i\ell_{i}})$ to $\mathcal{T}_{i}$.
+14:  $\ell_{i}\leftarrow\ell_{i}+1$.
+15:end while
+16:
+17:if $T_{i(\ell_{i}-1)}>C_{i}$ then
+18:  Remove the last pair: $\mathcal{T}_{i}\leftarrow\mathcal{T}_{i}[:-1]$.
+19:end if
+20:return $\mathcal{T}_{i}$
+The proof of the exactness of this algorithm is given below.
 Proof.
@@
 $=\prod_{s:\,(S_{i\ell_{i}},s)\in E}\mathbb{P}(T_{i(\ell_{i}+1)}^{s}>t\mid T_{i\ell_{i}},S_{i\ell_{i}}),$
-$=\exp\Big(-\sum_{s:\,(S_{i\ell_{i}},s)\in E}\Lambda_{i}^{s\mid S_{i(\ell_{i}-1)}}\big(t\mid\theta,X_{i},T_{i\ell_{i}},b_{i})\Big).$
+$=\exp\Big(-\sum_{s:\,(S_{i\ell_{i}},s)\in E}\Lambda_{i}^{s\mid S_{i(\ell_{i}-1)}}\big(t\mid b_{i},X_{i},T_{i\ell_{i}},\theta)\Big).$
 Furthermore:
@@
 $\propto p(\{T_{i(\ell_{i}+1)}^{S_{i(\ell_{i}+1)}}=t\}\cap\bigcap_{\begin{subarray}{c}s:\,(S_{i\ell_{i}},s)\in E,\\ s\neq S_{i(\ell_{i}+1)}\end{subarray}}\{T_{i(\ell_{i}+1)}^{s}\geq t\}\mid T_{i\ell_{i}},S_{i\ell_{i}}),$
-$=\lambda_{i}^{S_{i(\ell_{i}+1)}\mid S_{i\ell_{i}}}\big(t\mid\theta,X_{i},T_{i\ell_{i}},b_{i})\exp\Big(-\sum_{s:\,(S_{i\ell_{i}},s)\in E}\Lambda_{i}^{s\mid S_{i\ell_{i}}}\big(t\mid\theta,X_{i},T_{i\ell_{i}},b_{i})\Big).$
-Combining both, we retrieve the same joint density as in formula 3.3.
-If $t_{i}^{\text{surv}}>$-$\infty$, using the Chasles relation LABEL:rel:chasles, one can also check that both densities are equal.
+$=\lambda_{i}^{S_{i(\ell_{i}+1)}\mid S_{i\ell_{i}}}\big(t\mid\theta,X_{i},T_{i\ell_{i}},b_{i})\exp\Big(-\sum_{s:\,(S_{i\ell_{i}},s)\in E}\Lambda_{i}^{s\mid S_{i\ell_{i}}}\big(t\mid b_{i},X_{i},T_{i\ell_{i}},\theta)\Big).$
+Combining both, we retrieve the same joint density.
+If $t_{i}^{\text{surv}}>$-$\infty$, one can also check that both densities are equal.
 Note that the condtioning only affects the first transition, as the next time $\forall\ell_{i}>0,\,T_{i\ell_{i}}\geq t_{i}^{\text{surv}}\implies\forall\ell_{i}>0,\,T_{i\ell_{i}}\vee t_{i}^{\text{surv}}=T_{i\ell_{i}}$.
-[…]
+Appendix B Expression of the Semi-Markov likelihood
 Proof.
-The proof is short and relies on the fact that, using the Semi-Markovian assumption 6:
-$p\left((T_{i(l+1)},S_{i(l+1)})_{0\leq l\leq m_{i}-1}\mid\theta,X_{i},b_{i}\right)$ $=\prod_{l=0}^{m_{i}-1}p\left((T_{i(l+1)},S_{i(l+1)})\mid\theta,X_{i},(T_{il^{\prime}},S_{il^{\prime}})_{l^{\prime}\leq l},b_{i}\right),$
-$=\prod_{l=0}^{m_{i}-1}p\left((T_{i(l+1)},S_{i(l+1)})\mid\theta,X_{i},(T_{il},S_{il}),b_{i}\right).$
-On the other hand, the probability that no additional event is observed between $T_{im_{i}}$ and $C_{i}$ is $1$ if the last state reached is absorbing, and otherwise:
-$\mathbb{P}(T_{i(m_{i}+1)}>C_{i}\mid\theta,X_{i},(T_{im_{i}},S_{im_{i}}),b_{i})=\exp\left(-\sum_{s:\,(S_{im_{i}},s)\in E}\Lambda_{i}^{s\mid S_{im_{i}}}(C_{i}\mid\theta,X_{i},T_{im_{i}},b_{i})\right),$
+The proof is short and relies on the fact that, using the semi-Markov property 2
+$p\left((T_{i(\ell+1)},S_{i(\ell+1)})_{\ell\leq m_{i}-1}\mid b_{i},X_{i},\theta\right)$ $=\prod_{\ell=0}^{m_{i}-1}p\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid b_{i},X_{i},(T_{i\ell^{\prime}},S_{i\ell^{\prime}})_{\ell^{\prime}\leq\ell},\theta\right),$
+$=\prod_{\ell=0}^{m_{i}-1}p\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid b_{i},X_{i},(T_{i\ell},S_{i\ell}),\theta\right).$
+On the other hand, the probability that no additional event is observed between $T_{im_{i}}$ and $C_{i}$ is $1$ if the last state reached is absorbing, and otherwise
+$\mathbb{P}(T_{i(m_{i}+1)}>C_{i}\mid\theta,X_{i},(T_{im_{i}},S_{im_{i}}),b_{i})=\exp\left(-\sum_{s:\,(S_{im_{i}},s)\in E}\Lambda_{i}^{s\mid S_{im_{i}}}(C_{i}\mid b_{i},X_{i},T_{im_{i}},\theta)\right),$
 and so we recover the formula cited above with the convention that an empty sum is 0.
+Appendix C Stochastic optimization process
+(a) $\gamma$
+(b) $Q$
+(c) $R$
+(d) $\alpha^{1\mid 0}$
+(e) $\alpha^{2\mid 0}$
+(f) $\alpha^{2\mid 1}$
+(g) $\beta^{1\mid 0}$
+(h) $\beta^{2\mid 0}$
+(i) $\beta^{2\mid 1}$
+Figure 10: Evolution of the parameters during the optimization of the marginal log-likelihood using stochastic gradient descent. The dotted lines correspond to the true values.

### v3

@@
 Abstract
-Classical joint modeling approaches often rely on competing risks or recurrent event formulations to describe complex processes involving evolving longitudinal biomarkers and discrete event occurrences, but these frameworks typically capture only limited aspects of the underlying event dynamics.
+Conventional joint modeling approaches generally characterize the relationship between longitudinal biomarkers and discrete event occurrences within terminal, recurring or competing risk settings, thereby offering a limited representation of complex, multi-state trajectories.
 We propose a general multi-state joint modeling framework that unifies longitudinal biomarker dynamics with multi-state time-to-event processes defined on arbitrary directed graphs.
-The proposed framework accommodates arbitrary directed transition graphs, nonlinear longitudinal submodels, and scalable inference via stochastic gradient descent.
+The proposed framework also accomodates nonlinear longitudinal submodels and scalable inference via stochastic gradient descent.
 This formulation encompasses both Markovian and semi-Markovian transition structures, allowing recurrent cycles and terminal absorptions to be naturally represented.
 The longitudinal and event processes are linked through shared latent structures within nonlinear mixed-effects models, extending classical joint modeling formulations.
-We derive the complete likelihood, establish conditions for identifiability, and develop scalable inference procedures based on stochastic gradient descent to enable high-dimensional and large-scale applications.
+We derive the complete likelihood, model selection criteria, and develop scalable inference procedures based on stochastic gradient descent to enable high-dimensional and large-scale applications.
 In addition, we formulate a dynamic prediction framework that provides individualized state-transition probabilities and personalized risk assessments along complex event trajectories.
@@
 However, many real-world processes involve multiple possible outcomes, intermediate stages, or recurrent events, which cannot be fully captured by a single-event framework [KrolMauguen2017].
-In such settings, multi-state models provide a more flexible approach [ferrer2016joint].
+To address these limitations, specialized frameworks have been developed to handle competing risks and recurrent events.
+However, these remain constrained to specific event types and fail to accommodate more general multi-state transitions.
+In such settings, multi-state models provide a more general and flexible approach [ferrer2016joint].
 Multi-state models represent an overall joint process of clinical course as a series of discrete stages or health states that occur sequentially [lovblom2024modeling].
@@
 Often, these models assume the Markov property, where future transitions depend only on the current state, simplifying the transition intensity functions [asanjarani2022estimation].
-The Markov assumption can be relaxed by adopting a semi-Markov formulation, in which the transition probabilities depend not only on the current state but also on the sojourn time, effectively resetting the time scale after each transition.
+This assumption can be relaxed by adopting a semi-Markov formulation, in which the transition probabilities depend not only on the current state but also on the sojourn time, effectively resetting the time scale after each transition.
 The link between multi-state models and joint models arises when a multi-state process is integrated as a component within a broader joint modeling framework .
-While multi-state models, such as those described in the “sequential state framework” primarily focus on movements between discrete states, joint models operate under a “parallel trajectory framework” that combines a longitudinal process with a time-to-event process.
+While multi-state models, such as those described in the “sequential state framework” primarily focus on movements between discrete states, joint models operate under a “parallel trajectory framework” that combines a longitudinal process with an acyclic time-to-event process.
 The model proposed by Ferrer et al.
@@
 Section 2 reviews the background on multi-state and joint modeling frameworks.
-Section 3.2 introduces the proposed general multi-state joint modeling framework, detailing the likelihood formulation and identifiability conditions.
+Section 3.2 introduces the proposed general multi-state joint modeling framework, detailing the likelihood formulation and proposed model selection creteria.
 Section 4 describes the stochastic-gradient–based inference algorithm, while Section 5 presents the dynamic prediction framework.
-Section 6 reports simulation experiments that assess parameter recovery, identifiability, and convergence properties of the proposed model, and Section 7 applies the framework to data from the PAQUID cohort.
+Section 6 reports simulation experiments that assess parameter recovery and convergence properties of the proposed model, whereas Section 7 applies the framework to data from the PAQUID cohort.
 Finally, Section 8 concludes with perspectives and future research directions.
@@
 Such models have become central in biomedical research, particularly for dynamic prediction and personalized risk assessment [papageorgiou2019overview].
-However, these traditional joint models generally focus on a single terminal event, limiting their ability to represent more complex event histories that include intermediate, recurrent, or competing events.
+However, these traditional joint models generally focus either on a single terminal event, competing, or recurrent events, limiting their ability to represent more complex event histories that include intermediate, recurrent, or competing events.
 To address these limitations, multi-state extensions have been developed, providing a natural framework for describing transitions between multiple clinical or biological states over time [ferrer2016joint, KrolMauguen2017].
@@
 3 A General Multi-State Joint Modeling Framework
-Classical joint models typically link a longitudinal biomarker process to a single time-to-event outcome.
-While effective for terminal events, this design cannot represent more complex disease trajectories involving intermediate or recurrent events.
-To address these limitations, we extend the nonlinear joint modeling framework to incorporate a multi-state process, thereby enabling a unified analysis of longitudinal biomarkers and discrete transitions between multiple health states.
+To address the limitations imposed by traditional joint modeling frameworks, we extend them to incorporate a multi-state process, thereby enabling a unified analysis of longitudinal biomarkers and discrete transitions between multiple health states.
 The proposed model defines a joint process $\{Y_{i}(t),S_{i}(t)\}_{t\geq 0}$, where $Y_{i}(t)$ denotes the longitudinal biomarker trajectory and $S_{i}(t)$ the multi-state event process.
@@
 The vector $Y_{ij}\in\mathbb{R}^{d}$ represents longitudinal biomarker measurements, and $X_{i}\in\mathbb{R}^{p}$ the associated covariates.
-The latent random effects $b_{i}\sim\mathcal{N}(0,Q)$ capture individual heterogeneity, while $\psi_{i}=f(\gamma,X_{i},b_{i})$ denotes the subject-specific parameters in the longitudinal submodel.
+The latent random effects $b_{i}\sim\mathcal{N}(0,Q)$ capture individual heterogeneity, while $\psi_{i}\coloneqq f(\gamma,X_{i},b_{i})$ denotes the subject-specific parameters in the longitudinal submodel.
 The filtration $\mathcal{H}_{i}(t)$ represents the observed history of the longitudinal process up to time $t$.
 Transition times and states are denoted by $(T_{i\ell},S_{i\ell})$, with right-censoring time $C_{i}$.
-We also define the index of the last observed transition before time $t$ denoted by $m_{i}(t)\coloneq\sup\{\ell\geq 0:T_{i\ell}\leq t\}$.
+We also define the index of the last observed transition before time $t$ denoted by $m_{i}(t)\coloneqq\sup\{\ell\geq 0:T_{i\ell}\leq t\}$.
 3.2 Extending Multi-State Joint Models
-Joint models traditionally focus on a single time-to-event outcome, which limits their ability to represent complex disease trajectories involving multiple intermediate and recurrent events.
-To address this limitation, we extend the classical nonlinear joint modeling framework to incorporate multi-state processes, enabling the joint analysis of longitudinal biomarkers and transitions between multiple health states.
-This is achieved by introducing a directed graph structure encoding all permissible transitions and supports both Markov and semi-Markov assumptions.
+The joint model proposed by Ferrer et al.
+[ferrer2016joint] established a foundational link between longitudinal biomarkers and multi-state processes through shared random effects.
+However, their formalism relies on binary censoring indicators to characterize transitions, thereby inherently restricting the modeling framework to Directed Acyclic Graphs (DAGs) and excluding recurrent events.
+To overcome this limitation, we present a new method based on arbitrary graph transitions.
+By shifting away from binary indicators toward a more flexible graph-theoretic specification, our approach accommodates complex disease trajectories, including recurrent events and cyclic transitions that cannot be captured within a DAG-based structure.
 3.2.1 Graph Structure
@@
 Each individual $i$ follows a latent trajectory
-$\mathcal{T}_{i}^{*}=\big((T_{i0},S_{i0}),(T_{i1},S_{i1}),(T_{i2},S_{i2}),\dots\big),$
+$\mathcal{T}_{i}^{*}\coloneqq\big((T_{i0},S_{i0}),(T_{i1},S_{i1}),(T_{i2},S_{i2}),\dots\big),$
 where $T_{i\ell}$ denotes the $\ell$-th transition time and $S_{i\ell}\in V$ the corresponding state.
@@
 For any admissible transition $(k,k^{\prime})\in E$, let $\lambda_{i}^{k^{\prime}\mid k}(t\mid t_{0})$ denote the instantaneous risk of moving from $k$ to $k^{\prime}$ at time $t$, given entry time $t_{0}$
-$\lambda_{i}^{k^{\prime}\mid k}(t\mid t_{0})=\lim_{\delta\to 0^{+}}\frac{\mathbb{P}(T_{i,\ell+1}\leq t+\delta,S_{i,\ell+1}=k^{\prime}\mid T_{i,\ell+1}>t,T_{i\ell}=t_{0},S_{i\ell}=k)}{\delta}.$
-The cumulative intensity and the corresponding survival function are
+$\lambda_{i}^{k^{\prime}\mid k}(t\mid t_{0})\coloneqq\lim_{\delta\to 0^{+}}\frac{\mathbb{P}(T_{i,\ell+1}\leq t+\delta,S_{i,\ell+1}=k^{\prime}\mid T_{i,\ell+1}>t,T_{i\ell}=t_{0},S_{i\ell}=k)}{\delta}.$
+The cumulative risk and the corresponding survival function are
 $\Lambda_{i}^{k^{\prime}\mid k}(t\mid t_{0})=\int_{t_{0}}^{t}\lambda_{i}^{k^{\prime}\mid k}(w\mid t_{0})\,dw,$
@@
 The Akaike Information Criterion is computed as follows
-$\mathrm{AIC}=-2\log\mathcal{L}_{\text{marginal}}(\hat{\theta};X,Y,\mathcal{T},C)+2k,$
+$\mathrm{AIC}\coloneqq-2\log\mathcal{L}_{\text{marginal}}(\hat{\theta};X,Y,\mathcal{T},C)+2k,$
 where $k$ is the number of parameters, and $\hat{\theta}$ is a maximizer of the marginal likelihood.
@@
 It is to be noted that this integral is typically intractable in closed form due to the nonlinear nature of the model, necessitating numerical approximation methods such as Monte Carlo integration or Laplace approximation.
-Several optimization methods adapted from the nonlinear mixed-effects literature can be applied, including Stochastic EM [kuhn2004coupling], Laplace approximation [wolfinger1993laplace], Gauss-Hermite quadrature, stochastic gradient ascent with Robbins–Monro updates [robbins1951stochastic], and MCMC-based approximations such as Metropolis-Hastings and Hamiltonian Monte Carlo.
+Several optimization methods adapted from the nonlinear mixed-effects literature can be applied, including Stochastic EM [kuhn2004coupling], Laplace approximation [wolfinger1993laplace] and Gauss–Hermite quadrature.
 These strategies are implemented in software such as JMBayes [jmbayes22024, rizopoulos2020package].
-Another approach requiring mild regularity assumptions on the $\log$ marginal likelihood, but without the need for the model to belong in the exponential family, is to consider a stochastic gradient ascent scheme [caillebotte2025estimation, baey2023efficient] using Fisher’s identity and following the Robbins-Monro procedure [robbins1951stochastic].
-Under interchangeability of integral and differentiation, setting $x=(X,Y,\mathcal{T},C)$ for convenience, the Fisher identity writes
+Another approach requiring mild regularity assumptions on the $\log$ marginal likelihood, but without the need for the model to belong in the exponential family, is to consider a stochastic gradient ascent scheme [caillebotte2025estimation, baey2023efficient] using Fisher’s identity and following the Robbins-Monro procedure [robbins1951stochastic], alongside with MCMC-based approximations such as Metropolis-Hastings and Hamiltonian Monte Carlo.
+Under interchangeability of integral and differentiation, setting $x\coloneqq(X,Y,\mathcal{T},C)$ for convenience, the Fisher identity writes
 $\nabla_{\theta}\log\mathcal{L}_{\text{marginal}}(\theta;x)=\mathbb{E}_{b\sim p(\cdot\mid x,\theta)}\left(\nabla_{\theta}\log\mathcal{L}_{\text{complete}}(\theta;x,b)\right).$
@@
 Therefore, a Monte Carlo estimation scheme may be devised under certain restrictions, where we estimate $\chi^{*}(\mathcal{T}_{i}^{*})$ by its expected value:
-$\hat{\chi}_{i}=\mathbb{E}_{\mathcal{T}^{\prime}_{i}\sim\mathscr{L}(\mathcal{T}_{i}^{*}\mid\mathcal{Y}(t),\mathcal{T}_{i},t)}\left(\chi^{*}(\mathcal{T}^{\prime}_{i})\right)\approx\frac{1}{B}\sum_{k=1}^{B}\chi^{*}(\mathcal{T}_{i}^{(k)}),\,\mathcal{T}_{i}^{(k)}\overset{\text{i.i.d.}}{\sim}\mathscr{L}(\mathcal{T}_{i}^{*}\mid\mathcal{Y}(t),\mathcal{T}_{i},t).$
+$\hat{\chi}_{i}=\coloneqq\mathbb{E}_{\mathcal{T}^{\prime}_{i}\sim\mathscr{L}(\mathcal{T}_{i}^{*}\mid\mathcal{Y}(t),\mathcal{T}_{i},t)}\left(\chi^{*}(\mathcal{T}^{\prime}_{i})\right)\approx\frac{1}{B}\sum_{k=1}^{B}\chi^{*}(\mathcal{T}_{i}^{(k)}),\,\mathcal{T}_{i}^{(k)}\overset{\text{i.i.d.}}{\sim}\mathscr{L}(\mathcal{T}_{i}^{*}\mid\mathcal{Y}(t),\mathcal{T}_{i},t).$
 In general however, $\chi^{*}$ may very well depend on (countably) infinitely many transitions.
@@
 Assumption 1.
-Assume there exists $\chi:\bigcup_{n\geq 1}(\mathbb{R}\times\mathbb{S})^{n}\to\mathbb{Y}$ and $\tau_{i}$ a stopping time for the filtration $\mathcal{F}_{in}=\sigma\left((T_{il},S_{il})_{0\leq l\leq n}\right)$ such that:
+Assume there exists $\chi:\bigcup_{n\geq 1}(\mathbb{R}\times\mathbb{S})^{n}\to\mathbb{Y}$ and $\tau_{i}$ a stopping time for the filtration $\mathcal{F}_{in}\coloneqq\sigma\left((T_{il},S_{il})_{0\leq l\leq n}\right)$ such that:
 $\begin{cases}\tau_{i}<+\infty\text{ a.s.},\\ \chi^{*}(\mathcal{T}_{i}^{*})=\chi\left((T_{il},S_{il})_{1\leq l\leq\tau_{i}}\right)\end{cases}$
@@
 First, for a directed acyclic graph $G=(V,E)$, we define the depth of $G$, denoted by $\operatorname{depth}(G)$, as the length of the longest directed path in $G$, that is,
-$\operatorname{depth}(G)=\max_{(v_{0},v_{1},\dots,v_{k})}k,$
+$\operatorname{depth}(G)\coloneqq\max_{(v_{0},v_{1},\dots,v_{k})}k,$
 where $(v_{0},v_{1},\dots,v_{k})$ ranges over all directed paths in $G$.
@@
 6 Simulation Study
-The purpose of this simulation study is to evaluate the finite-sample performance, identifiability, and convergence properties of the proposed inference algorithm under controlled conditions where the true parameters are known.
+The purpose of this simulation study is to evaluate the finite-sample performance and convergence properties of the proposed inference algorithm under controlled conditions where the true parameters are known.
 We simulate data from a three-state semi-Markov process coupled with a nonlinear longitudinal biomarker trajectory, representing a simplified disease progression model.
@@
 $\beta_{1}$ -0.2990 0.9130 $\beta_{2}$ 0.0970 1.0300
-Figure 8: Inferred parameters and their estimated standard deviations. Bold values mean $p<0.05$.
+Figure 8: Inferred parameters and their estimated standard deviations. Bold values mean $p<0.05$ (variance parameters excluded).
 Notably, the estimated standard errors are consistent with the number of observations per transition, with larger standard errors for transitions with fewer observations.

### v4

@@
 We propose a general multi-state joint modeling framework that unifies longitudinal biomarker dynamics with multi-state time-to-event processes defined on arbitrary directed graphs.
-The proposed framework also accomodates nonlinear longitudinal submodels and scalable inference via stochastic gradient descent.
+The proposed framework also accommodates nonlinear longitudinal submodels and scalable inference via stochastic gradient ascent.
 This formulation encompasses both Markovian and semi-Markovian transition structures, allowing recurrent cycles and terminal absorptions to be naturally represented.
 The longitudinal and event processes are linked through shared latent structures within nonlinear mixed-effects models, extending classical joint modeling formulations.
-We derive the complete likelihood, model selection criteria, and develop scalable inference procedures based on stochastic gradient descent to enable high-dimensional and large-scale applications.
+We derive the complete likelihood, model selection criteria, and develop scalable inference procedures based on stochastic gradient ascent to enable high-dimensional and large-scale applications.
 In addition, we formulate a dynamic prediction framework that provides individualized state-transition probabilities and personalized risk assessments along complex event trajectories.
@@
 However, these remain constrained to specific event types and fail to accommodate more general multi-state transitions.
-In such settings, multi-state models provide a more general and flexible approach [ferrer2016joint].
+In such settings, multi-state models provide a more general and flexible approach [ferrer2016joint, you2024joint].
 Multi-state models represent an overall joint process of clinical course as a series of discrete stages or health states that occur sequentially [lovblom2024modeling].
@@
 The link between multi-state models and joint models arises when a multi-state process is integrated as a component within a broader joint modeling framework.
-While multi-state models, such as those described in the “sequential state framework” primarily focus on movements between discrete states, joint models operate under a “parallel trajectory framework” that combines a longitudinal process with an acyclic time-to-event process.
+While multi-state models, such as those described in the “sequential state framework” primarily focus on movements between discrete states, joint models operate under a “parallel trajectory framework” that combines a longitudinal process with a time-to-event process.
 The model proposed by Ferrer et al.
 [ferrer2016joint] exemplifies this link by presenting a joint model for a longitudinal process (e.g., Prostate-Specific Antigen (PSA) measurements) and a multi-state process (e.g., clinical progressions in prostate cancer).
-These two sub-models are interconnected by shared random effects, allowing the model to account for the correlation between the continuous longitudinal biomarker trajectory and the discrete transitions between health states.
+These two sub-models are interconnected by shared random effects, allowing the model to account for the correlation between a continuous longitudinal biomarker trajectory and acyclic discrete transitions between health states.
+Specifically, the previously proposed mechanism [ferrer2016joint, you2024joint] is based on binary censoring indicators for each state, implying that any given state can be entered at most once, therefore precluding the modeling of recurrent events.
 The model we propose operates on an arbitrary directed graph, enabling representation of complex state transitions and recurrent events.
@@
 Section 2 reviews the background on multi-state and joint modeling frameworks.
-Section 3.2 introduces the proposed general multi-state joint modeling framework, detailing the likelihood formulation and proposed model selection creteria.
+Section 3.2 introduces the proposed general multi-state joint modeling framework, detailing the likelihood formulation and proposed model selection criteria.
 Section 4 describes the stochastic-gradient–based inference algorithm, while Section 5 presents the dynamic prediction framework.
-Section 6 reports simulation experiments that assess parameter recovery and convergence properties of the proposed model, whereas Section 7 applies the framework to data from the PAQUID cohort.
+Section 6 reports simulation experiments that assess parameter recovery and convergence properties of the proposed model, whereas Section 7 applies the framework to data from the PAQUID cohort [Letenneur1994].
 Finally, Section 8 concludes with perspectives and future research directions.
@@
 In an MSMP, the process is defined by a directed graph $G=(V,E)$, where $V$ represents the set of states and $E$ the set of possible transitions.
-The transition intensities $\lambda_{k^{\prime}\mid k}(t\mid t_{0})$ from state $k$ to state $k^{\prime}$ at time $t$, given entry time $t_{0}$ are therefore assumed to be non-identically zero.
+The transition intensities $\lambda_{k\to k^{\prime}}(t\mid t_{0})$ from state $k$ to state $k^{\prime}$ at time $t$, given entry time $t_{0}$, are therefore assumed to be non identically zero.
 The sojourn time in state $k$ is not restricted to an exponential distribution, allowing for more realistic modeling of the time spent in each state.
 MSMPs have been widely applied in a range of disciplines.
-In reliability, they are used to modeling degradation and repair processes [limnios2001semi].
+In reliability, they are used to model degradation and repair processes [limnios2001semi].
 In biomedical studies, they have proven useful for analyzing illness-death models or disease progression with non-exponential transitions [commenges2006likelihood, fiocco2008modelling].
@@
 The multi-state sub-model assumes a non-homogeneous Markov process for the clinical progression, meaning that the future of the process depends only on the current state (Markov property) and the time elapsed since study entry (non-homogeneous property).
-Although Ferrer et al.
-[ferrer2016joint] note that their joint modeling framework could be adapted to semi-Markov processes in other contexts, their primary application to prostate cancer progression assumes a non-homogeneous Markov multi-state process.
-However, this framework is not well suited for arbitrary transition patterns other than the ones modeled by DAGs.
 While both joint models and multi-state models offer powerful tools for analyzing disease progression, they emphasize different aspects of the underlying process.
@@
 Existing joint multi-state models, such as that of Ferrer et al.
-[ferrer2016joint], are restricted to linear mixed-effects submodels and to acyclic Markov transition graphs, as their likelihood relies on a sequential factorization that is not valid when transitions form cycles.
+[ferrer2016joint], are restricted to linear mixed-effects submodels and to acyclic Markov transition graphs, since their reliance on binary censoring indicators inherently prevents modeling recurrent events.
 However, many biological processes involve cyclic or semi-Markov transitions and nonlinear biomarker dynamics.
@@
 To address the limitations imposed by traditional joint modeling frameworks, we extend them to incorporate a multi-state process, thereby enabling a unified analysis of longitudinal biomarkers and discrete transitions between multiple health states.
-The proposed model defines a joint process $\{Y_{i}(t),S_{i}(t)\}_{t\geq 0}$, where $Y_{i}(t)$ denotes the longitudinal biomarker trajectory and $S_{i}(t)$ the multi-state event process.
+The proposed model defines a joint process $\{Y_{i}(t),S_{i}(t)\}_{t}$, where $Y_{i}(t)$ denotes the longitudinal biomarker trajectory and $S_{i}(t)$ the multi-state event process.
 Both components are linked through shared latent structures, and the event process is represented on an arbitrary directed graph supporting both Markovian and semi-Markovian transition mechanisms.
 3.1 Notation Conventions
-Throughout this section, $i=1,\dots,n$ indexes individuals, and $t_{ij}$ denotes the $j$-th observation time of individual $i$.
+Throughout this section, let $i\in\llbracket n\rrbracket$ index $n\geq 1$ individuals, and let $t_{ij}$ denote the $j$-th observation time of individual $i$.
 The vector $Y_{ij}\in\mathbb{R}^{d}$ represents longitudinal biomarker measurements, and $X_{i}\in\mathbb{R}^{p}$ the associated covariates.
-The latent random effects $b_{i}\sim\mathcal{N}(0,Q)$ capture individual heterogeneity, while $\psi_{i}\coloneqq f(\gamma,X_{i},b_{i})$ denotes the subject-specific parameters in the longitudinal submodel.
-The filtration $\mathcal{H}_{i}(t)$ represents the observed history of the longitudinal process up to time $t$.
-Transition times and states are denoted by $(T_{i\ell},S_{i\ell})$, with right-censoring time $C_{i}$.
-We also define the index of the last observed transition before time $t$ denoted by $m_{i}(t)\coloneqq\sup\{\ell\geq 0:T_{i\ell}\leq t\}$.
-3.2 Extending Multi-State Joint Models
+The latent random effects $b_{i}\sim\mathcal{N}(0,Q)$ capture individual heterogeneity, $\gamma$ represent shared characteristics across all individuals, and the subject-specific parameters are defined as $\psi_{i}\coloneqq f(\gamma,X_{i},b_{i})$.
+The set $\mathcal{H}_{i}(t)\coloneqq\left\{(\sigma,h(s,\psi_{i})):s\leq t\right\}$ represents the true underlying history of the longitudinal process up to time $t$.
+Correspondingly, $\mathcal{Y}_{i}(t)$ denotes the observed noisy longitudinal history, and $\mathcal{S}_{i}(t)$ denotes the observed trajectory of the semi-Markovian dynamics.
+Transition times and states are denoted by $(T_{i\ell},S_{i\ell})\in\mathbb{R}^{+}\times V$, where $V$ is the space of admissible states, and are subject to right-censoring at time $C_{i}\in\bar{\mathbb{R}}^{+}$.
+We define the index of the last observed transition before time $t$ as $m_{i}(t)\coloneqq\sup\{\ell\geq 0:T_{i\ell}\leq t\}$.
+The complete sequence of transition times and states for individual $i$ is denoted by $\mathcal{T}_{i}^{*}$, whereas $\mathcal{T}_{i}$ denotes the corresponding right-censored sequence.
+3.2 Multi-State Processes on Arbitrary Graphs
 The joint model proposed by Ferrer et al.
@@
 However, their formalism relies on binary censoring indicators to characterize transitions, thereby inherently restricting the modeling framework to Directed Acyclic Graphs (DAGs) and excluding recurrent events.
+Indeed, the proposed model uses, for each state $k\in E$, a tuple $(T_{k},\delta_{k})$, where $T_{k}$ denotes the observed or censored time until leaving state $k$ and $\delta_{k}$ is an indicator equal to $1$ if the transition was observed and 0 if the observation was censored (i.e., the transition was not observed).
 To overcome this limitation, we present a new method based on arbitrary graph transitions.
@@
 3.2.1 Graph Structure
-Let $G=(V,E)$ be a directed graph, where $V=\{1,\dots,p\}$ denotes the set of states and $E\subset V\times V$ the set of admissible transitions.
+Let $G=(V,E)$ be a directed graph, where $V=\llbracket p\rrbracket$ denotes the set of states and $E\subseteq V\times V$ the set of admissible transitions.
 The graph encodes all possible paths of the multi-state process, allowing for competing, recurrent, or absorbing transitions.
-Figure 1 illustrates an example of a four-state model including an absorbing state (Death).
-The corresponding adjacency matrix $A=(A_{kk^{\prime}})_{1\leq k,k^{\prime}\leq|V|}$ satisfies $A_{kk^{\prime}}=1$ if $(k,k^{\prime})\in E$ and 0 otherwise.
+Figure 1 illustrates an example of a four-state model, including an absorbing state (Death).
+The corresponding adjacency matrix $A=(A_{kk^{\prime}})_{1\leq k,k^{\prime}\leq p}$ satisfies $A_{kk^{\prime}}=1$ if $(k,k^{\prime})\in E$ and 0 otherwise.
 Figure 1: Example of a 4-state transition graph $G=(V,E)$, including an absorbing state (4).
@@
 3.2.2 Individual Trajectories
-Each individual $i$ follows a latent trajectory
-$\mathcal{T}_{i}^{*}\coloneqq\big((T_{i0},S_{i0}),(T_{i1},S_{i1}),(T_{i2},S_{i2}),\dots\big),$
-where $T_{i\ell}$ denotes the $\ell$-th transition time and $S_{i\ell}\in V$ the corresponding state.
-The observed trajectory is right-censored at time $C_{i}$, so that only $\{(T_{i\ell},S_{i\ell}):T_{i\ell}\leq C_{i}\}$ is observed.
-Between transitions, the state $S_{i}(t)$ is piecewise constant.
-We note $m_{i}(t)\coloneq\sup\{\ell\geq 0:T_{i\ell}\leq t\}$ the index of the last observed transition before time $t$ and $m_{i}=m_{i}(C_{i})$.
+Each individual $i\in\llbracket n\rrbracket$ follows a latent trajectory
+$\mathcal{T}_{i}^{*}\coloneqq\left((T_{i0},S_{i0}),(T_{i1},S_{i1}),(T_{i2},S_{i2}),\dots\right),$
+where $T_{i\ell}\in\mathbb{R}^{+}$ denotes the $\ell$-th transition time and $S_{i\ell}\in V$ the corresponding state.
+The observed trajectory is right-censored at time $C_{i}\in\bar{\mathbb{R}}^{+}$, so that only $\mathcal{T}_{i}\coloneqq\left\{(T_{i\ell},S_{i\ell}):T_{i\ell}\leq C_{i}\right\}$ is observed.
+Between transitions, the state $S_{i}(t)$ is constant, thus the transitions are the discontinuity points of the map $t\mapsto\mathcal{S}_{i}(t)$, of which we assume there are at most only infinitely countably many.
+Letting $m_{i}(t)\coloneqq\sup\{\ell\geq 0:T_{i\ell}\leq t\}$ be the index of the last observed transition before time $t$ and $m_{i}\coloneqq m_{i}(C_{i})$, we thus assume that no further transitions occur for individual $i$ on the interval $(T_{im_{i}},C_{i}]$.
+We also define $\mathcal{S}_{i}(t)\coloneqq\left\{(s,S_{i}(s)):s\leq t\right\}$ as the observed trajectory up to time $t$, which is equivalent to specifying all transition times and their corresponding states up to $t$, together with the time $t$ itself.
+We emphasize once more that modeling the observed data solely as censored collections of transition times, as in Ferrer et al.
+[ferrer2016joint], is not sufficient to represent such trajectories, since it implicitly restricts each state to be visited at most once and therefore rules out recurrent transitions.
 The process is assumed to satisfy a (possibly time-inhomogeneous) semi-Markov property
-$\mathscr{L}\big((T_{i,\ell+1},S_{i,\ell+1})\mid\{(T_{i\ell^{\prime}},S_{i\ell^{\prime}})\}_{\ell^{\prime}\leq\ell}\big)=\mathscr{L}\big((T_{i,\ell+1},S_{i,\ell+1})\mid(T_{i\ell},S_{i\ell})\big),$ (1)
+$\mathscr{L}\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid\{(T_{i\ell^{\prime}},S_{i\ell^{\prime}})\}_{\ell^{\prime}\leq\ell}\right)=\mathscr{L}\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid(T_{i\ell},S_{i\ell})\right),$ (1)
 which is relaxed compared to a stricter Markov property by allowing sojourn-time dependence.
 3.2.3 Transition Intensities and Survival Functions
-For any admissible transition $(k,k^{\prime})\in E$, let $\lambda_{i}^{k^{\prime}\mid k}(t\mid t_{0})$ denote the instantaneous risk of moving from $k$ to $k^{\prime}$ at time $t$, given entry time $t_{0}$
-$\lambda_{i}^{k^{\prime}\mid k}(t\mid t_{0})\coloneqq\lim_{\delta\to 0^{+}}\frac{\mathbb{P}(T_{i,\ell+1}\leq t+\delta,S_{i,\ell+1}=k^{\prime}\mid T_{i,\ell+1}>t,T_{i\ell}=t_{0},S_{i\ell}=k)}{\delta}.$
-The cumulative risk and the corresponding survival function are
-$\Lambda_{i}^{k^{\prime}\mid k}(t\mid t_{0})=\int_{t_{0}}^{t}\lambda_{i}^{k^{\prime}\mid k}(w\mid t_{0})\,dw,$
-$\mathbb{P}(T_{i,\ell+1}>t\mid T_{i\ell},S_{i\ell})=\exp\left[-\sum_{s:(S_{i\ell},s)\in E}\Lambda_{i}^{s|S_{i\ell}}(t\mid T_{i\ell})\right].$
+Although one could equivalently define intensities for all ordered pairs of states and set those corresponding to inadmissible transitions identically to zero, explicitly restricting attention to a graph of admissible transitions provides a clearer and more coherent representation of the underlying state dynamics.
+For any admissible transition $(k,k^{\prime})\in E$ with nonzero probability, let $\lambda_{i}^{k\to k^{\prime}}(t_{0},t)$ denote the instantaneous risk at time $t$, given entry time $t_{0}\leq t$, and conditionally the transition $k\to k^{\prime}$
+$\lambda_{i}^{k\to k^{\prime}}(t_{0},t)\coloneqq\lim_{\delta\to 0^{+}}\frac{\mathbb{P}(T_{i(\ell+1)}\leq t+\delta\mid T_{i(\ell+1)}>t,T_{i\ell}=t_{0},S_{i\ell}=k,S_{i(\ell+1)}=k^{\prime})}{\delta}.$
+Assuming this quantity to be finite for any pair $(t_{0},t)$ so that the underlying distribution of transition times is continuous, the cumulative risk and the corresponding survival function are
+$\Lambda_{i}^{k\to k^{\prime}}(t_{0},t)=\int_{t_{0}}^{t}\lambda_{i}^{k\to k^{\prime}}(t_{0},s)\,ds,$
+$\forall t\geq T_{i\ell},\,\mathbb{P}(T_{i(\ell+1)}>t\mid T_{i\ell},S_{i\ell})=\exp\left(-\sum_{j:(S_{i\ell},j)\in E}\Lambda_{i}^{S_{i\ell}\to j}(T_{i\ell},t)\right).$
 Conditional transition probabilities are
-$\mathbb{P}(S_{i,\ell+1}=k^{\prime}\mid T_{i,\ell+1},T_{i\ell},S_{i\ell})=\frac{\lambda_{i}^{k^{\prime}|S_{i\ell}}(T_{i,\ell+1}\mid T_{i\ell})}{\sum_{s:(S_{i\ell},s)\in E}\lambda_{i}^{s|S_{i\ell}}(T_{i,\ell+1}\mid T_{i\ell})}.$
+$\mathbb{P}(S_{i(\ell+1)}=k^{\prime}\mid T_{i(\ell+1)},T_{i\ell},S_{i\ell})=\frac{\lambda_{i}^{S_{i\ell}\to k^{\prime}}(T_{i\ell},T_{i(\ell+1)})}{\sum_{j:(S_{i\ell},j)\in E}\lambda_{i}^{S_{i\ell}\to j}(T_{i\ell},T_{i(\ell+1)})},$
+so the joint density of a transition writes as
+$p\left((T_{i(\ell+1)},T_{i(\ell+1)})\mid(T_{i\ell},T_{i\ell})\right)=$ $\frac{\lambda_{i}^{S_{i\ell}\to S_{i(\ell+1)}}(T_{i\ell},T_{i(\ell+1)})}{\sum_{j:(S_{i\ell},j)\in E}\lambda_{i}^{S_{i\ell}\to j}(T_{i\ell},T_{i(\ell+1)})}$
+$\exp\left(-\sum_{j:(S_{i\ell},j)\in E}\Lambda_{i}^{S_{i\ell}\to j}(T_{i\ell},T_{i(\ell+1)})\right)$ (2)
+In practice, numerical integration can be efficiently carried out using quadrature methods such as the Gauss-Legendre quadrature [lether1978construction].
 3.3 Multi-State Joint Model Specification
-The proposed multi-state joint model with Gaussian prior and homoscedastic Gaussian noise is specified by:
-$\begin{cases}Y_{ij}=h(t_{ij},\psi_{i})+\epsilon_{ij},\quad\epsilon_{ij}\sim\mathcal{N}(0,R),\\ \psi_{i}=f(\gamma,X_{i},b_{i}),\quad b_{i}\sim\mathcal{N}(0,Q),\\ \lambda_{i}^{k^{\prime}\mid k}\left(t\mid X_{i},T_{im_{i}(t)},\mathcal{H}_{i}(t)\right)=\lambda_{0}^{k^{\prime}\mid k}(t\mid T_{im_{i}(t)})\exp\big(\alpha^{k^{\prime}\mid k}g^{k^{\prime}\mid k}(t,X_{i},\psi_{i})+\beta^{k^{\prime}\mid k}X_{i}\big).\end{cases}$
-Here, $h$ and $f$ define the nonlinear mixed-effects submodel, $g^{k^{\prime}\mid k}$ represents the link between biomarker dynamics and the transition intensity, and $\lambda_{0}^{k^{\prime}\mid k}$ denotes the baseline hazard associated with transition $(k\to k^{\prime})$.
+Omitting the implicit conditioning on the individual parameters and covariates for the sake of simple notation, the proposed multi-state joint model with a Gaussian prior and homoscedastic Gaussian noise is specified by
+$\begin{cases}Y_{ij}=h(t_{ij},\psi_{i})+\epsilon_{ij},\quad\epsilon_{ij}\sim\mathcal{N}(0,R),\\ \lambda_{i}^{k\to k^{\prime}}(t_{0},t)=\lambda_{0}^{k\to k^{\prime}}(t_{0},t^{\prime})\exp\left(\alpha^{k\to k^{\prime}}g^{k\to k^{\prime}}(t,X_{i},\psi_{i})+\beta^{k\to k^{\prime}}X_{i}\right),\\ \psi_{i}=f(\gamma,X_{i},b_{i}),\quad b_{i}\sim\mathcal{N}(0,Q).\end{cases}$
+Here, $h$ and $f$ define the nonlinear mixed-effects submodel, $g^{k\to k^{\prime}}$ represents the link between biomarker dynamics and the transition intensity, and $\lambda_{0}^{k\to k^{\prime}}$ denotes the baseline hazard associated with transition $(k\to k^{\prime})$.
 This structure extends the classical joint model to arbitrary directed graphs and to both Markovian and semi-Markov specifications.
-The coefficient $\alpha^{k^{\prime}\mid k}$ quantifies the effect of the longitudinal biomarker dynamics on the instantaneous risk of transitioning from state $k$ to state $k^{\prime}$.
-More concretely, for a one-unit increase in $g^{k^{\prime}\mid k}(t,X_{i},\psi_{i})$ (holding other covariates constant), the hazard ratio for transition $(k\to k^{\prime})$ is $\exp(\alpha^{k^{\prime}\mid k})$.
-Likewise, $\beta^{k^{\prime}\mid k}$ represents the effect of baseline (or time-varying) covariates $X_{i}$ on that same transition: a one-unit increase in a covariate $X_{ij}$ multiplies the hazard by $\exp(\beta^{k^{\prime}\mid k}_{j})$.
-Thus, $\alpha^{k^{\prime}\mid k}$ captures the degree to which the biomarker trajectory (via its latent parameter $\psi_{i}$ or function $g^{k^{\prime}\mid k}$) influences the transition risk, while $\beta^{k^{\prime}\mid k}$ captures the direct effect of covariates on that transition’s risk, beyond the biomarker pathway.
+The coefficient $\alpha^{k\to k^{\prime}}$ quantifies the effect of the longitudinal biomarker dynamics on the instantaneous risk of transitioning from state $k$ to state $k^{\prime}$.
+More concretely, for a one-unit increase in $g^{k\to k^{\prime}}(t,X_{i},\psi_{i})$ (holding other covariates constant), the hazard ratio for transition $(k\to k^{\prime})$ is $\exp(\alpha^{k\to k^{\prime}})$.
+Likewise, $\beta^{k\to k^{\prime}}$ represents the effect of baseline (or time-varying) covariates $X_{i}$ on that same transition: a one-unit increase in a covariate $X_{ij}$ multiplies the hazard by $\exp(\beta^{k\to k^{\prime}}_{j})$.
+Thus, $\alpha^{k\to k^{\prime}}$ captures the degree to which the biomarker trajectory (via its latent parameters $\psi_{i}$ or function $g^{k\to k^{\prime}}$) influences the transition risk, while $\beta^{k\to k^{\prime}}$ captures the direct effect of covariates on that transition’s risk, beyond the biomarker pathway.
 3.3.1 Model Variants
-Two baseline hazard conventions are common:
-$\lambda_{0}^{k^{\prime}\mid k}(t\mid T_{im_{i}(t)})=\begin{cases}\lambda_{0}^{k^{\prime}\mid k}(t-T_{im_{i}(t)})&\text{(clock reset)},\\ \lambda_{0}^{k^{\prime}\mid k}(t)&\text{(clock forward)}.\end{cases}$
+Two baseline hazard conventions are common and can be specified as
+$\lambda_{0}^{k\to k^{\prime}}(t_{0},t)=\begin{cases}\lambda_{0}^{k\to k^{\prime}}(t_{0},t)&\text{(clock-reset)},\\ \lambda_{0}^{k\to k^{\prime}}(t)&\text{(clock-forward)}.\end{cases}$
 The clock-reset form models risk as a function of time spent in the current state, whereas the clock-forward form measures risk with respect to global time since study entry.
+Common parametric baseline hazard specifications include the exponential ($\lambda_{0}(t)=\lambda$), the Weibull ($\lambda_{0}(t)=k\lambda^{k}t^{k-1}$), and the Gompertz ($\lambda_{0}(t)=ae^{bt}$) families, among others.
+Furthermore, under the sojourn-time formulation, transition times are no longer constrained to be positive, since only the elapsed duration between successive transitions enters the baseline hazard function.
 3.3.2 Likelihood Formulation
@@
 Censoring and process assumptions C1.
-Censoring times $(C_{i})_{i}$ are mutually independent and noninformative, i.e.
-$C_{i}\mathrel{\perp\!\!\!\perp}(Y_{i},\mathcal{T}_{i}^{*})\mid b_{i},X_{i}$.
+Censoring times $(C_{i})_{i}$ are mutually independent and noninformative, i.e., $C_{i}\mathrel{\perp\!\!\!\perp}(Y_{i},\mathcal{T}_{i}^{*})\mid b_{i},X_{i}$.
 C2.
-Conditionally on $b_{i},X_{i}$, event trajectories satisfy the semi-Markov property 1: $p(\mathcal{T}_{i}^{*}\mid b_{i},X_{i})=\prod_{\ell\geq 0}p\left((T_{i,\ell+1},S_{i,\ell+1})\mid T_{i\ell},S_{i\ell},b_{i},X_{i}\right)$.
-These grouped assumptions mirror those used in classical joint modeling [rizopoulos2012joint] and multi-state survival analysis [putter2007competing]: independence across subjects; the conditional independence structure linking the longitudinal and event submodels through shared random effects; noninformative censoring and (semi-)Markovian dynamics.
-Let $\theta=(\gamma,Q,R,\alpha,\beta)$ denote the vector of model parameters.
-For subject $i$, we observe the longitudinal measurements $Y_{i}=(Y_{i1},\dots,Y_{in_{i}})$ and the event trajectory $\mathcal{T}_{i}=\{(T_{i\ell},S_{i\ell})_{0\leq\ell\leq m_{i}}\}$.
+Conditionally on $b_{i},X_{i}$, event trajectories satisfy the semi-Markov property 1 given $b_{i},X_{i}$.
+These grouped assumptions mirror those used in classical joint modeling [rizopoulos2012joint] and multi-state survival analysis [putter2007competing]: independence across subjects, the conditional independence structure linking the longitudinal and event submodels through shared random effects, noninformative censoring, and semi-Markovian dynamics.
+Let $\theta\coloneqq(\gamma,Q,R,\alpha,\beta)$ denote the vector of model parameters.
+For subject $i\in\llbracket n\rrbracket$, we observe the longitudinal measurements $Y_{i}=(Y_{i1},\dots,Y_{in_{i}})$ and the event trajectory $\mathcal{T}_{i}=\left\{(T_{i\ell},S_{i\ell})_{\ell\leq m_{i}}\right\}$.
 The joint likelihood then factorizes as
-$p(Y_{i},\mathcal{T}_{i}\mid X_{i},C_{i},\theta)=\int p(Y_{i}\mid b_{i},X_{i},\theta)\,p(\mathcal{T}_{i}\mid b_{i},X_{i},C_{i},\theta)\,p(b_{i}\mid\theta)\,db_{i},$
-where $b_{i}$ are individual random effects.
+$p_{\theta}(Y_{i},\mathcal{T}_{i}\mid X_{i},C_{i})=\int p_{\theta}(Y_{i}\mid b_{i},X_{i})\,p_{\theta}(\mathcal{T}_{i}\mid b_{i},X_{i},C_{i})\,p_{\theta}(b_{i})\,db_{i}.$
 This formulation generalizes the joint likelihoods of wulfsohn1997joint and ferrer2016joint to arbitrary multi-state event structures.
-Each part can then be explicitely expressed using Assumptions 3.3.2, yielding an expression very similar to that obtained by rizopoulos2012joint.
+Each part can then be explicitly expressed using Assumptions 3.3.2, yielding an expression very similar to that obtained by rizopoulos2012joint.
 Prior likelihood:
-$p(b_{i}\mid\theta)=\frac{(2\pi)^{-q/2}}{\det(Q)^{1/2}}\exp\left(-\tfrac{1}{2}b_{i}^{T}Q^{-1}b_{i}\right),$
+$p_{\theta}(b_{i})=\frac{(2\pi)^{-q/2}}{\det(Q)^{1/2}}\exp\left(-\tfrac{1}{2}b_{i}^{T}Q^{-1}b_{i}\right),$
 Longitudinal likelihood:
-$p(Y_{i}\mid b_{i},X_{i},\theta)=\prod_{j=1}^{n_{i}}\frac{(2\pi)^{-d/2}}{\det(R)^{1/2}}\exp\left(-\tfrac{1}{2}\left(Y_{ij}-h(t_{ij},\psi_{i})\right)^{T}R^{-1}\left(Y_{ij}-h(t_{ij},\psi_{i})\right)\right),$
+$p_{\theta}(Y_{i}\mid b_{i},X_{i})=\prod_{j=1}^{n_{i}}\frac{(2\pi)^{-d/2}}{\det(R)^{1/2}}\exp\left(-\tfrac{1}{2}\left(Y_{ij}-h(t_{ij},\psi_{i})\right)^{T}R^{-1}\left(Y_{ij}-h(t_{ij},\psi_{i})\right)\right),$
 $\text{with }\psi_{i}=f(\gamma,X_{i},b_{i}).$
 Semi-Markov likelihood:
-$p(\mathcal{T}_{i}\mid b_{i},X_{i},C_{i},\theta)=$ $\prod_{l=0}^{m_{i}-1}p\left((T_{i(l+1)},S_{i(l+1)})\mid b_{i},X_{i},(T_{il},S_{il}),\theta\right)$ (2)
-$\exp\left(-\sum_{s:\,(S_{im_{i}},s)\in E}\Lambda_{i}^{k^{\prime}\mid S_{im_{i}}}(C_{i}\mid b_{i},X_{i},T_{im_{i}},\theta)\right).$
+$p_{\theta}(\mathcal{T}_{i}\mid b_{i},X_{i},C_{i},\theta)=$ $\prod_{\ell=0}^{m_{i}-1}p_{\theta}\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid b_{i},X_{i},(T_{i\ell},S_{i\ell})\right)$ (3)
+$\exp\left(-\sum_{j:(S_{im_{i}},j)\in E}\Lambda_{i,\theta}^{S_{im_{i}}\to j}(T_{im_{i}},C_{i}\mid b_{i},X_{i})\right).$
 The proof of the expression of the Semi-Markov likelihood (3) is provided in Appendix B.
+Each term $p_{\theta}\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid b_{i},X_{i},(T_{i\ell},S_{i\ell})\right)$ can be computed using Equation 2.
 The likelihood factorization above is valid and forms the basis for scalable inference procedures described in Section 4, leveraging the complete trajectory of biomarkers to refine predictions of transitions and survival [de2010mstate].
 Moreover, here, we assume that the initial state $S_{i0}$ is observed. However, the framework could also incorporate a multinomial model for unobserved initial states [yiu2018clustered].
+4 Statistical Inference and Model Selection
+In this section, we outline a practical framework for statistical inference in the proposed multi-state semi-Markov joint model.
+We first derive an optimization scheme for estimating model parameters through a stochastic gradient ascent procedure, and then introduce model selection criteria to assess competing specifications.
+These methods will be illustrated and evaluated on simulated and real datasets in Section 6.
+4.1 Stochastic Gradient Ascent
+The estimation of model parameters $\theta=(\gamma,Q,R,\alpha,\beta)$ relies on two likelihood formulations: the complete-data likelihood $\mathcal{L}_{\mathrm{complete}}$, which includes latent variables, and the marginal likelihood $\mathcal{L}_{\mathrm{marginal}}$, obtained by integrating them out
+$\mathcal{L}_{\mathrm{marginal}}(\theta;X,Y,\mathcal{T},C)=\int\mathcal{L}_{\mathrm{complete}}(\theta;X,Y,\mathcal{T},C,b)\,db,$
+where $b\coloneqq(b_{1},\dots,b_{n})$, and owing to the independence across individuals
+$\mathcal{L}_{\mathrm{complete}}(\theta;X,Y,\mathcal{T},C,b)=\prod_{i=1}^{n}p_{\theta}(Y_{i}\mid b_{i},X_{i})\,p_{\theta}(\mathcal{T}_{i}\mid b_{i},X_{i},C_{i})\,p_{\theta}(b_{i}).$
+It is to be noted that this integral is typically intractable in closed form due to the nonlinear nature of the model.
+Several optimization methods adapted from the nonlinear mixed-effects literature can be applied, including Stochastic EM [kuhn2004coupling], Laplace approximation [wolfinger1993laplace] and Gauss–Hermite quadrature, alongside MCMC-based approximations such as Metropolis–Hastings [hastings1970monte], Metropolis-within-Gibbs, Hamiltonian Monte Carlo [neal2011mcmc].
+These strategies are implemented in software such as JMBayes [jmbayes22024, rizopoulos2020package].
+Another approach requiring mild regularity assumptions on the $\log$ marginal likelihood, but without the need for the model to belong in the exponential family, is to consider a stochastic gradient ascent scheme [caillebotte2025estimation, baey2023efficient] using Fisher’s identity and following the Robbins-Monro procedure [robbins1951stochastic].
+Under interchangeability of integral and differentiation, setting $x\coloneqq(X,Y,\mathcal{T},C)$ for convenience, the Fisher identity writes
+$\nabla_{\theta}\log\mathcal{L}_{\mathrm{marginal}}(\theta;x)=\mathbb{E}_{b\sim p_{\theta}(\cdot\mid x)}\left[\nabla_{\theta}\log\mathcal{L}_{\mathrm{complete}}(\theta;x,b)\right].$
+This expectation is approximated using Monte Carlo samples from the posterior $p_{\theta}(\cdot\mid x)$, avoiding the need to evaluate the intractable marginal likelihood $\mathcal{L}_{\mathrm{marginal}}(\theta;x)$.
+To ensure the covariance matrices $Q$ and $R$ remain symmetric and positive-definite during optimization, we parametrize them using a $\log$-Cholesky decomposition of their inverse.
+By taking the logarithm of the diagonal elements of the Cholesky factors, we map the constrained covariance parameters to an unconstrained real-valued space, and avoid explicit matrix inversion for likelihood computation.
+This formulation naturally supports stochastic gradient ascent with Markov Chain Monte Carlo (MCMC)-based posterior sampling and enables parallelization across individuals, ensuring convergence to a critical point.
+The update rule is detailed in Algorithm 1.
+Algorithm 1 Stochastic gradient ascent model inference
+1:data $x$; initial parameters $\theta^{(0)}$; step sizes $(\eta_{t})_{t\geq 0}$ with $\sum_{t\geq 0}\eta_{t}=+\infty$, $\sum_{t\geq 0}\eta_{t}^{2}<+\infty$; number of parallel chains $K$; MCMC samplers $\{p_{k}^{\mathrm{MCMC}}\}_{k=1}^{K}$; stopping criterion.
+2:
+[…]
+4:while not converged do
+5:  For each $k\in\llbracket K\rrbracket$, draw $b_{k}\sim p_{\theta^{(t)}}(\cdot\mid x)\approx p_{k}^{\mathrm{MCMC}}$.
+6:  $\theta^{(t+1)}\leftarrow\theta^{(t)}+\frac{\eta_{t}}{K}\sum_{k=1}^{K}\nabla_{\theta}\log\mathcal{L}_{\mathrm{complete}}(\theta^{(t)};x,b_{k})$
+7:  $t\leftarrow t+1$
+8:end while
+9:return $\hat{\theta}$
+Preconditioning matrices may also be used, such as the Fisher Information Matrix in a natural gradient ascent framework, or even specialized optimizers such as Adam [kingma2014adam], NAdam [dozat2016incorporating] or Adagrad [duchi2011adaptive].
+In particular, the Adam optimizer will later be used in our numerical experiments.
 4.2 Model Selection and Information Criteria
-To compare competing model specifications, we rely on information criteria based on the marginal likelihood as defined in Section 4.
-The Akaike Information Criterion (AIC) [akaike2003new] and the Bayesian Information Criterion (BIC) [schwarz1978estimating] are standard tools.
-While the AIC relies on an asymptotically unbiased estimator of the $\log$-likelihood, the BIC aims to identify the true model with great probability as the sample size grows.
+To compare competing model specifications, one typically relies on information criteria based on the marginal likelihood.
+In practice, competing models are compared by selecting the specification that yields the smallest value of the chosen information criterion.
+For instance, the Akaike Information Criterion (AIC) [akaike2003new] and the Bayesian Information Criterion (BIC) [schwarz1978estimating] are standard tools.
+While the AIC relies on an asymptotically unbiased estimator of the $\log$-likelihood, the BIC aims to identify the true model with high probability as the sample size grows.
 The Akaike Information Criterion is computed as follows
-$\mathrm{AIC}\coloneqq-2\log\mathcal{L}_{\text{marginal}}(\hat{\theta};X,Y,\mathcal{T},C)+2k,$
-where $k$ is the number of parameters, and $\hat{\theta}$ is a maximizer of the marginal likelihood.
+$\mathrm{AIC}\coloneqq-2\log\mathcal{L}_{\mathrm{marginal}}(\hat{\theta};X,Y,\mathcal{T},C)+2\dim\hat{\theta}$,
+where $\dim\hat{\theta}$ is the number of parameters, and $\hat{\theta}$ is a maximizer of the marginal likelihood.
 The Bayesian Information Criterion is similarly defined but imposes a stronger penalty for model complexity, growing in $\log n$ where $n$ is the number of observations.
@@
 To alleviate this problem, we recall the derivation of the BIC from the Laplace approximation [lebarbier2006introduction].
-For a prior distribution $\pi$ on a set of models $\mathcal{M}$, the Laplace approximation approaches up to constant terms which could also be accounted to improve the approximation, the posterior probability given observed data $x$ defined in Section 4 by
+For a prior distribution $\pi$ on a set of models $\mathcal{M}$, the Laplace approximation yields, up to constant terms (which could also be included to improve the approximation), the approximation of the posterior probability given observed data $x$ reads
 […]
@@
 […]
-In both cases, one should aim at minimizing the chosen criterion.
-4 Statistical Inference
-The estimation of model parameters $\theta=(\gamma,Q,R,\alpha,\beta)$ relies on two likelihood formulations: the complete-data likelihood $\mathcal{L}_{\text{complete}}$, which includes latent variables, and the marginal likelihood $\mathcal{L}_{\text{marginal}}$, obtained by integrating them out
-$\mathcal{L}_{\text{marginal}}(\theta;X,Y,\mathcal{T},C)=\int\mathcal{L}_{\text{complete}}(\theta;X,Y,\mathcal{T},C,b)\,db,$
-where
-$\mathcal{L}_{\text{complete}}(\theta;X,Y,\mathcal{T},C,b)=\prod_{i=1}^{n}p(Y_{i}\mid b_{i},X_{i},\theta)\,p(\mathcal{T}_{i}\mid b_{i},X_{i},C_{i},\theta)\,p(b_{i}\mid\theta).$
-It is to be noted that this integral is typically intractable in closed form due to the nonlinear nature of the model, necessitating numerical approximation methods such as Monte Carlo integration or Laplace approximation.
-Several optimization methods adapted from the nonlinear mixed-effects literature can be applied, including Stochastic EM [kuhn2004coupling], Laplace approximation [wolfinger1993laplace] and Gauss–Hermite quadrature.
-These strategies are implemented in software such as JMBayes [jmbayes22024, rizopoulos2020package].
-Another approach requiring mild regularity assumptions on the $\log$ marginal likelihood, but without the need for the model to belong in the exponential family, is to consider a stochastic gradient ascent scheme [caillebotte2025estimation, baey2023efficient] using Fisher’s identity and following the Robbins-Monro procedure [robbins1951stochastic], alongside with MCMC-based approximations such as Metropolis-Hastings and Hamiltonian Monte Carlo.
-Under interchangeability of integral and differentiation, setting $x\coloneqq(X,Y,\mathcal{T},C)$ for convenience, the Fisher identity writes
-$\nabla_{\theta}\log\mathcal{L}_{\text{marginal}}(\theta;x)=\mathbb{E}_{b\sim p(\cdot\mid x,\theta)}\left(\nabla_{\theta}\log\mathcal{L}_{\text{complete}}(\theta;x,b)\right).$
-This expectation is approximated using Monte Carlo samples from the posterior $p(\cdot\mid x,\theta)$, avoiding the need to evaluate the intractable marginal likelihood $\mathcal{L}_{\text{marginal}}(\theta;x)$.
-This formulation naturally supports stochastic gradient ascent with MCMC-based posterior sampling and is amenable to minibatch parallelization across individuals, ensuring the convergence to a critical point.
-The update rule is as follows:
-Algorithm 1 Stochastic gradient ascent model inference
-1:data $x$; initial parameter $\theta^{(0)}$; step sizes $(\eta_{t})_{t\geq 0}$ with $\sum_{t}\eta_{t}=+\infty$, $\sum_{t}\eta_{t}^{2}<+\infty$; batch size $K$; sampler $p_{t}^{\mathrm{MCMC}}$; stopping criterion.
-2:
-3:while not converged do
-4:  Draw $b_{k}\sim p(\cdot\mid x,\theta_{n})\approx p_{n}^{\text{MCMC}}$.
-5:  $\theta^{(t+1)}\leftarrow\theta^{(t)}+\frac{\eta_{t}}{K}\sum_{k=1}^{K}\nabla_{\theta}\log\mathcal{L}_{\text{complete}}(\theta^{(t)};x,b_{k})$.
-6:end while
-7:return $\hat{\theta}$
-Preconditioning matrices may also be used, such as the Fisher Information Matrix in a natural gradient descent framework, or even specialized optimizers such as Adam [kingma2014adam], NAdam [dozat2016incorporating] or Adagrad [duchi2011adaptive].
-In particular, the Adam optimizer will later be used in our numerical experiments.
+However, as previously noted, the marginal $\log$-likelihood
+$\log\mathcal{L}_{\mathrm{marginal}}(\theta;x)=\log\int\mathcal{L}_{\mathrm{complete}}(\theta;x,b)\,db$
+is generally intractable in mixed effects and joint models, since the integral with respect to the latent variables does not admit a closed form.
+Although the gradient of the marginal $\log$-likelihood can be efficiently estimated using Fisher’s identity, the marginal $\log$-likelihood itself cannot be evaluated directly.
+One possibility is to estimate it using bridge sampling, which has been proven to provide consistent and accurate estimators of normalizing constants and marginal likelihoods based on MCMC output [meng1996simulating, gronau2017tutorial].
+An alternative consists in approximating the posterior distribution $p_{\hat{\theta}}(b\mid x)$ by a Gaussian distribution, for instance via a Laplace approximation. Using the entropy identity,
+$\log\mathcal{L}_{\mathrm{marginal}}(\hat{\theta};x)=\mathbb{E}_{b\sim p_{\hat{\theta}}(\cdot\mid x)}\left[\log\mathcal{L}(\hat{\theta};x,b)\right]-\mathbb{E}_{b\sim p_{\hat{\theta}}(\cdot\mid x)}\left[\log p_{\hat{\theta}}(b\mid x)\right],$
+we can further leverage independence across individuals, so that the entropy term decomposes as
+$\mathbb{E}_{b\sim p_{\hat{\theta}}(\cdot\mid x)}\left[\log p_{\hat{\theta}}(b\mid x)\right]=\sum_{i=1}^{n}\mathbb{E}_{b_{i}\sim p_{\hat{\theta}}(\cdot\mid x_{i})}\left[\log p_{\hat{\theta}}(b_{i}\mid x_{i})\right].$
+Approximating each posterior $p_{\hat{\theta}}(\cdot\mid x_{i})$ by a Gaussian distribution with covariance matrix $\Sigma_{i}$, we obtain
+$\log\mathcal{L}_{\mathrm{marginal}}(\hat{\theta};x)\approx\mathbb{E}_{b\sim p_{\hat{\theta}}(\cdot\mid x)}\left[\log\mathcal{L}(\hat{\theta};x,b)\right]+\sum_{i=1}^{n}\underbrace{\frac{1}{2}\left(\log\det\Sigma_{i}+q\left(\log(2\pi)+1\right)\right)}_{H_{i}},$
+where $q$ denotes the dimension of the latent vector.
+This approximation replaces the intractable posterior entropy with a sum of closed-form entropies of $q$-dimensional Gaussian distributions.
+Furthermore, each term $H_{i}$ depends only on its corresponding covariance matrix $\Sigma_{i}$, which can be easily estimated from MCMC samples of the latent variables using the empirical covariance estimator.
+The practical relevance of the proposed model selection criteria is discussed in Subsection 6.3.
 5 Dynamic prediction
-The simulation of individual trajectories using Algorithm 2 enables prediction in the multi-state joint modeling framework.
+In this section, we turn to dynamic prediction within the multi-state framework.
+Although the prediction of longitudinal trajectories does not fundamentally differ from the standard nonlinear joint modeling setting, the key difference arises from the structure of the event process.
+Indeed, because the joint distribution of multi-state trajectories has no tractable closed-form expression, direct calculation of predictive quantities is generally infeasible.
+To address this, we rely on simulation-based procedures which are rendered possible using Algorithm 3 (detailed in the Appendix).
 Specifically, predictions are derived from simulated event trajectories, allowing for individualized risk assessment and forecasting of future states or event times based on a subject’s observed biomarker and event history.
 This approach generalizes classical dynamic prediction by leveraging the rich structure of multi-state processes and the joint distribution of longitudinal and event data.
-Suppose we are interested in some functional of the true (unobserved) future trajectory,
+Let $i$ index a new individual. Suppose we are interested in some functional of the true (unobserved) future trajectory,
 $\chi^{*}(\mathcal{T}_{i}^{*})\in\mathbb{Y},$
-where $\mathbb{Y}$ is some generic space.
+where $\mathbb{Y}$ is some generic vector space.
 For example, one can think of the state taken by the individual $i$ at a time $u$.
-As is the case in traditional dynamic prediction however, we are only interested in predicting quantities or characteristics that depend on a yet unobserved future, given prior information, i.e.
-the trajectory up to some prediction time $t\leq u$ and observed longitudinal markers.
-Contrary to (single transition) joint models, the conditional probability distribution on $(\mathbb{R}\times\mathcal{S})^{\mathbb{N}}$ cannot be analytically derived.
+As is the case in traditional dynamic prediction however, we are only interested in predicting quantities or characteristics that depend on a yet unobserved future, given prior information, i.e., the trajectory $\mathcal{S}_{i}(t)$ up to some prediction time $t\leq u$ and observed longitudinal markers $\mathcal{Y}_{i}(t)\coloneq\left\{(t_{ij},Y_{ij}):t_{ij}\leq t\right\}$.
+Contrary to (single transition) joint models, the conditional probability distribution on $(\mathbb{R}^{+}\times V)^{\mathbb{N}}$ cannot be analytically derived.
 Nonetheless, as seen in Algorithm 3, we are able to accurately simulate each trajectory.
-Therefore, a Monte Carlo estimation scheme may be devised under certain restrictions, where we estimate $\chi^{*}(\mathcal{T}_{i}^{*})$ by its expected value:
-$\hat{\chi}_{i}=\coloneqq\mathbb{E}_{\mathcal{T}^{\prime}_{i}\sim\mathscr{L}(\mathcal{T}_{i}^{*}\mid\mathcal{Y}(t),\mathcal{T}_{i},t)}\left(\chi^{*}(\mathcal{T}^{\prime}_{i})\right)\approx\frac{1}{B}\sum_{k=1}^{B}\chi^{*}(\mathcal{T}_{i}^{(k)}),\,\mathcal{T}_{i}^{(k)}\overset{\text{i.i.d.}}{\sim}\mathscr{L}(\mathcal{T}_{i}^{*}\mid\mathcal{Y}(t),\mathcal{T}_{i},t).$
-In general however, $\chi^{*}$ may very well depend on (countably) infinitely many transitions.
-As a result, since the proposed estiamtion relies on simulted samples, we require the quantity to depend only on a finite subset of transitions.
+Therefore, a single or double Monte Carlo estimation scheme may be devised under certain restrictions, where we estimate $\chi^{*}(\mathcal{T}_{i}^{*})$ by its conditional expected value (or other statistics depending on the posterior distribution, such as the median)
+$\hat{\chi}_{i}^{(1)}\coloneqq\mathbb{E}_{\hat{\theta}}\left[\chi^{*}(\mathcal{T}^{\prime}_{i})\mid\mathcal{Y}_{i}(t),\mathcal{S}_{i}(t)\right]\approx\frac{1}{B}\sum_{k=1}^{B}\chi^{*}(\mathcal{T}_{i}^{(k)}),\,\mathcal{T}_{i}^{(k)}\overset{\mathrm{i.i.d.}}{\sim}\mathscr{L}_{\hat{\theta}}\left(\mathcal{T}_{i}^{*}\mid\mathcal{Y}(t),\mathcal{S}_{i}(t)\right),$
+where the samples $\{\mathcal{T}_{i}^{(k)}\}_{k=1}^{B}$ are generated using a MCMC algorithm targeting the distribution $\mathscr{L}_{\hat{\theta}}\left(b_{i}\mid\mathcal{Y}_{i}(t),\mathcal{S}_{i}(t)\right)$ in conjunction with Algorithm 3.
+Another approach from Bayesian perspective treats the model parameters $\theta$ as random variables by imposing a prior distribution together with the training data $\mathcal{D}$ [rizopoulos2011dynamic], and thus in this case our new estimator may be defined
+$\hat{\chi}_{i}^{(2)}\coloneqq\mathbb{E}_{\theta^{\prime}\sim\mathscr{L}(\theta\mid\mathcal{D})}\left[\mathbb{E}_{\theta^{\prime}}\left[\chi^{*}(\mathcal{T}^{\prime}_{i})\mid\mathcal{Y}(t),\mathcal{S}_{i}(t)\right]\right],$
+and approximated by a double Monte Carlo scheme accordingly where $\mathscr{L}(\theta\mid\mathcal{D})$ is the posterior distribution of $\theta$.
+In the case of longitudinal values, the procedure simplifies: for each draw of $b_{i}$, one usually predicts a single deterministic functional by substituting $\chi$ with $h(\cdot,\psi_{i})$, which already corresponds to the conditional mean of $Y_{i}$ given $b_{i}$.
+However, in general multi-state prediction, $\chi^{*}$ may very well depend on (countably) infinitely many transitions.
+As a result, since the proposed estimation relies on simulated samples, we require the quantity to depend only on a finite subset of transitions.
 Assumption 1.
-Assume there exists $\chi:\bigcup_{n\geq 1}(\mathbb{R}\times\mathbb{S})^{n}\to\mathbb{Y}$ and $\tau_{i}$ a stopping time for the filtration $\mathcal{F}_{in}\coloneqq\sigma\left((T_{il},S_{il})_{0\leq l\leq n}\right)$ such that:
-$\begin{cases}\tau_{i}<+\infty\text{ a.s.},\\ \chi^{*}(\mathcal{T}_{i}^{*})=\chi\left((T_{il},S_{il})_{1\leq l\leq\tau_{i}}\right)\end{cases}$
-Essentially, given Assumption 1, with probability one we are guaranteed that the quantity of interested may be computed in a finite number of simulation steps.
-Indeed, the prediction algorithm for a single sample may be summarized as follows:
+There exists $\chi:\bigcup_{n\geq 1}(\mathbb{R}^{+}\times V)^{n}\to\mathbb{Y}$ and $\tau_{i}$ a stopping time for the filtration $\mathcal{F}_{in}\coloneqq\sigma\left((T_{il},S_{il})_{\ell\leq n}\right)$ such that
+[…]
+Essentially, given Assumption 1, with probability one we are guaranteed that the quantity of interest may be computed in a finite number of simulation steps.
+Indeed, the prediction algorithm for a single sample may be summarized as in Algorithm 2.
+Algorithm 2 Prediction algorithm
 […]
@@
 […]
-4:while $n<\tau_{i}$ do
-5:  Append simulated $(T_{in},S_{in})$ to $\mathcal{T}_{i}$ if $n>\operatorname{len}(\mathcal{T}_{i})$
-6:  $n\leftarrow n+1$
+4:while $\ell\leq\tau_{i}$ do
+5:  Append simulated $(T_{i\ell},S_{i\ell})$ to $\mathcal{T}_{i}$
+6:  $\ell\leftarrow\ell+1$
 7:end while
-8:return $\chi(\mathcal{T}_{i})$
+8:return $\chi\left((T_{il},S_{il})_{\ell\leq\tau_{i}}\right)$
 Numerous quantities of interest may be encompassed by this framework.
@@
 First, for a directed acyclic graph $G=(V,E)$, we define the depth of $G$, denoted by $\operatorname{depth}(G)$, as the length of the longest directed path in $G$, that is,
-$\operatorname{depth}(G)\coloneqq\max_{(v_{0},v_{1},\dots,v_{k})}k,$
-where $(v_{0},v_{1},\dots,v_{k})$ ranges over all directed paths in $G$.
+$\operatorname{depth}(G)\coloneqq\max_{(v_{0},\dots,v_{k})}k,$
+where $(v_{0},\dots,v_{k})$ ranges over all directed paths in $G$.
 Example 1 (State at time $u$).
-Let $G=(V,E)$ be a directed acyclic graph, $u\in\mathbb{R}$ be a fixed time, and $\mathbb{Y}=V$ be the set of possible states.
+Let $G=(V,E)$ be a finite directed acyclic graph, $u\in\mathbb{R}^{+}$ be a fixed time, and $\mathbb{Y}=\Delta(V)$ be the simplex of distributions over states.
 Let $\tau_{i}=\inf\{n\in\mathbb{N}:T_{in}\geq u\text{ or }S_{in}\text{ is absorbing}\}$.
 Clearly, $\tau_{i}\leq\operatorname{depth}(G)<+\infty$.
-Then, take $\chi_{u}^{*}(\mathcal{T}_{i}^{*})=S_{i\tau_{i}}$, which corresponds to the state of the individual $i$ at time $u$.
-In particular, if $V=\{0,1\}$ and $E=\{(0,1)\}$, we recover the special case used for survival probability estimation in standard joint models [rizopoulos2011dynamic].
+For each state $k\in V$, let $\mathds{1}_{k}\in\mathbb{Y}$ denote the vector that is $1$ at entry $k$ and 0 elsewhere (a one-hot encoding of $k$).
+Then, we set $\chi_{u}^{*}(\mathcal{T}_{i}^{*})=\mathds{1}_{S_{i\tau_{i}}}\in\mathbb{Y}$ to represent the state occupied by individual $i$ at time $u$.
+In particular, if $V=\{0,1\}$ and $E=\left\{(0,1)\right\}$, we recover the special case used for survival probability estimation in standard joint models [rizopoulos2011dynamic].
 Example 2 (Hitting time).
-Let $G=(V,E)$ be a directed acyclic graph, $\mathbb{A}\subset V$ a non-empty subset of states, and $\mathbb{Y}=\mathbb{R}\cup\{+\infty\}$.
-For any two non-empty sets $\mathbb{A},\,\mathbb{B}\subset V$, we note $\mathbb{A}\rightsquigarrow\mathbb{B}$ if there exists a path from $\mathbb{A}$ to $\mathbb{B}$ in $G$.
+Let $G=(V,E)$ be a finite directed acyclic graph, $\mathbb{A}\subseteq V$ a non-empty subset of states, and $\mathbb{Y}=\bar{\mathbb{R}}^{+}$.
+For any two non-empty sets $\mathbb{A},\,\mathbb{B}\subseteq V$, we note $\mathbb{A}\rightsquigarrow\mathbb{B}$ if there exists a path from $\mathbb{A}$ to $\mathbb{B}$ in $G$.
 Let $\tau_{i}=\inf\{n\in\mathbb{N}:S_{in}\in\mathbb{A}\vee\{S_{in}\}\not\rightsquigarrow\mathbb{A}\}$.
-Then, $\tau_{i}\leq\operatorname{depth}(G)<+\infty$ and take $\chi_{\mathbb{A}}^{*}(\mathcal{T}_{i}^{*})=T_{i\tau_{i}}+\mathds{1}_{\{S_{in}\}\not\rightsquigarrow\mathbb{A}}(+\infty)$ represents the hitting time for the set $\mathbb{A}$.
-In Example  1, the stopping time $\tau_{i}$ captures the step at which the process for individual $i$ reaches or exceeds a given time $u$, or enters an absorbing state.
+Then, $\tau_{i}\leq\operatorname{depth}(G)<+\infty$ and we set $\chi_{\mathbb{A}}^{*}(\mathcal{T}_{i}^{*})=T_{i\tau_{i}}+\mathds{1}_{\{S_{i\tau_{i}}\}\not\rightsquigarrow\mathbb{A}}(+\infty)$ represents the hitting time for the set $\mathbb{A}$.
+In Example 1, the stopping time $\tau_{i}$ captures the step at which the process for individual $i$ reaches or exceeds a given time $u$, or enters an absorbing state.
 The resulting value $\chi^{*}(\mathcal{T}_{i}^{*})=S_{i\tau_{i}}$ therefore represents the state of the individual at that specific time.
@@
 It enables individualized forecasting of future state occupancy, event risks, sojourn times, and other clinically relevant outcomes, fully exploiting the subject’s observed biomarker trajectory and event history.
-In Section 7, we demonstrate its application to forecasting dependency trajectories in the PAQUID cohort.
+In Section 7, we demonstrate its application to forecasting dependency trajectories in the PAQUID cohort (see Subsection 7.3.2).
 6 Simulation Study
@@
 We simulate data from a three-state semi-Markov process coupled with a nonlinear longitudinal biomarker trajectory, representing a simplified disease progression model.
-This setup allows us to assess both statistical accuracy and computational scalability of the stochastic-gradient estimation procedure.
-All simulations, estimations, and figures reported in this section were produced using the open-source jmstate Python package, available on PyPI, which we developed to implement the multi-state joint modeling framework and inference algorithms described in Sections 3.2 and 4.
+This setup allows us to assess both statistical estimation error (see Table 5) and computational scalability of the stochastic gradient estimation procedure.
+We further evaluate the model selection performance of the proposed methodology, as outlined in Subsection 4.2.
+To ensure statistical reliability, we perform $n_{\mathrm{runs}}=100$ independent simulations on datasets, comprised of $n=500$ individuals, generated from the same underlying distribution.
+All simulations, estimations, and figures reported in this section were produced using the open-source jmstate Python package version 0.15.2, available on PyPI, which we developed to implement the multi-state joint modeling framework and inference algorithms described in Sections 3.2 and 4.
 6.1 Model Specification
-The simulated model involves three states—Healthy (0), Sick ($1$), and Terminal (22)—connected by the transitions $0\rightarrow 1$, $0\rightarrow 2$, and $1\rightarrow 2$ (Figure 3).
-This configuration captures the typical monotone evolution of a disease process, with both direct and indirect paths to the terminal state.
-Transition intensities follow an exponential baseline hazard under a clock-reset convention.
-The longitudinal process follows a continuous piecewise-affine regression:
-$h(t,\psi)=\psi_{1}+\psi_{2}t+\mathds{1}_{t>\tau}(\psi_{3}-\psi_{2})(t-\tau),$
-with a known inflection time $\tau=6$.
-The individual-specific parameters $\psi_{i}=(\psi_{i1},\psi_{i2},\psi_{i3})$ arise from a nonlinear mixed-effects model with Gaussian random effects $b_{i}\sim\mathcal{N}(0,Q)$ and independent Gaussian noise $\epsilon_{ij}\sim\mathcal{N}(0,R)$.
-The link function combines the current biomarker value and its slope,
-$g(t,X_{i},\psi_{i})=\big(h(t,\psi_{i}),\frac{\partial h}{\partial t}(t,\psi_{i})\big),$
-shared across transitions. Covariates $X_{i}$ are normally distributed and enter linearly through $\beta^{k^{\prime}\mid k}X_{i}$.
-Longitudinal data were collected at up to $20$ random time points per subject and censored at $C_{i}\sim\mathcal{U}([10,15])$.
-Censoring also truncates the biomarker trajectories.
-The true parameter vector lies in $\mathbb{R}^{16}$.
-Each simulated dataset comprised $n=1000$ individuals, generating approximately balanced transition counts (Figure 4).
-This trajectory mimics a biomarker whose evolution slows after treatment or onset of symptoms, while preserving continuity at the change point.
-Figure 3: State transition diagram of the simulated three-state model.
-6.2 Estimator Convergence and Accuracy
-Parameter estimation was performed using the stochastic-gradient inference procedure of Section 4, with the Adam optimizer (learning rate $0.5$, batch size $K=100$) and an adaptive stopping criterion based on the first and second moments of parameter differences.
-Convergence was typically achieved in 400400500500 iterations and required around $12$ seconds per run on an AMD Ryzen 5 processor.
-Across $n_{\text{runs}}=100$ independent replications, the mean relative bias of all parameters remained below $2\%$, and the root-mean-square error (RMSE) scaled approximately as $n^{-1/2}$.
-Parameters governing the biomarker trajectory $(\gamma,Q,R)$ were consistently well identified, while association coefficients $(\alpha^{k^{\prime}\mid k},\beta^{k^{\prime}\mid k})$ exhibited slightly higher variability for rare transitions.
-Representative convergence trajectories are shown in Figure 10, and quantitative results are summarized in Table 5.
-Furthermore, longitudinal measurements were taken at $m_{i}\leq 20$ time points, and censoring times were drawn from $C_{i}\sim\mathcal{U}\left([10,15]\right)$.
-Longitudinal observations were also truncated at the censoring times.
+The simulated model represents a toy example based on a pharmacological scenario where drug absorption can influence seizure activity and patient stability.
+The model involves two states: Seizure ($1$) and Stable ($2$), which is absorbing.
+Transitions occur both from $1\rightarrow 1$ and $1\rightarrow 2$ (Figure 3), reflecting how a patient may either experience repeated seizure episodes or stabilize after treatment.
+If the drug is well absorbed, it helps mitigate seizure activity and promotes stabilization.
+Conversely, poor absorption increases the likelihood of recurrent seizures, highlighting the importance of pharmacokinetics in disease control.
+Transition intensities follow an exponential baseline hazard.
+Figure 3: State transition diagram of the simulated recurrent two-state model.
+The longitudinal biomarker trajectory is modeled using a bi-exponential function:
+$h(t,\psi)=\psi_{1}\left(e^{-\psi_{2}t}-e^{-\psi_{3}t}\right),\quad\psi_{i}=\gamma e^{b_{i}},\quad b_{i}\sim\mathcal{N}\left(0,\mathrm{diag}(Q_{1},Q_{2},Q_{3})\right),$
+and the observation noise is assumed Gaussian, independent across individuals and time points.
+This bi-exponential structure is commonly used in pharmacokinetic and longitudinal modeling (e.g., kerioui2022modelling), although alternative parametrizations are sometimes employed.
+Incidentally, one of the central points in kerioui2022modelling was the deliberate exclusion of recurrent transitions, regarded as difficult to estimate reliably.
+We show empirically that such transitions can in fact be estimated within our framework.
+The link functions are applied consistently across all transitions and defined as
+$g^{1\to 1}(t,\psi)=g^{1\to 2}(t,\psi)=\int_{0}^{t}h(s,\psi)\,ds,$
+integrating the current biomarker value, thus providing a cumulative measure. Covariates $X_{i}$ are normally distributed and enter linearly via $\beta^{k\to k^{\prime}}X_{i}$.
+Longitudinal data were collected at $20$ equally spaced time points between $t=0$ and $t=15$ per subject, with censoring at $C_{i}\sim\mathcal{U}([10,15])$ alongside the individual trajectories.
+Since the base hazard rates $\lambda^{1\to 1}$ and $\lambda^{1\to 2}$ are jointly optimized with the other parameters, the final parameter vector lies in $\mathbb{R}^{13}$, and its values used in simulation are shown in Table 5.
+Each simulated dataset comprised a moderate number of $n=500$ individuals, generating roughly balanced transition counts (Figure 4).
+Thanks to careful parametrization, almost all parameters are mapped to the whole real axis unconstrained.
 A short summary of the longitudinal process as well as the trajectories is given by the Figure 4 below.
-From $\rightarrow$ To # Observations
-$0\rightarrow 1$ 613
-$0\rightarrow 2$ 375
-$1\rightarrow 2$ 592
-Figure 4: Simulated data: on the left, a sample of longitudinal measurements from $50$ individuals; on the right, observed transitions between states for the complete population of $1000$ individuals.
+Note that unlike traditional joint modeling, the notion of censored transitions does not apply here.
+If a transition does not occur within a trajectory $\mathcal{T}_{i}$, it is simply omitted from the representation.
+Figure 4: Simulated data: on the left, a sample of longitudinal measurements from $50$ individuals; on the right, observed transitions histograms for the complete population.
+$76$ patients experienced repeated seizure.
 6.2 Estimator Convergence
-To illustrate the convergence of the estimator, we examine the optimization process for a particular run with $n=1000$ (see Figure 10), as well as RMSE values computed from $n_{\text{runs}}=100$ independent runs on different simulated datasets, all generated from the same underlying distribution.
-The results of these simulations are shown in Table 5.
-To fit the model, we first have to define initial parameters, which can be particularly important for the optimization process if the model is not well-behaved.
-In the present case, without prior knowledge of the true parameters, we zero the values and use the identity matrix for both covariance matrices.
-Formally:
-$m_{i}^{(t)}\leftarrow\beta_{i}m_{i}^{(t-1)}+(1-\beta_{i})(\theta^{(t)}-\theta^{(t-1)})^{i},\quad\hat{m}_{i}^{(t)}=\frac{m_{i}^{(t)}}{1-\beta_{i}^{t}},$
-and the optimization process is terminated when
-$|\hat{m}_{1}^{(t)}|\leq 10^{-6}+10^{-1}\sqrt{\hat{m}_{2}^{(t)}}.$
-The Figure 10 shows the evolution of the parameters during the optimization process, and Table 5 summarizes the accuracy of the inferred parameters.
-Parameter True value Mean Standard error RMSE
-$\gamma_{1}$ 2.5000 2.4960 0.0370 0.0372
-$\gamma_{2}$ -1.3000 -1.3039 0.0257 0.0260
-$\gamma_{3}$ 0.2000 0.1936 0.0267 0.0275
-$\tilde{Q}_{1}$ 0.2554 0.2243 0.0442 0.0540
-$\tilde{Q}_{2}$ 0.8047 0.8009 0.0260 0.0263
-$\tilde{Q}_{3}$ 0.6020 0.6001 0.0276 0.0277
-$\tilde{R} -$0.2653 -0.2684 0.0083 0.0089
-$\alpha_{1}^{1\mid 0}$ -0.5000 -0.49997 0.0486 0.0486
-$\alpha_{2}^{1\mid 0}$ -3.0000 -2.9962 0.0800 0.0801
-$\alpha_{1}^{2\mid 0}$ -1.0000 -0.9990 0.0847 0.0847
-$\alpha_{2}^{2\mid 0}$ -5.0000 -4.9897 0.1191 0.1196
-$\alpha_{1}^{2\mid 1}$ 0.0000 0.0011 0.0327 0.0327
-$\alpha_{2}^{2\mid 1}$ -1.2000 -1.2045 0.0429 0.0432
-$\beta^{1\mid 0}$ -1.3000 -1.3015 0.0501 0.0501
-$\beta^{2\mid 0}$ -0.9000 -0.8982 0.0670 0.0670
-$\beta^{2\mid 1}$ -0.7000 -0.6981 0.0551 0.0551
-Figure 5: Comparison of true and estimated parameters (averaged over 100 runs).
-The simulation confirms that the proposed inference method accurately recovers all parameters of the multi-state joint model under realistic sample sizes.
-The optimization demonstrates stable convergence (Figure 10), with minimal sensitivity to initialization.
-The algorithm’s computational efficiency allows for extensive simulation or cross-validation studies at negligible cost.
+For each run, model parameters were estimated using the stochastic gradient ascent procedure described in Section 4, using a Metropolis-within-Gibbs sampling algorithm.
+Although the initial parameter values can strongly influence the optimization process, we set $\gamma^{(0)}=(1,1,1)^{T}$ to prevent division by zero, initialize both covariance matrices $Q$ and $R$ as identity matrices, set $\lambda^{1\to 1}$ and $\lambda^{1\to 2}$ to $1$, and initialize all linear coefficients to zero, thereby avoiding any assumptions about prior knowledge of the parameters.
+We used the Adam optimizer alongside an adaptive stopping criterion based on a local linear trend diagnostic applied component-wise to the parameters optimization history.
+Formally, given a fixed window size $T\geq 2$ and any $t\geq T-1$, let
+$\{\theta_{j}^{(s)}\}_{s=t-T+1}^{t}$
+denote the last $T$ iterates of the $j$-th component. We consider the linear model with intercept
+$\theta_{j}^{(s)}=a_{j}+c_{j}s+\varepsilon_{j}^{(s)},\qquad s\in\{t-T+1,\dots,t\},$
+and compute the associated coefficient of determination $R_{j,t}^{2}$.
+Since in such a simple linear regression, the $R^{2}$ coefficient coincides with the squared correlation between the response and the regressor, small values of $R_{j,t}^{2}$ indicate that the recent evolution of $\theta_{j}$ is not better explained by a linear trend than by a constant model, and is therefore consistent with local stationarity.
+The optimization procedure is stopped as soon as
+$\frac{1}{\dim\theta^{(t)}}\sum_{j=1}^{\dim\theta^{(t)}}R_{j,t}^{2}\leq\delta,$
+where $\delta\in[0,1]$ is a prescribed threshold (in practice $\delta=0.1$).
+With a relatively high value of $T=100$, this criterion that over the last $T$ iterations, no component displays a statistically meaningful linear drift, so the parameter vector is locally indistinguishable from a constant trajectory.
+In our findings, it proved to be robust in spite of the stochastic nature of the optimization path, and is scale-invariant.
+To illustrate the convergence of the estimator, we examine the optimization process for a particular run in Figure 10, as well as RMSE values computed from $n_{\mathrm{runs}}=100$, as shown in Table 5.
+Parameter True value Bias Std. deviation RMSE
+$\gamma_{1}$ 2.000 -0.018 0.174 0.175
+$\gamma_{2}$ 0.200 0.012 0.120 0.121
+$\gamma_{3}$ 1.000 -0.009 0.101 0.101
+$\frac{1}{2}\log(Q_{1})$ 0.949 -0.017 0.159 0.160
+$\frac{1}{2}\log(Q_{2})$ 1.498 -0.035 0.237 0.240
+$\frac{1}{2}\log(Q_{3})$ 1.151 -0.040 0.264 0.267
+$\frac{1}{2}\log(R)$ 1.498 0.000 0.009 0.009
+$\log\lambda^{1\to 1}$ -1.609 -0.009 0.152 0.152
+$\log\lambda^{1\to 2}$ -4.605 -0.035 0.176 0.180
+$\alpha_{1}^{1\to 1}$ -1.000 0.007 0.148 0.148
+$\alpha_{1}^{1\to 2}$ 0.500 -0.010 0.144 0.144
+$\beta^{1\to 1}$ -1.000 0.025 0.153 0.155
+$\beta^{1\to 2}$ 0.500 0.018 0.143 0.144
+Figure 5: Comparison of true and estimated parameters over $100$ runs with $n=500$ simulated individuals (three decimal places).
+Since the covariance matrix $Q$ is assumed diagonal, the $\log$-Cholesky reduces to the vector of negative $\log$-standard deviations.
+The simulation results indicate that the proposed estimation procedure accurately recovers all parameters of the multi-state joint model for moderate sample sizes.
+Across the $100$ replications, the biases remain remarkably low, and the RMSE values are small compared to the magnitude of the corresponding parameters.
+Moreover, the observed variability is probably largely attributable to the intrinsic sampling variability of the true maximum likelihood estimator rather than to optimization error.
+The population-level parameters $\gamma$ and the measurement noise parameter $R$ display comparatively lower dispersion, indicating stronger sensitivity for these components within the model.
+The optimization procedure exhibits stable convergence (Figure 10) and further supports that the interpretation that residual fluctuations across replications are primarily driven by the variance of the estimator itself.
+6.3 Model Selection
+To assess the performance of the information criteria described in Subsection 4.2, we compared the Akaike Information Criterion and the modified $\mathrm{BIC}\mathcal{I}$.
+Following the same procedure as in the rest of the current section, we performed $n_{\mathrm{runs}}=100$ independent simulations, generating for each run a dataset of size $n=500$ from the above specified model.
+For each dataset, we then fitted four competing models:
+  1. 1.
+    Correct model: the model used to generate the data, with correctly specified parameters.
+  2. 2.
+    Shared-$\alpha$ model: a simplified model in which the link parameters $\alpha^{k\to k^{\prime}}$ are constrained to be equal across all transitions, i.e., $\alpha^{1\to 1}=\alpha^{1\to 2}=\alpha$.
+  3. 3.
+    Expanded-$\beta$ model: a slightly more complex model in which three additional normally distributed covariates are added to each individual without any association.
+  4. 4.
+    Link misspecified: a model in which the link functions are incorrectly specified as the bi-exponential function instead of the cumulative exposure.
+For each run, the model with the smallest value of AIC or $\mathrm{BIC}_{\mathcal{I}}$ was selected. Table 6 summarizes the results across the $100$ simulations.
+Model AIC selections $\mathrm{BIC}_{\mathcal{I}}$ selections
+Correct 69 99
+Expanded-$\beta$ 31 1
+Shared-$\alpha$ 0 0
+Misspecified link 0 0
+Figure 6: Number of times each model was selected based on AIC and $\mathrm{BIC}_{\mathcal{I}}$ over $100$ simulation runs.
+The results indicate that both information criteria successfully identify the correct model in the majority of cases.
+However, $\mathrm{BIC}_{\mathcal{I}}$ favors the correctly specified model far more frequently than AIC, aligning with its stronger penalty for model complexity, and is able to correctly identify the model $99\%$ of the times.
+Indeed, the increase in the AIC penalty for the expanded model is only $6$ relative to the correctly specified model.
+Yet the negative log-likelihood is of the order of 2 0002\,000.
+In such a regime, the additional penalty appears negligible compared with the scale of the likelihood term, so the AIC does not provide a sufficient criterion in this situation.
+Both the shared-$\alpha$ model and the misspecified link models were never selected by the AIC nor the $\mathrm{BIC}_{\mathcal{I}}$.
+Overall, these findings demonstrate that our model selection criteria very effectively distinguish between correctly specified and misspecified multi-state joint models, providing reliable guidance for model choice in finite-sample settings.
 7 Application to the PAQUID Cohort
-7.1 Introduction
 The data used in this section originates from the PAQUID cohort [Letenneur1994], a large prospective population-based study initiated in southwestern France in 1988, aimed at understanding the determinants and trajectories of aging.
 A subsample of $n=500$ individuals was followed over a period of up to $20$ years [Proust-Lima2017], with repeated measures of cognitive and physical health, as well as socio-demographic characteristics.
-In particular, global cognitive functioning was assessed through the Mini-Mental State Examination (MMSE, see Figure 6), while physical dependency was evaluated using the HIER scale, which classifies subjects into four ordered states (see Figure 7): 0 (no dependency), $1$ (mild dependency), $2$ (moderate dependency), and $3$ (severe dependency).
+In particular, global cognitive functioning was assessed through the Mini-Mental State Examination (MMSE), while physical dependency was evaluated using the HIER scale, which classifies subjects into four ordered states (see Figure 7): 0 (no dependency), $1$ (mild dependency), $2$ (moderate dependency), and $3$ (severe dependency).
+Individual covariates correspond to a pair of binary variables indicating whether or not the individual has a diploma and is a male $\left(\mathds{1}_{\mathrm{CEP}}(i),\mathds{1}_{\mathrm{male}}(i)\right)\in\{0,1\}^{2}$.
 In this work, we focus on the association between cognitive decline and the progression of physical dependency by jointly modeling the longitudinal trajectory of MMSE scores and the transitions between the four HIER states using a joint multi-state model.
-More specifically, the state of an individual $i$ at a given time $t$ is defined as the highest HIER dependency score between trial entry and $t$.
+More specifically, the state of an individual at a given time $t$ is defined as the highest HIER dependency score between trial entry and $t$, and progression is therefore monotone.
 This approach allows us to characterize the dynamic interrelationship between cognitive and physical aging processes within a unified statistical framework.
-Both time and longitudinal values were normalized in [1,1][-1,1]. This technique ensured stable and fast convergence of the optimization process by keeping all parameters on roughly the same scale.
-From $\rightarrow$ To # Observations
-$0\rightarrow 1$ 85
-$1\rightarrow 2$ 135
-$2\rightarrow 3$ 86
-$0\rightarrow 2$ 44
-$1\rightarrow 3$ 15
-$0\rightarrow 3$ 11
-Figure 6: On the left, each light gray line represents one of the $500$ observed sequences of (normalized) MMSE scores, and three functional clusters are also represented; on the right, observed transitions between states.
-Therefore, the chosen transition graph is as follows, and only allows monotonic transitions from a lower level of physical dependency to a higher one.
+7.1 Preprocessing
+Longitudinal values were normalized to the interval $[0,1]$ by dividing each score by its maximum value of $30$, and the measurement times were normalized to represent the elapsed time since age $65$ in decades.
+This technique ensured stable and fast convergence of the optimization process by keeping all parameters on roughly the same scale.
+7.2 Model Specification
+The transition graph was chosen as follows in Figure 7, and only allows monotonic transitions from a lower level of physical dependency to a higher one, as some transitions occur so infrequently that they cannot be meaningfully analyzed.
+We considered Weibull baseline hazards for all transitions, consistent with previous observations by Proust-Lima2017, with a clock-reset (sojourn time) specification.
+The parameters of these baseline hazards were jointly optimized with the model parameters.
+It should also be noted that not all patients begin in state 0, thus, in this respect, our model inherently accounts for left censoring.
 Figure 7: Schematic representation of the four ordered HIER states, with possible transitions. HIER quantifies physical dependency (0: no dependency to 33: severe dependency) [Proust-Lima2017].
-We consider an exponential baseline hazard model for transitions, with a clock reset specification.
-The parameters of these exponential baseline hazards are jointly optimized with the model parameters.
-Moreover, we impose a particular substructure on the model, where $\forall(k,k^{\prime})\in E,\,\beta^{k^{\prime}\mid k}=\beta$.
-In the light of Figure 6, the regression and link functions were taken to be scaled and shifted hyperbolic tangent and its derivative respectively such that the normalized MMSE score always has a value of $1$ when $t\to$-$\infty$ and is not increasing
-$\begin{cases}h(t,\psi)=\psi_{1}\tanh\left(\frac{\psi_{3}-t}{\psi_{2}}\right)+(1-\psi_{1}),\\ g(t,\psi)=\frac{\partial}{\partial t}h,\\ \psi=\begin{pmatrix}\sigma(\gamma_{1}+b_{1})\\ \exp(\gamma_{2}+b_{2})\\ \gamma_{3}+b_{3}\end{pmatrix},\quad\sigma(z)=\frac{1}{1+e^{-z}}.\end{cases}$
-At last, an individual covariates correspond to a pair of binary variables indicating wheter or not the individual has a diploma and is a male $\left(\mathds{1}_{\text{CEP}}(i),\mathds{1}_{\text{male}}(i)\right)\in\{0,1\}^{2}$.
+The regression and link functions were taken to be shifted and scaled logistic curve such that the normalized MMSE score is always is non-increasing
+[…]
+For every transition $(k,k^{\prime})\in E$, we define $g^{k\to k^{\prime}}:=g$.
+The parameter $\Psi_{1}$ governs the amplitude of the logistic function, $\Psi_{2}$ introduces a time shift, and $\Psi_{3}$ controls the slope at the inflection point.
+Moreover, we imposed a specific substructure on the model by setting $\forall(k,k^{\prime})\in E,\,\beta^{k\to k^{\prime}}=\beta$, which helped mitigate issues arising from the small number of observations in some groups.
+Additionally, since no prior information was available for the covariance matrix of the latent effects $Q\in\mathbb{R}^{3\times 3}$, we used a full $\log$-Cholesky parametrization of the precision matrix $Q^{-1}$, with six parameters denoted $\tilde{Q}_{1}$ through $\tilde{Q}_{6}$.
+The full parametrization is essential to capture the potential cross-dependencies between the three latent effects.
 7.3 Results
+In this study, the full dataset was used to estimate the inferential results presented in Subsection 7.3.1.
+For the dynamic prediction analysis in Subsection 7.3.2, we employed a 5-fold cross-validation procedure, corresponding to splits of $400$ observations for model training and $100$ for validation in each fold, as to obtain unbiased estimates of the predictive performance metrics.
 7.3.1 Inference
-Fitting was performed using $80\%$ of the data, with the remaining $20\%$ used for testing.
-The model was fitted using the Adam optimizer.
-The optimization process was run until convergence based on the same criterion as the simulation study with a tolerance of $5\%$, which was obtained after a little less than $500$ iterations with $10$ parallel chains.
-Before displaying the fitted parameters, the Fisher Information Matrix was computed, implementing the method described in [baey2023efficient], directly taking into account shared parameters.
+The model was fitted on the complete dataset using the Adam optimizer and the optimization process was run until convergence based on the Equation 6.2 as in the simulation study.
+The Fisher Information Matrix was computed, implementing the method described in [baey2023efficient].
 The standard errors were then computed as the square root of the diagonal of the inverse of the Fisher Information Matrix.
-The results are summarized in Table 8.
+Since, under the chosen parametrization, the value 0 lies in the interior of the admissible parameter space, Wald statistics can be used to test the nullity of each coefficient through the usual asymptotic normal approximation.
+However, we report only the significance results for the linear coefficients coefficients in either the covariates or the link functions, as they allow a direct interpretation.
+They are indicated in bold in the summarized Table 8.
+The optimization process itself is provided in the Appendix (see Figure 11).
 Parameter Inferred value Std. error Parameter Inferred value Std. error
-$\gamma_{1}$ 2.1600 0.0470 $\gamma_{2}$ 0.1590 0.0140
-$\gamma_{3}$ 1.0010 0.0030 $\tilde{Q}_{1}$ -0.5700 0.0430
-$\tilde{Q}_{2}$ 0.9570 0.0960 $\tilde{Q}_{3}$ 0.3820 0.0430
-$\tilde{Q}_{4}$ -0.2050 0.1570 $\tilde{Q}_{5}$ -2.3310 0.1380
-$\tilde{Q}_{6}$ 0.5220 0.0470 $\tilde{R}$ 2.0830 0.0260
-$\alpha^{3\mid 1}$ 0.7220 0.7150 $\alpha^{1\mid 0}$ -0.5400 0.4030
-$\alpha^{2\mid 1}$ -1.0030 0.2000 $\alpha^{2\mid 0}$ -0.1210 0.0380
-$\alpha^{3\mid 2}$ -0.0480 0.2520 $\alpha^{3\mid 0}$ -0.3070 0.0740
-$\beta_{1}$ -0.2990 0.9130 $\beta_{2}$ 0.0970 1.0300
-Figure 8: Inferred parameters and their estimated standard deviations. Bold values mean $p<0.05$ (variance parameters excluded).
-Notably, the estimated standard errors are consistent with the number of observations per transition, with larger standard errors for transitions with fewer observations.
-Therefore, parameters with too few observations should be interpreted with caution, if at all.
-The linear covariate parameters also exhibit large standard errors, which indicates the coefficients are not significative in this setting.
-The association parameters $\alpha$ being negative, we can infer that cognitive decline (i.e. rate of change) is associated with an increase in physical dependency.
-7.2.2 Dynamic Prediction
-Using the $n_{\text{test}}=100$ individuals not included in the model fitting, we performed dynamic prediction.
-For each individual, longitudinal measurements and trajectories were truncated at various times $t$, and predictions were made for future time points $u\geq t$ beyond the truncation.
-Accuracy was then assessed by comparing the most likely predicted state with the true state at each corresponding future time point $u$, as illustrated in Figure 9.
-Formally, the prediction function $\chi_{u}^{*}$ is defined by Example 1, and the accuracy measure at time $t$ for predicted states $(s_{i})_{i}$ is defined as
-$\text{accuracy}(u)=\frac{1}{n_{\text{test}}}\sum_{i=1}^{n_{\text{test}}}\mathds{1}_{\chi_{u\wedge C_{i}}(\mathcal{T}_{i})=s_{i}}.$
-Although $\mathcal{T}_{i}^{*}$ is unkown, the quantity $\chi_{u\wedge C_{i}}(\mathcal{T}_{i})$ may be computed based on the censored trajectories.
-However, incorporating censoring times into prediction functions relies on information that is, in principle, not available at prediction time.
-Nevertheless, this inclusion is necessary to ensure fairness in accuracy metrics, since individuals who are not censored are generally more likely to have progressed to more advanced states of physical dependency than those who are censored.
-Figure 9: Accuracy of the predicted state at each time point with respect to the true state for different truncation times.
-Here, the initial drops correspond to the truncation times, before which the accuracy is always $100\%$.
-As expected from dynamic predictions, the accuracy increases as more data is available, therefore when the truncation time grows.
-We also observe that the longer the prediction horizon, the lower the accuracy of the predictions.
+$\gamma_{1}$ 1.201 0.035 $\boldsymbol{\gamma_{2}}$ 0.221 0.042
+$\boldsymbol{\gamma_{3}}$ 0.155 0.041 $\gamma_{4}$ 3.321 0.076
+$\gamma_{5}$ 0.181 0.094 $\boldsymbol{\gamma_{6}}$ 0.568 0.119
+$\gamma_{7}$ 0.636 0.188 $\boldsymbol{\gamma_{8}}$ 0.663 0.229
+$\boldsymbol{\gamma_{9}}$ 1.168 0.233 $\tilde{Q}_{1}$ 2.537 0.079
+$\tilde{Q}_{2}$ -0.032 0.321 $\tilde{Q}_{3}$ 0.644 0.086
+$\tilde{Q}_{4}$ -1.909 0.134 $\tilde{Q}_{5}$ -0.752 0.100
+$\tilde{Q}_{6}$ -0.690 0.065 $\frac{1}{2}\log(R)$ 3.023 0.008
+$\log\lambda^{0\to 1}$ -1.174 1.090 $\log k^{0\to 1}$ 0.269 0.125
+$\log\lambda^{0\to 2}$ 1.160 0.585 $\log k^{0\to 2}$ 0.648 0.142
+$\log\lambda^{0\to 3}$ 1.663 0.769 $\log k^{0\to 3}$ 0.876 0.352
+$\log\lambda^{1\to 2}$ 1.260 0.519 $\log k^{1\to 2}$ 0.434 0.097
+$\log\lambda^{1\to 3}$ 2.744 0.926 $\log k^{1\to 3}$ 0.520 0.283
+$\log\lambda^{2\to 3}$ 2.239 0.346 $\log k^{2\to 3}$ 0.509 0.107
+$\alpha^{0\to 1}$ 1.254 1.614 $\boldsymbol{\alpha^{0\to 2}}$ -3.844 1.281
+$\boldsymbol{\alpha^{0\to 3}}$ -8.487 1.362 $\boldsymbol{\alpha^{1\to 2}}$ -2.167 0.898
+$\boldsymbol{\alpha^{1\to 3}}$ -8.393 0.957 $\boldsymbol{\alpha^{2\to 3}}$ -5.315 0.516
+$\beta_{1}$ -0.159 0.315 $\beta_{2}$ 0.152 0.233
+Figure 8: Inferred parameters and their estimated standard deviations. Bold values mean $p<0.05$ (only the linear coefficients included).
+All link parameters $\alpha^{k\to k^{\prime}}$ for transitions $(k,k^{\prime})\in E$ are negative and statistically significant, except that from state 0 to $1$, and the most negative coefficients correspond to transitions reflecting rapid deterioration ($0\to 3$, $1\to 3$, and $2\to 3$).
+This pattern indicates that the current value of the MMSE score strongly influences the risk of progressing toward physical dependence, and lower cognitive performance is associated with a substantially increased transition intensity, especially for direct or fast declines.
+Consistent with other analyses conducted on this dataset, both education level (CEP) and female gender are associated with more favorable cognitive and functional trajectories.
+Specifically, the inferred values of $\gamma_{2}$, $\gamma_{5}$, and $\gamma_{8}$ indicate that individuals with a primary school certificate (CEP) not only start from a significantly higher baseline level of MMSE, but also experience a delayed and slower progression of dementia.
+Similarly, the positive values of $\gamma_{3}$, $\gamma_{6}$, and $\gamma_{9}$ suggest the same remarks for women relative to men.
+Regarding the association coefficient $\beta$, the relatively large standard errors imply that the remaining effects are modest and that the link function already captures most of the relevant information.
+7.3.2 Multi-State Dynamic Prediction
+After fitting the model on each training fold, we performed multi-state dynamic prediction on the corresponding validation fold.
+For each validation fold $V_{k}$, with $k\in\llbracket 5\rrbracket$, predictions were computed using a single Monte Carlo integration scheme as described in Section 5.
+We assessed the performance of our model using two metrics: the Brier Skill Score (BSS), which is one of the most standard metric in this context, as well as the accuracy.
+More specifically, for each new individual $i$, longitudinal measurements and trajectories were truncated at various landmarks $t$, and predictions were made for future time points $u\geq t$ beyond each truncation.
+We considered the quantity $\chi_{u}^{*}(\mathcal{T}_{i}^{*})$ as defined by Example 1, which is estimated using the information available up to $t$, i.e., $\left(\mathcal{Y}_{i}(t),\mathcal{S}_{i}(t)\right)$.
+Because the complete trajectories are not fully observed in the test sample and the quantity $\chi_{u}^{*}(\mathcal{T}_{i}^{*})$ is not available for $u>C_{i}$, predictions must account for censoring.
+For a given landmark time $t$, we obtain a collection of probability estimates
+$\{\hat{\chi}_{i,u}(t)\}\subset\Delta\left(\{0,1,2,3\}\right),$
+each computed from the information available up to $t$. The Brier Skill Score for the fold $V_{k}$ with $k\in\llbracket 5\rrbracket$ is then defined by
+$\mathrm{BSS}_{k}(t,u)\coloneqq 1-\frac{\mathrm{BS}_{k}(t,u)}{\mathrm{BS}_{k}^{(0)}(u)},$
+where
+$\begin{cases}\mathrm{BS}_{k}(t,u)\coloneqq\frac{1}{\#V_{k}}\sum_{i\in V_{k}}\left\|\hat{\chi}_{i,u\wedge C_{i}}(t)-\chi_{u\wedge C_{i}}^{*}(\mathcal{T}_{i}^{*})\right\|^{2},\\ \mathrm{BS}_{k}^{(0)}(u)\coloneqq\min_{p}\frac{1}{\#V_{k}}\sum_{i\in V_{k}}\left\|p-\chi_{u\wedge C_{i}}^{*}(\mathcal{T}_{i}^{*})\right\|^{2}.\end{cases}$
+Hence, values of $\mathrm{BSS}_{k}(t,u)$ close to $1$ indicate a substantial improvement over the optimal constant prediction.
+Similarly, the accuracy at $(t,u)$ on the fold $V_{k}$ is defined as
+$\mathrm{Acc}_{k}(t,u)\coloneqq\frac{1}{\#V_{k}}\sum_{i\in V_{k}}\mathds{1}\left\{\operatorname*{\arg\max}_{j\in\{0,1,2,3\}}\left(\hat{\chi}_{i,u\wedge C_{i}}(t)\right)_{j}=\operatorname*{\arg\max}_{j\in\{0,1,2,3\}}\left(\chi_{u\wedge C_{i}}^{*}(\mathcal{T}_{i}^{*})\right)_{j}\right\},$
+which corresponds to the proportion of correctly identified states.
+Figure 9 illustrates our dynamic predictions framework evaluated at three distinct landmark times using both metrics estimated from our cross-validation procedure.
+Both metrics are averaged across the $5$ folds.
+Figure 9: Average values on the 55-fold cross-validation of the Brier Skill Score (top) and accuracy (bottom) at each time point for different prediction landmarks (dotted lines).
+Colored bands correspond to $\mathrm{mean}\pm 1.65\mathrm{sd}$.
+Across the three prediction landmarks, the Brier Skill Score remains above $40\%$ at $10$ years beyond the prediction landmark.
+This indicates a very strong improvement over the best constant baseline model, despite the limited information available at the time of prediction.
+Similarly, the classification accuracy consistently exceeds $60\%$, often higher, thus the model correctly identifies the majority of states.
+This level of performance persists even for long-term horizons extending more than $20$ years into the future, including during the early phases of the study.
+As expected, predictions made immediately following each truncation are the most precise, while uncertainty naturally increases as the prediction horizon extends further.
+Both metrics also get substantially better as the prediction landmarks include more data.
+Both the Brier Skill Score (BSS) and classification accuracy remain robust over time, providing strong evidence of the reliability and stability of our predictive framework.
 8 Conclusion
@@
 Simulation studies confirmed robust parameter recovery, and an application to the PAQUID cohort highlighted the ability to capture the interplay between cognitive decline and physical dependency.
-Since the model may potentially involve many transitions and therefore a large number of parameters, a strategy of parameter sharing across transitions could help reduce the overall number of parameters.
+Since the model may potentially involve many transitions, and therefore a large number of parameters, a strategy of parameter sharing across transitions could help reduce the overall number of parameters.
 This is especially beneficial in settings with limited data for certain transitions, or when biological or clinical knowledge suggests similar effects across multiple transitions.
@@
 This may be achieved through bisection or other root-finding algorithms.
-The simulation of the transition process $\mathcal{T}_{i}^{*}$ conditionally on $b_{i}$, $X_{i}$, and $(T_{i0},S_{i0})$ can be achieved drawing on the semi-Markov property by considering one transition at a time, according to Algorithm 2.
-The algorithme is similar to Gillespie’s algorithm [gillespie2007stochastic], and can also include a survival condition $T_{i1}\geq t_{i}^{\text{surv}}$ that uses the Chasles relation.
+The simulation of the transition process $\mathcal{T}_{i}^{*}$ conditionally on $b_{i}$, $X_{i}$, and $(T_{i0},S_{i0})$ can be achieved drawing on the semi-Markov property 1 by considering one transition at a time, according to Algorithm 3.
+The algorithm is similar to Gillespie’s algorithm [gillespie2007stochastic], and can also include a survival condition $T_{i1}\geq t_{i,\mathrm{cond}}$ that uses the Chasles relation.
+Here, we deliberately avoid imposing any conditioning in order to preserve maximal generality, since the algorithm is intended to be applicable beyond the specific framework of our model, only assuming continuous transition times distributions and the semi-Markov property.
 Algorithm 3 Simulation of trajectory $i$
-1:subject-specific censoring time $C_{i}\in\mathbb{R}\cup\{+\infty\}$; covariates $X_{i}$; latent variables $b_{i}$; initial time-state pair $(T_{i0},S_{i0})$; optional survival condition $T_{i1}\geq t_{i}^{\text{surv}}$, defaults to $t_{i}^{\text{surv}}=$-$\infty$.
+1:subject-specific censoring time $C_{i}\in\bar{\mathbb{R}}^{+}$; initial time-state pair $(T_{i0},S_{i0})$; optional survival condition $T_{i1}\geq t_{i,\mathrm{cond}}$, defaults to $t_{i,\mathrm{cond}}=T_{i0}$.
 2:
-3:Initialize $\ell_{i}\leftarrow 1$, $\mathcal{T}_{i}\leftarrow\left((T_{i0},S_{i0})\right)$.
+3:Initialize $\ell\leftarrow 1$, $\mathcal{T}_{i}\leftarrow\left((T_{i0},S_{i0})\right)$
 4:
-5:while $T_{i(\ell_{i}-1)}<C_{i}$ and $\{s:(S_{i(\ell_{i}-1)},s)\in E\}\neq\emptyset$ do
-6:  for each $s$ with $(S_{i(\ell_{i}-1)},s)\in E$ do
-7:   Draw $T_{i\ell_{i}}^{s}$ such that
-8:      $\log\mathbb{P}(T_{i\ell_{i}}^{s}>t)=\int_{T_{i(\ell_{i}-1)}\vee t_{i}^{\text{surv}}}^{t}\lambda_{i}^{s\mid S_{i(\ell_{i}-1)}}(w\mid b_{i},X_{i},T_{i(\ell_{i}-1)},\theta)\,dw,$
-9:      $\forall t\geq T_{i(\ell_{i}-1)}\vee u.$
-10:  end for
-11:
-12:  Set $T_{i\ell_{i}}\leftarrow\min_{m}T_{i\ell_{i}}^{s}$, and $S_{i\ell_{i}}\leftarrow\arg\min_{m}T_{i\ell_{i}}^{s}$.
-13:  Append $(T_{i\ell_{i}},S_{i\ell_{i}})$ to $\mathcal{T}_{i}$.
-14:  $\ell_{i}\leftarrow\ell_{i}+1$.
-15:end while
-16:
-17:if $T_{i(\ell_{i}-1)}>C_{i}$ then
-18:  Remove the last pair: $\mathcal{T}_{i}\leftarrow\mathcal{T}_{i}[:-1]$.
-19:end if
-20:return $\mathcal{T}_{i}$
+5:while $T_{i(\ell-1)}<C_{i}$ and $\{j:(S_{i(\ell-1)},j)\in E\}\neq\emptyset$ do
+6:  for each $j$ with $(S_{i(\ell-1)},j)\in E$ do
+7:   Draw $T_{i\ell}^{(j)}$ such that $\forall t\geq T_{i(\ell-1)}\vee t_{i,\mathrm{cond}}$
+8:      $\log\mathbb{P}(T_{i\ell}^{(j)}>t)=\int_{T_{i(\ell-1)}\vee t_{i,\mathrm{cond}}}^{t}\lambda_{i}^{S_{i(\ell-1)}\to j}(s\mid T_{i(\ell-1)})\,ds$
+9:  end for
+10:
+11:  Set $T_{i\ell}\leftarrow\min_{j}T_{i\ell}^{(j)}$, and $S_{i\ell}\leftarrow\arg\min_{j}T_{i\ell}^{(j)}$
+12:  Append $(T_{i\ell},S_{i\ell})$ to $\mathcal{T}_{i}$
+13:  $\ell\leftarrow\ell+1$
+14:end while
+15:
+16:if $T_{i(\ell-1)}>C_{i}$ then
+17:  Remove the last pair: $\mathcal{T}_{i}\leftarrow\mathcal{T}_{i}[:-1]$
+18:end if
+19:return $\mathcal{T}_{i}$
 The proof of the exactness of this algorithm is given below.
 Proof.
-First consider the case $t_{i}^{\text{surv}}=$-$\infty$. Let $t>T_{i\ell_{i}}$:
-$\mathbb{P}(T_{i(\ell_{i}+1)}>t\mid T_{i\ell_{i}},S_{i\ell_{i}})$ $=\mathbb{P}(\bigcap_{s:\,(E_{i(\ell_{i}-1)},s)\in E}T_{i(\ell_{i}+1)}^{s}>t\mid T_{i\ell_{i}},S_{i\ell_{i}}),$
-$=\prod_{s:\,(S_{i\ell_{i}},s)\in E}\mathbb{P}(T_{i(\ell_{i}+1)}^{s}>t\mid T_{i\ell_{i}},S_{i\ell_{i}}),$
-$=\exp\Big(-\sum_{s:\,(S_{i\ell_{i}},s)\in E}\Lambda_{i}^{s\mid S_{i(\ell_{i}-1)}}\big(t\mid b_{i},X_{i},T_{i\ell_{i}},\theta)\Big).$
-Furthermore:
-$\mathbb{P}(S_{i(\ell_{i}+1)}\mid T_{i(\ell_{i}+1)}=t,T_{i\ell_{i}},S_{i\ell_{i}})$
-$\propto p(\{T_{i(\ell_{i}+1)}^{S_{i(\ell_{i}+1)}}=t\}\cap\bigcap_{\begin{subarray}{c}s:\,(S_{i\ell_{i}},s)\in E,\\ s\neq S_{i(\ell_{i}+1)}\end{subarray}}\{T_{i(\ell_{i}+1)}^{s}\geq t\}\mid T_{i\ell_{i}},S_{i\ell_{i}}),$
-$=\lambda_{i}^{S_{i(\ell_{i}+1)}\mid S_{i\ell_{i}}}\big(t\mid\theta,X_{i},T_{i\ell_{i}},b_{i})\exp\Big(-\sum_{s:\,(S_{i\ell_{i}},s)\in E}\Lambda_{i}^{s\mid S_{i\ell_{i}}}\big(t\mid b_{i},X_{i},T_{i\ell_{i}},\theta)\Big).$
-Combining both, we retrieve the same joint density.
-If $t_{i}^{\text{surv}}>$-$\infty$, one can also check that both densities are equal.
-Note that the condtioning only affects the first transition, as the next time $\forall\ell_{i}>0,\,T_{i\ell_{i}}\geq t_{i}^{\text{surv}}\implies\forall\ell_{i}>0,\,T_{i\ell_{i}}\vee t_{i}^{\text{surv}}=T_{i\ell_{i}}$.
+First, consider the case $t_{i,\mathrm{cond}}=T_{i0}$. Let $\ell\geq$ and $t\geq T_{i\ell}$. One has
+$\mathbb{P}(T_{i(\ell+1)}>t\mid T_{i\ell},S_{i\ell})$ $=\mathbb{P}\left(\bigcap_{j:(S_{i(\ell-1)},j)\in E}T_{i(\ell+1)}^{(j)}>t\mid T_{i\ell},S_{i\ell}\right),$
+$\overset{\mathrel{\perp\!\!\!\perp}}{=}\prod_{j:(S_{i\ell},j)\in E}\mathbb{P}(T_{i(\ell+1)}^{(j)}>t\mid T_{i\ell},S_{i\ell}),$
+$=\exp\left(-\sum_{j:(S_{i\ell},j)\in E}\Lambda_{i}^{S_{i(\ell-1)}\to j}(T_{i\ell},t)\right).$
+Furthermore, for any $k^{\prime}\in V$ such that $(S_{i\ell},k^{\prime})\in E$, calling $J_{i\ell}^{(k^{\prime})}\coloneqq\{j:(S_{i\ell},j)\in E\}\setminus\{k^{\prime}\}$, we have
+$\mathbb{P}(S_{i(\ell+1)}=k^{\prime}\mid T_{i(\ell+1)}=t,T_{i\ell},S_{i\ell})$
+$\propto p\left(T_{i(\ell+1)}^{(k^{\prime})}=t,\forall j\in J_{i\ell}^{(k^{\prime})}\,T_{i(\ell+1)}^{(j)}\geq t\mid T_{i\ell},S_{i\ell}\right)$
+$=\lambda_{i}^{S_{i\ell}\to k^{\prime}}(T_{i\ell},t)\exp\left(-\sum_{j:(S_{i\ell},j)\in E}\Lambda_{i}^{S_{i\ell}\to j}(T_{i\ell},t)\right)$
+$\propto\lambda_{i}^{S_{i\ell}\to k^{\prime}}(T_{i\ell},t).$
+Combining both, we retrieve the same joint density as in Equation 2.
+If $t_{i,\mathrm{cond}}>T_{i0}$, one can also check that both densities are equal by simply integrating from $t_{\mathrm{cond}}$ instead of $T_{i\ell}$.
+Note that the condtioning only affects the first transition, as the next time $\forall\ell>0,\,T_{i\ell}\geq t_{i,\mathrm{cond}}\implies\forall\ell>0,\,T_{i\ell}\vee t_{i,\mathrm{cond}}=T_{i\ell}$.
+∎
 Appendix B Expression of the Semi-Markov likelihood
 Proof.
-The proof is short and relies on the fact that, using the semi-Markov property 2
-$p\left((T_{i(\ell+1)},S_{i(\ell+1)})_{\ell\leq m_{i}-1}\mid b_{i},X_{i},\theta\right)$ $=\prod_{\ell=0}^{m_{i}-1}p\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid b_{i},X_{i},(T_{i\ell^{\prime}},S_{i\ell^{\prime}})_{\ell^{\prime}\leq\ell},\theta\right),$
-$=\prod_{\ell=0}^{m_{i}-1}p\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid b_{i},X_{i},(T_{i\ell},S_{i\ell}),\theta\right).$
-On the other hand, the probability that no additional event is observed between $T_{im_{i}}$ and $C_{i}$ is $1$ if the last state reached is absorbing, and otherwise
-$\mathbb{P}(T_{i(m_{i}+1)}>C_{i}\mid\theta,X_{i},(T_{im_{i}},S_{im_{i}}),b_{i})=\exp\left(-\sum_{s:\,(S_{im_{i}},s)\in E}\Lambda_{i}^{s\mid S_{im_{i}}}(C_{i}\mid b_{i},X_{i},T_{im_{i}},\theta)\right),$
-and so we recover the formula cited above with the convention that an empty sum is 0.
-Appendix C Stochastic optimization process
-(a) $\gamma$
-(b) $Q$
-(c) $R$
-(d) $\alpha^{1\mid 0}$
-(e) $\alpha^{2\mid 0}$
-(f) $\alpha^{2\mid 1}$
-(g) $\beta^{1\mid 0}$
-(h) $\beta^{2\mid 0}$
-(i) $\beta^{2\mid 1}$
-Figure 10: Evolution of the parameters during the optimization of the marginal log-likelihood using stochastic gradient descent. The dotted lines correspond to the true values.
+The proof is short and relies on the fact that, using the semi-Markov property 1
+$p_{\theta}\left((T_{i(\ell+1)},S_{i(\ell+1)})_{\ell\leq m_{i}-1}\mid b_{i},X_{i}\right)$ $=\prod_{\ell=0}^{m_{i}-1}p_{\theta}\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid b_{i},X_{i},(T_{i\ell^{\prime}},S_{i\ell^{\prime}})_{\ell^{\prime}\leq\ell}\right),$
+$=\prod_{\ell=0}^{m_{i}-1}p_{\theta}\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid b_{i},X_{i},(T_{i\ell},S_{i\ell})\right).$
+On the other hand, the probability that no additional event is observed between $T_{im_{i}}$ and $C_{i}$ is $1$ if the last state reached is absorbing, and otherwise, since censoring is assumed non-informative, it can be treated as an independent constant
+$\mathbb{P}_{\theta}\left(T_{i(m_{i}+1)}>C_{i}\mid b_{i},X_{i},(T_{im_{i}},S_{im_{i}})\right)=\exp\left(-\sum_{j:(S_{im_{i}},j)\in E}\Lambda_{i,\theta}^{S_{im_{i}}\to j}(T_{im_{i}},C_{i}\mid b_{i},X_{i})\right),$
+so we recover the formula cited above, with the convention that an empty sum is 0. ∎
+Appendix C Stochastic optimization process on simulated data
+Figure 10: Evolution of the parameters during the optimization of the marginal $\log$-likelihood using stochastic gradient ascent.
+Dotted horizontal lines correspond to the true values, while vertical lines mark the start of each window used to evaluate the stopping criterion.
+Appendix D Stochastic optimization process on the PAQUID dataset
+Figure 11: Evolution of the parameters during the optimization of the marginal $\log$-likelihood using stochastic gradient ascent.
+Vertical lines mark the start of each window used to evaluate the stopping criterion.